### v1

# Singular Bayesian Neural Networks

Abstract
Bayesian neural networks promise calibrated uncertainty but require $O(mn)$ parameters for standard mean-field Gaussian posteriors. We argue this cost is often unnecessary, particularly when weight matrices exhibit fast singular value decay. By parameterizing weights as $W=AB^{\top}$ with $A\in\mathbb{R}^{m\times r}$, $B\in\mathbb{R}^{n\times r}$, we induce a posterior that is singular with respect to the Lebesgue measure, concentrating on the rank-$r$ manifold. This singularity captures structured weight correlations through shared latent factors, geometrically distinct from mean-field’s independence assumption. We derive PAC-Bayes generalization bounds whose complexity term scales as $\sqrt{r(m+n)}$ instead of $\sqrt{mn}$, and prove loss bounds that decompose the error into optimization and rank-induced bias using the Eckart-Young-Mirsky theorem. We further adapt recent Gaussian complexity bounds for low-rank deterministic networks to Bayesian predictive means. Empirically, across MLPs, LSTMs, and Transformers on standard benchmarks, our method achieves predictive performance competitive with 5-member Deep Ensembles while using up to $15\times$ fewer parameters. Furthermore, it substantially improves OOD detection and often improves calibration relative to mean-field and perturbation baselines.
Machine Learning, ICML
1 Introduction
Bayesian neural networks (BNNs) provide principled uncertainty quantification by maintaining distributions over weights rather than point estimates (MacKay, 1992; Neal, 1996). This capability is increasingly critical for high-stakes applications (Amodei et al., 2016) across healthcare (Ruhe et al., 2019; Dusenberry et al., 2020a, b), autonomous systems (Kendall and Gal, 2017), and domains requiring robust generalization under distribution shift (Hendrycks and Gimpel, 2016; Louizos and Welling, 2017; Ovadia et al., 2019; Abdar et al., 2021). Recent calls for Bayesian approaches in large-scale AI (Papamarkou et al., 2024) underscore the growing recognition that uncertainty-aware models are essential for trustworthy deployment.
However, scaling BNNs to modern architectures remains challenging. Exact Bayesian inference is intractable for neural networks, necessitating approximate methods such as variational inference (Peterson and Anderson, 1987; Hinton and Van Camp, 1993; Jordan et al., 1999) or MCMC sampling (Neal, 1996). While MCMC methods and their modern stochastic variants like SGLD (Welling and Teh, 2011) and stochastic HMC (Chen et al., 2014) provide theoretical guarantees, they require costly iterative sampling that is prohibitive for large-scale models. On the other hand, traditional Mean-field variational inference (MFVI) implementations (Graves, 2011; Kingma and Welling, 2014; Blundell et al., 2015) parameterize each weight with a location-scale family (e.g., Gaussian $\mathcal{N}(\mu,\sigma^{2})$), requiring two variational parameters per weight thus doubling the parameter count relative to deterministic networks. Additionally, they impose a fully factorized posterior discarding the structured correlations that may be important for expressiveness, performance and theoretical guarantees. Recent work has further revealed that weight-space inference in modern architectures such as transformers suffers from fundamental pathologies related to prior specification and the difficulty of mapping weight-space to function-space distributions (Cinquin et al., 2021). These limitations have largely confined BNNs to small-scale problems, while modern deep neural networks, from convolutional architectures (He et al., 2016) and LSTMs (Hochreiter and Schmidhuber, 1997) to transformers (Vaswani et al., 2017), routinely scale to billions of parameters.
Recent work has begun addressing scalability through low-rank parameterizations. Indeed, neural networks have been observed to exhibit low intrinsic dimensionality (Hinton and Van Camp, 1993; Li et al., 2018; Izmailov et al., 2020; Aghajanyan et al., 2021), suggesting that full-rank parameterizations may be unnecessarily expensive. Current approaches to exploiting this structure fall into three camps: (i) post-hoc low-rank perturbations (Dusenberry et al., 2020a; Doan et al., 2025) that add stochastic noise to deterministic backbones, (ii) low-rank posterior covariance approximations (Tomczak et al., 2020; Swiatkowski et al., 2020) that enrich mean-field VI by replacing diagonal covariances with low-rank-plus-diagonal structures while maintaining full-rank weight parameterizations, and (iii) parameter-efficient low-rank fine-tuning methods like LoRA (Hu et al., 2022) for fine-tuning pretrained models which inspired bayesian variants (Yang et al., 2024; Meo et al., 2024; Wang et al., 2024). However, all three approaches either require pretrained backbones, sacrificing the end-to-end learning of uncertainty, or, in the case of covariance methods, still parameterize $\mathcal{O}(mn)$ weight means rather than directly exploiting low-rank matrix structure. An in-depth comparison with similar approaches is provided in Appendix A.
To the best of our knowledge, no prior work trains low-rank BNNs end-to-end across diverse architectures with rigorous theoretical guarantees.
We introduce a fully end-to-end low-rank variational inference framework for BNNs that jointly learns factorized weight uncertainties from initialization. By parameterizing each weight matrix $W\in\mathbb{R}^{m\times n}$ as $W=AB^{\top}$ where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$, we reduce variational parameters from $\mathcal{O}(mn)$ to $\mathcal{O}(r(m+n))$ while capturing structured correlations between weights. We provide both theoretical and empirical contributions:
Theory. We prove that the induced posterior $q(W)$ on weights is singular with respect to Lebesgue measure, concentrating entirely on the measure-zero manifold of rank-$r$ matrices (See Figure 2). This geometric constraint distinguishes our approach from full-rank mean-field methods and captures non-trivial weight correlations: $\text{Cov}(W_{ij},W_{i^{\prime}j^{\prime}})\neq 0$ for weights sharing latent factors. Building on this structure, we establish three theoretical guarantees: (i) deterministic loss approximation bounds via the Eckart-Young-Mirsky (EYM) theorem (Eckart and Young, 1936), with excess risk controlled by tail singular values $\sum_{i>r}\sigma_{i}^{2}$ of the target weight matrix, (ii) probabilistic risk decomposition for learned low-rank factors separating learning error from rank bias, and (iii) Tighter generalization bounds due to the complexity reduction. These results establish that low-rank structure provides not just efficiency but provably tighter theoretical guarantees when weight matrices exhibit fast singular value decay, a property we verify empirically holds for modern architectures (see Section 4 and Appendix H).
Implementation. We implement our method from scratch for three architecture families: multilayer perceptrons (MLPs)(Rumelhart et al., 1986), transformers (Vaswani et al., 2017), and LSTMs (Hochreiter and Schmidhuber, 1997), avoiding black-box libraries to ensure full control over the variational parameterization. Our implementation handles the unique challenges of each architecture: position-wise factorization for transformers, tied weight matrices for LSTMs (Fortunato et al., 2017), and standard feedforward layers for MLPs. Our variational layers serve as drop-in replacements for standard Keras layers, enabling seamless integration into arbitrary architectures.
Experiments. We evaluate our approach on tabular, text, time-series, and image benchmarks (See section 4 an Appendix I). On classification and regression tasks, rank-$r$ factorization often outperforms full-rank mean-field variational inference (Blundell et al., 2015) across Negative log-likelihood, accuracy, and AUROC. When compared to a 5-member Deep Ensemble (Lakshminarayanan et al., 2017), a single rank-$r$ model achieves competitive predictive performance with significantly reduced computational cost. We further demonstrate improved out-of-distribution detection and selective prediction capabilities than MFVI and deterministic-backbone with low-rank perturbations(Dusenberry et al., 2020a) (see Section 4).
Our work establishes that low-rank variational inference is not only a computational convenience but a principled approach with provable benefits, enabling practical Bayesian deep learning across modern architectures.
2 Methods
2.1 Low-Rank Variational Inference
We parameterize each weight matrix $W\in\mathbb{R}^{m\times n}$ via a low-rank factorization $W=AB^{\top}$ where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$. We place independent priors on the factors, $p(A,B)=p_{A}(A)\,p_{B}(B)$, and employ a mean-field variational posterior over the factors, $q(A,B)=q_{A}(A)\,q_{B}(B)$. The weight distribution is induced through the transformation $(A,B)\mapsto AB^{\top}$. This concentrates the posterior support on the rank-$r$ matrices manifold, whose geometric and generalization properties we analyze in Section 3. We employ a scale mixture prior following Blundell et al. (2015): $p_{A}(A)=\prod_{j}\left[\pi\mathcal{N}(a_{j}\mid 0,\sigma_{1}^{2})+(1-\pi)\mathcal{N}(a_{j}\mid 0,\sigma_{2}^{2})\right],$ and similarly for $p_{B}(B)$, where $\mathcal{N}(x\mid\mu,\sigma^{2})$ denotes the Gaussian density with mean $\mu$ and variance $\sigma^{2}$, $j$ indexes the $mr$ elements of $A$ (and $nr$ elements of $B$), and $\sigma_{1}^{2}\gg\sigma_{2}^{2}$ creates a heavy-tailed prior that encourages sparse structure while allowing occasional large weights. The variational posteriors are mean-field Gaussians: $q_{A}(A;\theta_{A})=\prod_{j=1}^{mr}\mathcal{N}(a_{j}\mid\mu_{A,j},\sigma_{A,j}^{2})$ and $q_{B}(B;\theta_{B})=\prod_{j=1}^{nr}\mathcal{N}(b_{j}\mid\mu_{B,j},\sigma_{B,j}^{2})$, where $\theta_{A}=\{\mu_{A},\sigma_{A}\}$ and $\theta_{B}=\{\mu_{B},\sigma_{B}\}$ are the variational parameters learned through backpropagation.
2.2 Evidence Lower Bound
[…]
$\mathcal{L}(q)$ $=\underbrace{\mathbb{E}_{q_{A}q_{B}}[\log p(\mathcal{D}\mid AB^{\top})]}_{\text{data fit}}-\underbrace{\beta(\text{KL}(q_{A}\|p_{A})+\text{KL}(q_{B}\|p_{B}))}_{\text{regularization}}.$ (1)
where $\beta$ is the KL scale (or temperature parameter) that controls the strength of the regularization penalty. The key insight is that the KL divergence decomposes into two independent terms enabling efficient parallel computation. All three terms in Eq. (1) are estimated via Monte Carlo. For the expected log-likelihood, we sample $A\sim q_{A}$ and $B\sim q_{B}$, form $W=AB^{\top}$, and evaluate $\log p(\mathcal{D}\mid W)$. The KL terms are computed similarly via sampling, as the scale mixture prior precludes closed-form computation.
2.3 Gradient Estimation and Optimization
We optimize the ELBO (1) using gradient-based methods with the reparameterization trick (Kingma and Welling, 2014). Following the Bayes by Back-propagation algorithm from Blundell et al. (2015), we parameterize the standard deviations as $\sigma=\log(1+\exp(\rho))$ to ensure positivity. To sample from the posteriors, we draw $\epsilon_{A},\epsilon_{B}\sim\mathcal{N}(0,I)$ and compute: $A=\mu_{A}+\log(1+\exp(\rho_{A}))\circ\epsilon_{A},\quad B=\mu_{B}+\log(1+\exp(\rho_{B}))\circ\epsilon_{B}$ where $\circ$ denotes elementwise multiplication. This makes sampling differentiable with respect to the variational parameters, allowing gradients to flow through Monte Carlo estimates of the ELBO. We use Adam optimizer (Kingma and Ba, 2015) for all models.
2.4 Architecture-Specific Implementations
We apply our factorization to MLPs, LSTMs, and Transformers.
Multilayer Perceptrons. For fully connected layers $W_{\ell}\in\mathbb{R}^{d_{\ell}\times d_{\ell+1}}$, we factorize as $W_{\ell}=A_{\ell}B_{\ell}^{\top}$ where $A_{\ell}\in\mathbb{R}^{d_{\ell}\times r_{\ell}}$ and $B_{\ell}\in\mathbb{R}^{d_{\ell+1}\times r_{\ell}}$. Each layer maintains its own rank $r_{\ell}$, tunable independently or uniformly. This requires no architecture-specific modifications.
Transformers. We factorize query, key, value projections ($\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$) and feed-forward weights ($\mathbf{W}_{1}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{ff}}}$, $\mathbf{W}_{2}\in\mathbb{R}^{d_{\text{ff}}\times d_{\text{model}}}$). Since these are linear operations, we use the same variational layers as MLPs. For embeddings $W_{emb}\in\mathbb{R}^{V\times d}$, we exploit batch sparsity: only rows of $A$ for tokens in the current batch are sampled, reducing cost from $\mathcal{O}(Vd)$ to $\mathcal{O}(|U|r+dr)$ where $|U|$ is unique tokens. Independent factors per head preserve multi-head structure.
LSTMs. We factorize input-to-hidden ($\mathbf{W}_{ih}\in\mathbb{R}^{d_{\text{in}}\times 4h}$) and hidden-to-hidden ($\mathbf{W}_{hh}\in\mathbb{R}^{h\times 4h}$) matrices, where the factor of 4 accounts for input, forget, output, and cell gates. Following Fortunato et al. (2017), factors $A$ and $B$ are sampled once per batch, $W=AB^{\top}$ is cached across time steps, then resampled for the next batch. This ensures KL divergence is computed once per sequence rather than per time step, maintaining the correct variational objective.
3 Theoretical Analysis
We now establish the theoretical foundations of our approach. We show that the factorization $W=AB^{\top}$ induces a different posterior geometry than mean-field methods (Section 3.1), captures structured weight correlations (Section 3.2), provides bounded approximation error under optimal rank selection (Section 3.3), and yields provably tighter generalization bounds (Section 3.4).
3.1 The Induced Posterior: Geometry and Singularity
When we parameterize the posterior over factors $A$ and $B$ rather than directly over $W$, the resulting distribution $q_{W}$ on weight matrices has a geometric structure determined by the factorization constraint. We formalize this through the induced posterior and characterize its support.
Definition 3.1 (Induced Posterior).
Let $A$ and $B$ be drawn from the variational posterior $q(A,B)$. The induced posterior (pushforward measure (Bogachev, 2007)) $q_{W}$ is the distribution of the random variable $W=AB^{\top}$.
Lemma 3.2 (Constrained Support ).
Let $\mathcal{R}_{r}\subset\mathbb{R}^{m\times n}$ be the set of matrices with rank at most $r$. $\mathcal{R}_{r}=\left\{W\in\mathbb{R}^{m\times n}\mid\text{rank}(W)\leq r\right\}$ and let $W=AB^{\top}$ where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$. Then: (i) $\text{rank}(W)\leq r$ and (ii) For any variational posterior $q(A,B)$, $q_{W}$, satisfies: $q_{W}(\mathcal{R}_{r})=1$ (see Appendix B.4 for complete proof)
Lemma 3.3 (Measure Zero of $\mathcal{R}_{r}$).
The set $\mathcal{R}_{r}$ has Lebesgue measure zero in $\mathbb{R}^{m\times n}$ when $r<\min(m,n)$. (see Appendix B.5 for complete proof)
Theorem 3.4 (Singularity of the Induced Posterior).
Let $r<\min(m,n)$ and let $q(A,B)$ be any variational distribution over factors $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$. Then the induced posterior $q_{W}$ is singular with respect to Lebesgue measure on $\mathbb{R}^{m\times n}$.
Proof.
We prove that the induced posterior $q_{W}$ is singular with respect to Lebesgue measure by showing it cannot admit a density function.
(i) From Lemma 3.2, we have already established that: $q_{W}(\mathcal{R}_{r})=1,$ which means the induced posterior $q_{W}$ is entirely supported on the low-rank manifold $\mathcal{R}_{r}$. From Lemma 3.3, we have that the Lebesgue measure of this manifold is zero: $\lambda(\mathcal{R}_{r})=0.$
(ii) Proof by contradiction: Suppose for the sake of contradiction that $q_{W}$ admits a probability density function $f:\mathbb{R}^{m\times n}\rightarrow[0,\infty)$ with respect to the Lebesgue measure $\lambda$. By the definition of a density, the probability of the set $\mathcal{R}_{r}$ would be given by the integral of the density over that set: $q_{W}(\mathcal{R}_{r})=\int_{\mathcal{R}_{r}}f(W)\,d\lambda(W).$ Since the domain of integration has measure zero ($\lambda(\mathcal{R}_{r})=0$), standard integration theory tells us this integral must be zero: $\int_{\mathcal{R}_{r}}f(W)\,d\lambda(W)=0.$ However, this contradicts our finding in step (i) that $q_{W}(\mathcal{R}_{r})=1$. Therefore, no such density function $f$ exists with respect to $\lambda$ on $\mathbb{R}^{m\times n}$. This means $q_{W}$ is singular with respect to $\lambda$. ∎
Interpretation. This result reveals an important geometric distinction: standard mean-field posteriors $q(W)=\prod_{ij}\mathcal{N}(w_{ij}|\mu_{ij},\sigma_{ij}^{2})$ have positive density everywhere in $\mathbb{R}^{m\times n}$, while our induced posterior $q_{W}$ concentrates on the rank-$r$ manifold $\mathcal{R}_{r}$, which has zero volume in the full weight space (See Figure 2). This geometric constraint impacts inductive bias. As argued by Wilson and Izmailov (2020), generalization in probabilistic models depends on two properties: the support of the posterior and the inductive biases it encodes. Mean-field imposes a bias of independence, treating weights as freely adjustable parameters. Our factorization imposes a bias of correlation by restricting support to $\mathcal{R}_{r}$, acting as an implicit regularizer: updating $W_{ij}=\sum_{k=1}^{r}A_{ik}B_{jk}$ requires modifying shared factors affecting entire rows and columns. This prevents local memorization and enables coherent uncertainty propagation.
3.2 Structured Weight Correlations
We now characterize the covariance structure of $W$ and show how the rank parameter $r$ controls the expressiveness of these correlations.
Lemma 3.5 (Covariance of Weight Entries).
For weight entries $W_{ij}=\sum_{k=1}^{r}A_{ik}B_{jk}$ and $W_{i^{\prime}j^{\prime}}=\sum_{\ell=1}^{r}A_{i^{\prime}\ell}B_{j^{\prime}\ell}$, under mean-field factorization $q(A,B)=q_{A}(A)q_{B}(B)$ with independent entries, we have:
$\mathrm{Cov}(W_{ij},W_{i^{\prime}j^{\prime}})=\sum_{k=1}^{r}\mathrm{Cov}(A_{ik}B_{jk},A_{i^{\prime}k}B_{j^{\prime}k}).$ (2)
In general, this is nonzero when weights share common latent factors. See Appendix B.7 for complete proof.
Implications. While $A,B$ are under mean-field factorization, enabling us to leverage existing mean field algorithms, entries of W are not independent. Weights $W_{ij}$ and $W_{i^{\prime}j^{\prime}}$ that share latent factors (indexed by $k$) exhibit correlated uncertainties, capturing global structure in weight space (see Figure 1). This filters high-frequency noise that does not align with the dominant low-rank structure, providing generalization benefits that independent weight posteriors can not capture. The rank $r$ controls this expressiveness: higher rank enables richer correlation patterns while maintaining $O(r(m+n))$ parameters.
Figure 1: Weight correlation structure. Comparison of Full-Rank BBB (left) and Low-Rank Gaussian r=15 (right). Full-Rank exhibits diagonal correlations; Low-Rank captures block structure.
3.3 Loss Approximation Guarantees
We now establish conditions under which the low-rank model closely approximates the full-rank optimum.
Assumptions. We assume the loss function $\ell(\cdot,y)$ satisfies: (i)$L$-Lipschitz: $|\ell(z,y)-\ell(z^{\prime},y)|\leq L\|z-z^{\prime}\|_{2}$ and (ii) Bounded inputs: $\|x\|_{2}\leq R$ almost surely
These conditions hold for common regression and classification losses (see Appendix C.5 for extended analysis)
Let $W^{*}$ be the optimal full-rank weight matrix (learned via back-propagation) and $W^{*}_{r}$ be the optimal rank-$r$ matrix of $W^{*}$ obtained via truncated SVD (Eckart and Young, 1936))
Theorem 3.6 (Bounded Loss Approximation ).
Under the assumptions above, the expected loss difference is bounded:
$|\mathbb{E}[\ell(W^{*}x,y)]-\mathbb{E}[\ell(W^{*}_{r}x,y)]|\leq LR\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})},$ (3)
where $\sigma_{i}(W^{*})$ are the singular values of $W^{*}$ and $\rho=$rank($W^{*}$). In practice, our learned matrix $W=AB^{\top}$ may not achieve the optimal rank-$r$ approximation $W^{*}_{r}$ of $W^{*}$. Thus, we decompose the total approximation error into two components.
Theorem 3.7 (Decomposition of Approximation Error).
Let $W=AB^{\top}$ be optimal learned rank-$r$ matrix and let $\sigma_{>r}^{2}=\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})$. Then for any $(x,y)$,
$|\ell(Wx,y)-\ell(W^{*}x,y)|\leq LR\bigl(\|W-W^{*}_{r}\|_{F}+\sigma^{2}_{>r}\bigr).$ (4)
See Appendix C.1 for complete proofs and the Bayesian extension with $\beta$-smooth losses.
Implications. Theorem 3.6 establishes that the optimal rank-$r$ approximation incurs minimal loss when tail singular values $\sigma_{>r}^{2}$ decay rapidly. Theorem 3.7 reveals that the total error for our learned factorization $W=AB^{\top}$ decomposes into two components: (i) learning error $\|W-W_{r}\|_{F}$, which measures how well training finds the optimal rank-$r$ solution and can be reduced through better optimization, and (ii) rank bias $\sigma_{>r}^{2}$, the unavoidable approximation error from rank restriction. In practice, rank $r$ of each layer can be tuned via ablation studies with reduced computational budget (fewer epochs, fewer MC samples). When pretrained deterministic weights are available, singular value decay analysis can optionally narrow the search range or validate ablation results (See Figures 7, 9, and 15). This connects our theoretical result directly to implementation decisions.
Figure 2: Geometric distinction between mean-field and low-rank posteriors. 3D projection of weight space $\mathbb{R}^{m\times n}$. Left: Rank-$r$ manifold $\mathcal{M}_{r}$ (blue surface, dimension r(m+nr)r(m+n-r)). Middle: Mean-field posterior $q_{\text{MF}}(W)$ has full-dimensional support (volume). Right: Low-rank posterior $q_{\text{LR}}(A,B)$ concentrates on the manifold surface.
3.4 PAC-Bayes Generalization Bounds
Theorem 3.8 (Tighter Bounds for Low-Rank Posteriors).
Consider full-rank mean-field with $D_{\text{full}}=mn$ and low-rank factorization with $D_{\text{LR}}=r(m+n)$ parameters. The ratio of complexity terms in the PAC-Bayes bound (McAllester, 1998) satisfies:
$\frac{\text{Complexity}(Q_{\text{LR}})}{\text{Complexity}(Q_{\text{full}})}\simeq\sqrt{\frac{r(m+n)}{mn}}\simeq\sqrt{r\left(\frac{1}{m}+\frac{1}{n}\right)}\ll 1$ (5)
when $r\ll\min(m,n)$. Figure 3 shows the empirical difference. See Appendix D for complete proof.
3.5 Gaussian Complexity Generalization Bounds
While PAC-Bayes bounds (Section 3.4) capture generalization through KL divergence, we provide a complementary perspective using Gaussian complexity, a data-dependent measure exploiting the geometric structure of rank-constrained weight matrices. Leveraging recent work on vector-valued Gaussian complexity for deterministic low-rank networks (Pinto et al., 2025), we establish guarantees that explicitly depend on spectral norm bounds and rank constraints.
Theorem 3.9 (Gaussian-complexity bound for low-rank BNN predictors).
Consider a depth-$D$ network with weights $W_{i}=A_{i}B_{i}^{\top}$ where $A_{i}\in\mathbb{R}^{h_{i}\times r_{i}}$, $B_{i}\in\mathbb{R}^{h_{i-1}\times r_{i}}$ and $r_{i}\leq\min(h_{i},h_{i-1})$. Let the variational posterior factorize as $q(A,B\mid\theta)=\prod_{i=1}^{D}q_{A}(A_{i}\mid\theta_{i}^{A})q_{B}(B_{i}\mid\theta_{i}^{B})$ and assume $\|A_{i}\|_{2}\leq C_{i}^{A}$, $\|B_{i}\|_{2}\leq C_{i}^{B}$ almost surely. Define $C_{i}:=C_{i}^{A}C_{i}^{B}$ and the posterior-mean predictor $f_{\mathrm{BNN}}(x;\theta)=\mathbb{E}_{(A,B)\sim q}[f_{A,B}(x)]$. Let $\phi$ be $1$-Lipschitz and let $\ell(\cdot,y)\in[0,1]$ be $\mathcal{L}$-Lipschitz in its first argument. Then for $S\sim\mathcal{D}^{N}$, with probability at least $1-\delta$,
$\mathbb{E}\big[\ell(f_{\mathrm{BNN}}(x;\theta),y)\big]\leq\frac{1}{N}\sum_{j=1}^{N}\ell(f_{\mathrm{BNN}}(x_{j};\theta),y_{j})+$ (6)
$\sqrt{\pi\mathcal{L}}\;\hat{\mathcal{G}}_{S}\!\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)+3\sqrt{\frac{\log(2/\delta)}{2N}},$
where $\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)$ is the deterministic class of depth-$D$ networks whose layers satisfy $\|W_{i}\|_{2}\leq C_{i}$ and $\mathrm{rank}(W_{i})\leq r_{i}$. Moreover, $\hat{\mathcal{G}}_{S}(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r))$ admits the explicit bound of (Pinto et al., 2025)(Thm.7). See Appendix E for complete proof .
Interpretation. Although Gaussian complexity is not the canonical Bayesian tool (PAC-Bayes is), it offers insight into capacity control: low-rank layers impose geometric constraints that reduce complexity beyond parameter-count arguments. As shown in Figure 3, the Gaussian complexity bounds are vacuous for both methods and do not provide practical generalization certificates. However, they formalize how rank constraints and spectral norms jointly control model capacity, suggesting that the rank-induced restrictions contribute to the good generalization often observed in low-rank neural networks.
Figure 3: Empirical generalization bounds for low-rank Bayesian neural networks. PAC-Bayes (left) and Gaussian complexity (right) bounds use empirical values from trained LSTM model and training data. PAC-Bayes bound exhibits critical rank $r^{*}\approx 11$ transitioning from non-vacuous to vacuous (>1>1, dashed red line). Gaussian complexity decreases from 45.56 (full-rank) to 18.97 with rank reduction. Full-rank Bayesian models yield vacuous PAC-Bayes bounds regardless of configuration.
4 Experiments
We compare six approaches: (i) Deterministic baseline, (ii) Deep Ensemble (Lakshminarayanan et al., 2017) ($M=5$), (iii) Full-Rank BBB (Blundell et al., 2015), (iv) Low-Rank (random init), (v) Low-Rank SVD init (see Appendix F.9), and (vi) Rank-1 multiplicative (Dusenberry et al., 2020a). Low-Rank methods (iv-v) use factorization $W=AB^{\top}$ with rank $r$. See Appendix G for all implementation details.
4.1 Evaluation Metrics
Predicted probabilities are Monte Carlo averages over posterior weight samples. For classification, we report Accuracy, AUROC (discrimination), NLL and ECE (calibration). For regression, we report MAE and $R^{2}$ (accuracy), PICP and MPIW (calibration). Uncertainty is quantified via mutual information (MI) (Houlsby et al., 2011): MI Ratio (OOD vs in-domain mean MI), AUPR-Err (error detection), and AUPR-Succ (selective prediction). OOD detection uses AUROC-OOD, AUPR-OOD (OOD as positive), and AUPR-In (in-domain as positive). Params denotes trainable parameters.
For each architecture, we select the rank $r$ via ablation studies with reduced computational budget (fewer epochs and Monte Carlo samples), and when deterministic baselines are available, we further validate the chosen ranks using singular value decay analysis of the corresponding weight matrices (See Appendix F.9 and Figures 9, 15, 7).
4.2 MIMIC-III: ICU Mortality Prediction
We first evaluate our method on the MIMIC-III clinical database (Johnson et al., 2016) to predict binary ICU mortality from patient vitals and laboratory values. The dataset comprises 40,40640{,}406 training and 4,4904{,}490 test samples with $44$ clinical features extracted from the first 48 hours of ICU stays ($8.4\%$ mortality rate). For out-of-distribution evaluation, we test on 5,3575{,}357 newborn ICU records ($1.1\%$ mortality), representing significant distributional shift from the adult training population. All methods use a 2-layer MLP with hidden dimension $128$. Low-Rank methods used $r=15$. Complete experimental details are provided in Appendix G.3.
[…]
Table 1: MIMIC-III evaluation metrics averaged over 5 independent runs. Best in bold, second underlined. Arrows indicate optimization direction: $\uparrow$ (higher better), $\downarrow$ (lower better).
[…]
Deterministic .922±.003 .145±.035 .500±.000 .544±.000 .456±.000 .284±.017 22.4k
Deep Ens. .929±.002 .237±.035 .738±.061 .754±.048 .721±.053 .300±.021 112k
Full-Rank .895±.004 .412±.029 .770±.051 .759±.055 .807±.035 .401±.033 44.8k
Low-Rank .895±.001 .540±.027 .802±.018 .788±.034 .824±.013 .433±.020 13.6k
LR-SVD .898±.001 .486±.155 .713±.069 .719±.046 .738±.074 .452±.035 13.6k
Rank-1 .901±.010 .322±.019 .705±.073 .742±.062 .663±.080 .317±.024 23.3k
Figure 4: Model comparison on MIMIC-III (averaged across 5 seeds). Low-Rank Gaussian r=15 (orange) achieves superior OOD detection. Deep Ensemble maintains better calibration and in-domain discrimination. Rank-1 multiplicative (green) achieves better calibration but weaker OOD detection. Full-Rank BBB (blue) shows balanced but moderate performance across metrics.
4.3 Beijing Air Quality: Time Series Forecasting
Figure 5: Selective prediction on Beijing Air Quality LSTM. While Deep Ensemble (red) achieves best point predictions at 100% retention, Bayesian methods outperform when discarding the most uncertain samples. Low-Rank (green) achieves largest improvement (17.4% MAE reduction at 80% retention), demonstrating superior uncertainty quality for selective prediction.
Next we evaluate our approach on the Beijing Multi-Site Air Quality dataset (Chen, 2015) for PM2.5 forecasting from meteorological features. Using 24-hour sliding windows, the dataset has 29,213 training, 6,260 validation, and 6,260 test samples with 15 features. For OOD evaluation, we use 20,050 samples from Guangzhou (Chen, 2016), a city with distinct subtropical climate representing distributional shift from Beijing’s continental climate. All methods use 2-layer LSTMs with 64 hidden units. Low-Rank methods decompose input-to-hidden ($\mathbf{W}_{ih}\in\mathbb{R}^{15\times 256}$) and hidden-to-hidden ($\mathbf{W}_{hh}\in\mathbb{R}^{64\times 256}$) weight matrices as $\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}$ with ranks $r=14$ for the first layer and $20$ for the second. Complete experimental details are provided in Appendix G.4.
Results. Table 2 and Figure 6 show that Low-Rank Bayesian LSTM achieves the best coverage (PICP=0.790) and near-best calibration (ECE=0.114, second only to Full-Rank BBB’s 0.111) while using 64% fewer parameters than Full-Rank BBB. For OOD detection, Low-Rank attains second-best performance (AUROC=0.710, AUPR=0.861) after Deep Ensemble (0.730, 0.883) despite 86% fewer parameters. Figure 5 shows Bayesian methods outperform Deep Ensemble when filtering uncertain samples, with Low-Rank achieving 17.4% error reduction at 80% retention. The calibration gap is notable: Low-Rank achieves PICP near the nominal 95% level, while Deep Ensemble (0.310) and Rank-1 (0.449) severely undercover. This confirms that structured weight correlations from low-rank factorization (Section 3.2) propagate uncertainty more reliably. See Appendix H.3 for extended results and analysis.
Table 2: Beijing Air Quality results averaged over 4 independent runs. Best in bold, second underlined. Arrows indicate optimization direction: $\uparrow$ (higher better), $\downarrow$ (lower better).
[…]
Deterministic 10.79 , , 0.5 0.5 33K
Full-Rank BBB 10.55 0.111 0.788 0.492 0.743 132K
Low-Rank (r=14,20r{=}14,20) 10.63 0.114 0.790 0.710 0.861 47K
LR-SVD 10.70 0.139 0.773 0.687 0.812 47K
Rank-1 Mult. 10.80 0.307 0.449 0.580 0.751 66K
Deep Ensemble (5) 10.45 0.317 0.310 0.730 0.883 330K
Figure 6: Model comparison on Beijing Air Quality LSTM (averaged across 5 seeds). Deep Ensemble (brown) achieves best OOD detection but poor coverage. Low-Rank BBB (green) shows best coverage (PICP) and near-best calibration (ECE) with second-best OOD performance. Low-Rank SVD (red) shows strong OOD detection with good coverage. Full-Rank BBB (orange) achieves best calibration but moderate coverage. Rank-1 BBB (purple) shows poor uncertainty performance.
4.4 Text Classification with Transformers
Finally, we evaluate our approach on binary sentiment classification using the Stanford Sentiment Treebank (SST-2) dataset (Socher et al., 2013), which contains 67,349 training samples and 872 validation samples of movie reviews. For out-of-distribution detection, we use the AG News dataset (Zhang et al., 2015) with 7,600 samples from the news domain, which provides a meaningful distributional shift from the sentiment classification task. We implement a 4-layer Transformer encoder with $d_{\text{model}}=256$, 4 attention heads, and feed-forward dimension $d_{\text{ff}}=512$, using the BERT tokenizer with vocabulary size 30,522. The architecture was inspired by BERT-mini (Turc et al., 2019). Singular value decay analysis (Figure 7) show that the embedding layer, which represents 70% of the parameter counts for all models, has particularly rapid decay. Applying rank $r=16$ for all layers yields 86% parameter reduction vs. deterministic and 93% vs. full-rank BBB.
Results. Table 3 and Figure 8 show Low-Rank BBB achieves 0.806 accuracy with best AUPR-In (0.302) and second-best AUROC-OOD (0.640), while Low-Rank SVD attains highest AUPR-Succ (0.927) among Bayesian methods. Deep Ensemble achieves best overall accuracy (0.825) and calibration (NLL=0.434) with higher MI Ratio (1.55 vs 1.35), while Full-Rank BBB underperforms across metrics (0.752 accuracy, 0.552 NLL), consistent with (Cinquin et al., 2021) findings. Low-Rank BBB uses $13\times$ fewer parameters than Full-Rank (1.5M vs 19.8M) and $33\times$ fewer parameters than Deep Ensemble (49.6M). Training times reflect this efficiency: Low-Rank trains in 8.2 minutes versus 23.1 for Full-Rank BBB and 64.7 for Deep Ensemble. See Appendix H.4 for extended results and analysis.
Figure 7: Singular value decay for embedding and value projection matrices showing rapid decay.
Table 3: SST-2 evaluation metrics averaged over 4 independent runs. Best in bold, second underlined. Arrows indicate optimization direction: $\uparrow$ (higher better), $\downarrow$ (lower better).
[…]
Deterministic .812±.006 .490±.003 .500±.000 .000±.000 .102±.000 .812±.006 9.9M
Deep Ens. .825±.006 .434±.023 .657±.009 1.55±0.05 .267±.016 .933±.003 49.6M
Full-Rank .752±.011 .552±.010 .622±.019 1.31±0.07 .222±.028 .888±.006 19.8M
Low-Rank .806±.003 .527±.006 .640±.028 1.35±0.10 .302±.027 .917±.005 1.5M
LR-SVD .800±.005 .516±.002 .616±.011 1.22±0.10 .272±.012 .923±.003 1.5M
Rank-1 .795±.013 .507±.005 .613±.015 1.11±0.09 .272±.027 .931±.004 10.0M
Figure 8: Model comparison on SST-2 Transformer comparison (averaged across 4 seeds). Deep Ensemble (green) achieves best overall performance. Low-Rank BBB (red) shows good OOD detection and highest epistemic uncertainty discrimination (MI Ratio). Full-Rank BBB (orange) demonstrates moderate balanced performance. Rank-1 BBB (brown) maintains good calibration but weaker OOD metrics.
4.5 Experimental Insights
Calibration–OOD Detection Tradeoff. We observe a consistent tradeoff: low-rank achieves competitive OOD detection but exhibits higher NLL and ECE than Deep Ensembles. We hypothesize that the rank constraint maintains broader uncertainty distributions that better distinguish OOD inputs at the cost of predictive sharpness. This tradeoff favors low-rank in safety-critical settings where reliable uncertainty quantification matters more than marginal likelihood gains.
Computational Efficiency. Parameter reduction on small or medium scales architecture provides substantial memory savings, enabling deployment in resource-constrained settings, but minimal training speedups (Table 9). However, efficiency gains emerge at scale on the larger Transformer architecture: our low-rank approaches train in 7.95 minutes (SVD init) and 8.22 minutes (random init) versus 23.1 for Full-Rank BBB and 64.7 for Deep Ensembles (Table 12).
Hyperparameter Sensitivity. The KL weight $\beta$ and rank $r$ require tuning. We recommend starting with $\beta=1/N_{\text{train}}$ or $\beta=1/N_{\text{batches}}$ depending on layer implementation, then adjusting based on validation performance. Higher $\beta$ may improve OOD detection but degrade in-distribution metrics; we use KL annealing for stability. Rank $r$ impacts the calibration-OOD tradeoff (see Figures 10, 13, 14) and can be tuned via ablation and singular value analysis when pretrained weights are available.
SVD Initialization. SVD warm-starting yields modest gains on specific metrics (MIMIC-III AUROC: 0.898 vs 0.895; SST-2 AUPR-Succ: 0.927, highest among Bayesian methods) but shows no consistent benefit across tasks. On Beijing LSTMs, random initialization performs better overall. SVD initialization is only worthwhile when pretrained weights already exist, as the improvements do not warrant training an additional deterministic model. See Appendix H.1 for detailed analysis.
5 Discussion and Conclusion
We presented a low-rank variational inference framework for Bayesian neural networks that parameterizes weights as $W=AB^{\top}$ with learned posteriors on the factors. This induces a posterior singular with respect to Lebesgue measure, concentrating on the rank-$r$ manifold, a geometric property distinguishing our approach from mean-field VI, low-rank covariance approximations, and perturbation-based methods. The factorization reduces parameters from $O(mn)$ to $O(r(m+n))$ while capturing structured weight correlations through shared latent factors.
Our theoretical analysis established tighter PAC-Bayes bounds scaling with rank, loss decomposition via Eckart-Young-Mirsky, and complexity transfer from deterministic to Bayesian predictors. Empirically, the method achieved competitive predictive performance across MLPs, LSTMs, and Transformers.
On MIMIC-III, low-rank attained superior OOD detection; on SST-2, it achieved second-best OOD performance after Deep Ensembles while substantially outperforming Full-Rank BBB and Rank-1 baselines. On Beijing Air Quality (LSTM), low-rank achieved best prediction interval coverage and second-best OOD detection after Deep Ensembles. We observed a consistent calibration–OOD tradeoff: the rank constraint maintains broader epistemic uncertainty at the cost of predictive sharpness, favoring applications where reliable uncertainty matters more than marginal likelihood gains.
Several limitations exist. The calibration gap relative to ensembles may limit applicability in likelihood-focused settings, though ensembling multiple low-rank models offers potential mitigation at lower cost than standard ensembles. Architecture choices reflect our computational constraints, not method limitations. In fact, Computational benefits emerge primarily at larger scales where GPU parallelism can be exploited.
Rank and KL weight require tuning, and while improvements are directionally consistent, some confidence intervals overlap across seeds. A broader experimental scope including comparisons with SWAG and low-rank covariance methods would strengthen the positioning of this work.
Our method is agnostic to the distributional family of the posterior. While we employ Gaussian posteriors for tractability, the factorization $W=AB^{\top}$ naturally extends to other location-scale distributions such as the Laplace distribution, which provides heavier tails that may improve uncertainty quantification. Additional experiments using a Laplace posterior are presented in Appendix I.
Future work includes low-rank ensembles, adaptive rank selection, and scaling to larger language models. We are exploring spike-and-slab priors that enforce sparsity while permitting large weights where needed however an in-depth analysis of prior choice and its effect on performance across different data regimes and tasks would provide valuable guidance for practitioners. The framework also extends naturally to computer vision, including medical imaging and mineral prospectivity mapping. The singular posterior geometry provides both theoretical foundations and practical benefits for uncertainty quantification, establishing low-rank factorization as a principled path toward scalable Bayesian deep learning.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Comparison with Related Low-Rank Approaches
Table 4: Comparison with related low-rank approaches for BNNs. Our method is the only approach that (i) learns both backbone and uncertainty end-to-end without pretrained models, (ii) reduces both mean and variance parameterization, and (iii) induces a geometrically singular posterior.
Method Backbone Parameter count Posterior Weight
training (mean + var) support correlations
Mean-field VI (Blundell et al., 2015) Learned Bayes. $2mn$ Full $\mathbb{R}^{m\times n}$ Independent
Low-rank cov. (Swiatkowski et al., 2020; Tomczak et al., 2020) Learned Bayes. $mn+r(m+n)$ Full $\mathbb{R}^{m\times n}$ Via covariance
Rank-1 mult. (Dusenberry et al., 2020a) Learned det. + pert. $mn+2(m+n)$ Full $\mathbb{R}^{m\times n}$ Rank-1 structure
Bella (Doan et al., 2025) Frozen + adapt $r(m+n)$ Affine subspace Via $\Delta\theta_{i}$
LoRA + Bayes (Yang et al., 2024) Frozen + adapt $r(m+n)$ Affine subspace Via $\Delta W$
Ours Learned Bayes. $r(m+n)$ Rank-$r$ manifold $\mathcal{R}_{r}$ Via shared factors
A.1 Mean-Field Variational Inference (Blundell et al., 2015)
Mean-field variational inference places independent Gaussian posteriors over each weight: $q(W)=\prod_{ij}\mathcal{N}(w_{ij}|\mu_{ij},\sigma_{ij}^{2})$. This requires $\mathcal{O}(2mn)$ variational parameters per layer (mean and variance for each of $mn$ weights) and imposes an independence assumption: weight uncertainties are decoupled.
Our method differs in several respects. We place distributions directly over the factor matrices $A$ and $B$, optimizing $\mathcal{L}(q)=\mathbb{E}_{q_{A}q_{B}}[\log p(\mathcal{D}|AB^{\top})]-\mathrm{KL}(q_{A}\|p_{A})-\mathrm{KL}(q_{B}\|p_{B})$ with only $\mathcal{O}(r(m+n))$ variational parameters. The induced posterior $q_{W}$ concentrates on the rank-$r$ manifold $\mathcal{R}_{r}$ (Theorem 3.4, Lemma 3.2), which has measure zero in $\mathbb{R}^{m\times n}$. In contrast, mean-field has full-dimensional support.
Additionally, our factorization induces non-trivial covariances between weights sharing latent factors (Lemma 3.5), capturing structured dependencies. Mean-field assumes complete independence. The rank constraint prevents overfitting to individual examples by forcing shared factor modifications rather than allowing arbitrary independent weight adjustments (Section 3.1).
A.2 Low-Rank Covariance Approximations (Swiatkowski et al., 2020), (Tomczak et al., 2020)
Low-rank covariance methods maintain full mean parameters $\mu_{W}\in\mathbb{R}^{m\times n}$ but approximate the posterior covariance via low-rank factors: $\mathrm{Cov}(W)\approx LL^{\top}$ where $L\in\mathbb{R}^{mn\times r}$. This requires $\mathcal{O}(mn+r(m+n))$ parameters.
Our approach differs in parameterization and support. These methods learn the full mean $\mu_{W}$ independently and use low-rank factors solely to approximate off-diagonal covariances. We jointly parameterize both mean and uncertainty through the same factorization $W=AB^{\top}$. Their posterior retains full-dimensional support in $\mathbb{R}^{m\times n}$ with only the covariance structure constrained, whereas our support is restricted to $\mathcal{R}_{r}$, a lower-dimensional manifold.
The parameter count difference is substantial: they scale as $mn+r(m+n)$ due to storing the full mean vector, while we achieve $r(m+n)$ by encoding both mean and covariance structure in the factorization. This translates to reduced memory requirements and more efficient gradient computation.
A.3 Rank-1 Multiplicative Perturbations (Dusenberry et al., 2020a)
Rank-1 Bayesian neural networks learn a deterministic weight matrix $W_{0}$ and scale it by multiplicative rank-1 factors: $W=(1+rs^{\top})\odot W_{0}$, where $r\in\mathbb{R}^{m}$ and $s\in\mathbb{R}^{n}$ are learned stochastically. The posterior is placed over these perturbation vectors, requiring $\mathcal{O}(2(m+n))$ additional variational parameters for means and variances on $r,s$.
Several distinctions emerge. First, rank-1 methods learn end-to-end on a deterministic backbone, whereas we place uncertainty on both factors from initialization without a fixed mean weight matrix. Second, they parameterize $q(r,s)$ with $W$ emerging through element-wise multiplicative perturbations, while we directly parameterize $q_{A}(A)q_{B}(B)$ with $W=AB^{\top}$ emerging through matrix multiplication.
The induced posteriors exhibit different uncertainty structures: theirs has multiplicative structure (perturbations scale deviations from $W_{0}$), while ours has additive low-rank structure (sums of rank-1 components). This affects uncertainty propagation. Furthermore, rank-1 methods are restricted to scaling-based perturbations, whereas our factorization permits arbitrary rank-$r$ matrices when $r>1$, enabling richer expressiveness (Lemma 3.5).
A.4 Bayesian LoRA via SVGD (Doan et al., 2025)
Bella adapts a frozen pretrained backbone via low-rank LoRA factors and applies Stein Variational Gradient Descent (SVGD) over an ensemble of rank-$r$ particles. Each particle $i$ has weights $\theta_{i}=\theta_{0}+B_{i}A_{i}^{\top}$, where $\theta_{0}$ is frozen and only $(A_{i},B_{i})$ are updated. This requires $\mathcal{O}(r(m+n))$ parameters per particle, totaling $n_{\text{particles}}\times r(m+n)$.
Thi approach differs from ours in training scheme and posterior representation. Bella is post-hoc, requiring a pretrained backbone, whereas ours learns end-to-end from random initialization. Bella uses SVGD (kernel-based repulsion over discrete particles) to approximate the posterior over adapters; we employ mean-field variational inference over factors. SVGD maintains $n_{\text{particles}}$ distinct solutions; variational inference maintains a continuous distribution.
Regarding parameterization, Bella freezes $\theta_{0}$ and optimizes $(A_{i},B_{i})$ for each particle independently. We learn variational distributions $q_{A}(A),q_{B}(B)$ over shared factors, with individual samples drawn implicitly. Bella’s posterior is an empirical distribution over $n_{\text{particles}}$ affine subspaces around $\theta_{0}$, while ours concentrates on a single rank-$r$ manifold $\mathcal{R}_{r}$ in the full weight space.
Computationally, Bella requires $n_{\text{particles}}$ forward and backward passes; we require one per minibatch, leveraging reparameterization for efficient gradient estimation. Additionally, Bella depends on a high-quality pretrained model, whereas we work from random initialization, enabling application to novel architectures without pretrained checkpoints.
A.5 Post-hoc Laplace Approximation on LoRA (Yang et al., 2024)
This approach freezes a pretrained backbone, finetunes LoRA adapters to a maximum a posteriori (MAP) estimate, then applies the Laplace approximation (K-FAC-based Hessian curvature) around that point in the LoRA subspace. Computing and storing the Hessian approximation requires $\mathcal{O}(r(m+n))$ parameters with additional structural overhead.
The training schemes differ substantially: Laplace-LoRA is post-hoc (requiring pretrained backbone, finetuning, then Hessian computation), whereas ours is end-to-end from initialization. The posterior types also differ: Laplace assumes a local quadratic approximation around the MAP point, while our mean-field variational posterior does not rely on local curvature assumptions.
In terms of expressiveness, the Laplace approximation is accurate near the MAP but can be poor in regions of high curvature. Our factorization captures global structure via the low-rank constraint (Lemma 3.5). Regarding scalability, Hessian computation in high dimensions is computationally expensive even with K-FAC approximations; we avoid this entirely through reparameterization.
A.6 Summary
Our method is distinguished by combining (i) end-to-end variational learning with (ii) parameter-efficient factorization and (iii) geometrically-constrained posterior support on a rank-$r$ manifold. The rank-$r$ manifold constraint (Theorem 3.4) provides theoretical advantages in generalization bounds (Section 3.4) and empirical benefits in out-of-distribution detection and calibration, as demonstrated in our experiments. Table 4 summarizes the key difference across the approaches we discussed.
Appendix B Background and Proofs
B.1 Variational Inference for Bayesian Neural Networks
[…]
B.2 Low-Rank Matrix Factorization
Theorem B.1 (Existence of Low-Rank Factorization).
Let $W\in\mathbb{R}^{m\times n}$ be a matrix of rank $r\leq\min(m,n)$ Then there exist matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$ such that $W=AB^{T}$(Ford, 2014). The EYM theorem further establishes that for any full-rank matrix with singular value decomposition $W_{\text{full}}=\sum_{i=1}^{\ell}\sigma_{i}u_{i}v_{i}^{\top}$, where $\sigma_{i}$ are the singular values (in decreasing order) $u_{i}\in\mathbb{R}^{m}$ are the left singular vectors, and $v_{i}\in\mathbb{R}^{n}$ are the right singular vectors, the truncated SVD $W_{r}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{\top}$ provides the optimal rank-$r$ approximation in Frobenius norm. (See  B.3 for complete proof.) with approximation error $\|W_{\text{full}}-W_{r}\|_{F}^{2}=\sum_{i=r+1}^{\ell}\sigma_{i}^{2}$.
B.3 Proof of Theorem B.1: Existence of Low-Rank Factorization
Proof.
We prove that for any matrix $W\in\mathbb{R}^{m\times n}$ with $\mathrm{rank}(W)\leq r$, there exist matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$ such that $W=AB^{\top}$. The proof proceeds by considering two cases: (i) when $\mathrm{rank}(W)=r$ exactly, and (ii) when $\mathrm{rank}(W)=k<r$ is strictly less than $r$.
Case 1: $\mathrm{rank}(W)=r$.
Assume $\mathrm{rank}(W)=r$. Then the columns of $W$ span an $r$-dimensional subspace of $\mathbb{R}^{m}$. Let $A=[a_{1},\ldots,a_{r}]$ be the matrix whose columns form a basis for this column space, so $A\in\mathbb{R}^{m\times r}$.
For every column $w_{j}$ of $W$ (for $j=1,\ldots,n$), there exists a set of coefficients $(b_{j,i})_{i=1}^{r}$ such that $w_{j}$ is a linear combination of the basis vectors $a_{1},\ldots,a_{r}$:
$w_{j}=\sum_{i=1}^{r}b_{j,i}a_{i},\quad\text{where}\quad a_{i}=\begin{bmatrix}a_{1i}\\ a_{2i}\\ \vdots\\ a_{mi}\end{bmatrix}$.
Now we define $B\in\mathbb{R}^{n\times r}$ with rows $B_{j}=[b_{j,1},\ldots,b_{j,r}]$:
$B=\begin{bmatrix}b_{1,1}&\cdots&b_{1,r}\\ \vdots&\ddots&\vdots\\ b_{n,1}&\cdots&b_{n,r}\end{bmatrix}$.
For each column $w_{j}$, we have $w_{j}=AB_{j}^{\top}$:
$w_{j}=\begin{bmatrix}a_{1,1}&\cdots&a_{1,r}\\ \vdots&\ddots&\vdots\\ a_{m,1}&\cdots&a_{m,r}\end{bmatrix}\begin{bmatrix}b_{j,1}\\ \vdots\\ b_{j,r}\end{bmatrix}=\begin{bmatrix}\sum_{i=1}^{r}a_{1,i}b_{j,i}\\ \vdots\\ \sum_{i=1}^{r}a_{m,i}b_{j,i}\end{bmatrix}$.
Thus $w_{j}=AB_{j}^{\top}$ for each column $j=1,\ldots,n$, which yields $W=AB^{\top}$.
Case 2: $\mathrm{rank}(W)=k<r$.
Now assume $\mathrm{rank}(W)=k<r$. Then by Case 1, there exist matrices $A_{k}\in\mathbb{R}^{m\times k}$ and $B_{k}\in\mathbb{R}^{n\times k}$ such that $W=A_{k}B_{k}^{\top}$.
Define $A_{r}=[A_{k}\mid O_{m\times(r-k)}]$ by augmenting $A_{k}$ with rkr-k zero columns, and similarly define $B_{r}=[B_{k}\mid O_{n\times(r-k)}]$ by augmenting $B_{k}$ with rkr-k zero columns. Then:
$A_{r}B_{r}^{\top}$ $=\bigl[\,A_{k}\mid O_{m\times(r-k)}\,\bigr]\begin{bmatrix}B_{k}^{\top}\\ O_{(r-k)\times n}\end{bmatrix}$
$=A_{k}B_{k}^{\top}+O_{m\times n}$
$=A_{k}B_{k}^{\top}$
$=W.$
Therefore $W=A_{r}B_{r}^{\top}$ where $A_{r}\in\mathbb{R}^{m\times r}$ and $B_{r}\in\mathbb{R}^{n\times r}$.
Conclusion.
We have shown that for any $W$ with $\mathrm{rank}(W)\leq r$, there exist $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$ such that $W=AB^{\top}$, completing the proof. ∎
B.4 Proof of Lemma 3.2: Constrained Support
Proof.
Let $q(A,B)$ be any variational distribution over $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$. We show that the induced posterior $q_{W}$ places all its probability mass on the set of rank-at-most-$r$ matrices.
By Theorem B.1, any matrix of the form $W=AB^{\top}$ with $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$ satisfies $\mathrm{rank}(W)\leq r$.
The distribution of the induced posterior, $q_{W}$, is the image measure (or pushforward measure) of $q(A,B)$ under the deterministic function:
$f(A,B)=AB^{\top}$
defined as:
$f:\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\longrightarrow\mathbb{R}^{m\times n},\qquad(A,B)\longmapsto AB^{\top}.$
Thus every sample $(A,B)\sim q(A,B)$ is mapped to $W=AB^{\top}$, which by construction satisfies $\mathrm{rank}(W)\leq r$.
Therefore, $q_{W}$ places all its probability mass on the set
$\mathcal{R}_{r}=\{W\in\mathbb{R}^{m\times n}:\mathrm{rank}(W)\leq r\},$
that is, $q_{W}(\mathcal{R}_{r})=1$. ∎
B.5 Proof of Lemma 3.3: Measure Zero of Low-Rank Manifold
Proof.
We prove that the set of matrices with rank at most $r$ has Lebesgue measure zero in $\mathbb{R}^{m\times n}$.
Recall from the characterization of matrix rank (see, e.g., (Ford, 2014)) that for any $W\in\mathcal{R}_{r}$ (i.e., $\mathrm{rank}(W)\leq r$), we have:
$\det(W_{I,J})=0\quad\forall\ I\subset\{1,\ldots,m\}\ \text{and}\ J\subset\{1,\ldots,n\}$
such that $|I|=|J|=r+1$. That is, all $(r+1)\times(r+1)$ minors must vanish.
Let us take a given pair of indices $(I_{0},J_{0})$ of size $r+1$. Define the function $g:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ as
$g(W)=\det(W_{I_{0},J_{0}}).$
Since the determinant is a sum of products of entries, $g(W)$ is a polynomial in the entries of $W$.
Claim: The polynomial $g(W)$ is not identically zero.
Proof of claim: There exists at least one matrix $W\in\mathbb{R}^{m\times n}$ such that $g(W)\neq 0$. To see this, construct $W^{*}$ such that the submatrix $W^{*}_{I_{0},J_{0}}=I_{r+1}$ (the identity matrix of size $r+1$) and all other entries are zero. Then:
$g(W^{*})=\det(I_{r+1})=1\neq 0.$
Thus $g$ is not identically zero.
We now apply the following classical result from real analysis (see (Folland, 1999)):
Theorem B.2 (Zero Sets of Polynomials).
If $p:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a polynomial that is not identically zero, then the set of its zeros $\{x\in\mathbb{R}^{d}:p(x)=0\}$ has Lebesgue measure zero.
Applying this theorem to our function $g$, the set
$Z_{I_{0},J_{0}}=\{W\in\mathbb{R}^{m\times n}\mid g(W)=0\}$
has Lebesgue measure zero:
$\lambda(Z_{I_{0},J_{0}})=0.$
Now, since a matrix has rank at most $r$ if and only if all its $(r+1)\times(r+1)$ minors vanish, we can write:
$\mathcal{R}_{r}=\bigcap_{\begin{subarray}{c}I,J:\\ |I|=|J|=r+1\end{subarray}}Z_{I,J}.$
[…]
$\lambda(\mathcal{R}_{r})=0.$
This completes the proof. ∎
B.6 Additional Remarks on Theorem 3.4: Singularity
The singularity of $q_{W}$ with respect to Lebesgue measure has several important implications beyond what we discussed in the main text.
Relationship to Measure-Theoretic Concepts: The induced posterior $q_{W}$ is mutually singular with respect to Lebesgue measure, meaning it places all its mass on a set ($\mathcal{R}_{r}$) that has zero Lebesgue measure. When $q(A,B)$ is chosen to be continuous (e.g., factorized Gaussians), $q_{W}$ is an example of a singular continuous measure i.e it has no atoms (point masses) yet cannot be represented by a density with respect to Lebesgue measure. This is distinct from absolutely continuous measures (which have densities) and discrete measures (which have atoms).
Comparison to Mean-Field. Mean-field Gaussian posteriors are absolutely continuous with respect to Lebesgue measure, meaning they admit a density function $f:\mathbb{R}^{m\times n}\rightarrow[0,\infty)$. This difference in measure-theoretic character explains why the two approaches have such different inductive biases: mean-field spreads mass across the entire space with positive density everywhere, while our induced posterior concentrates entirely on the low-rank manifold.
Practical Implications. Despite being singular with respect to Lebesgue measure, $q_{W}$ is perfectly well-defined and tractable. Samples can be drawn efficiently by sampling $(A,B)\sim q(A,B)$ and computing $W=AB^{\top}$. Expectations with respect to $q_{W}$ can be computed via Monte Carlo sampling. The singularity is a geometric property that constrains the support to $\mathcal{R}_{r}$, not a computational obstacle. In fact, this concentration of support is precisely what provides the beneficial inductive bias discussed in Section 3.1.
B.7 Proof of Lemma 3.5: Covariance Structure
We derive the covariance structure between weight entries induced by the low-rank factorization $W=AB^{\top}$.
B.7.1 Setup
Consider two weight entries $W_{ij}$ and $W_{i^{\prime}j^{\prime}}$ where
$W_{ij}=\sum_{k=1}^{r}A_{ik}B_{jk},\qquad W_{i^{\prime}j^{\prime}}=\sum_{\ell=1}^{r}A_{i^{\prime}\ell}B_{j^{\prime}\ell}.$
We seek to compute $\mathrm{Cov}(W_{ij},W_{i^{\prime}j^{\prime}})$ under a factorized variational posterior:
$q(A,B)=\left(\prod_{i,k}q(A_{ik})\right)\left(\prod_{j,k}q(B_{jk})\right),$
which implies $A\perp B$ and independence between distinct entries within each factor.
B.7.2 Auxiliary Lemma: Covariance of Products
Lemma B.3 (Covariance of Independent Products).
Let $A$ and $B$ be independent random variables, and let $X,U$ be measurable functions of $A$ and $Y,V$ be measurable functions of $B$. Then
$\mathrm{Cov}(XY,UV)=\mathbb{E}[XU]\mathbb{E}[YV]-\mathbb{E}[X]\mathbb{E}[U]\mathbb{E}[Y]\mathbb{E}[V].$
Proof.
By definition of covariance:
$\mathrm{Cov}(XY,UV)=\mathbb{E}[XYUV]-\mathbb{E}[XY]\mathbb{E}[UV].$
Since $X,U$ depend only on $A$ and $Y,V$ depend only on $B$, and $A\perp B$, the expectations factorize:
$\mathbb{E}[XYUV]=\mathbb{E}[XU]\mathbb{E}[YV].$
Similarly:
$\mathbb{E}[XY]=\mathbb{E}[X]\mathbb{E}[Y],\qquad\mathbb{E}[UV]=\mathbb{E}[U]\mathbb{E}[V].$
Substituting:
$\mathrm{Cov}(XY,UV)=\mathbb{E}[XU]\mathbb{E}[YV]-\mathbb{E}[X]\mathbb{E}[Y]\mathbb{E}[U]\mathbb{E}[V].$
B.7.3 Derivation of Weight Covariance
Define $Z_{k}=A_{ik}B_{jk}$ and $Z^{\prime}_{\ell}=A_{i^{\prime}\ell}B_{j^{\prime}\ell}$. Then:
$\mathrm{Cov}(W_{ij},W_{i^{\prime}j^{\prime}})=\mathrm{Cov}\left(\sum_{k=1}^{r}Z_{k},\sum_{\ell=1}^{r}Z^{\prime}_{\ell}\right).$
By bilinearity of covariance:
$\mathrm{Cov}(W_{ij},W_{i^{\prime}j^{\prime}})=\sum_{k=1}^{r}\sum_{\ell=1}^{r}\mathrm{Cov}(Z_{k},Z^{\prime}_{\ell}).$
Case 1: $k\neq\ell$.
By independence of distinct entries and Lemma B.3:
$\mathrm{Cov}(A_{ik}B_{jk},A_{i^{\prime}\ell}B_{j^{\prime}\ell})$ $=\mathbb{E}[A_{ik}A_{i^{\prime}\ell}]\mathbb{E}[B_{jk}B_{j^{\prime}\ell}]-\mathbb{E}[A_{ik}]\mathbb{E}[A_{i^{\prime}\ell}]\mathbb{E}[B_{jk}]\mathbb{E}[B_{j^{\prime}\ell}]$
$=\mathbb{E}[A_{ik}]\mathbb{E}[A_{i^{\prime}\ell}]\mathbb{E}[B_{jk}]\mathbb{E}[B_{j^{\prime}\ell}]-\mathbb{E}[A_{ik}]\mathbb{E}[A_{i^{\prime}\ell}]\mathbb{E}[B_{jk}]\mathbb{E}[B_{j^{\prime}\ell}]$
$=0.$
Case 2: $k=\ell$.
For the diagonal terms:
$\mathrm{Cov}(A_{ik}B_{jk},A_{i^{\prime}k}B_{j^{\prime}k})$ $=\mathbb{E}[A_{ik}A_{i^{\prime}k}]\mathbb{E}[B_{jk}B_{j^{\prime}k}]-\mathbb{E}[A_{ik}]\mathbb{E}[A_{i^{\prime}k}]\mathbb{E}[B_{jk}]\mathbb{E}[B_{j^{\prime}k}].$
Therefore:
$\mathrm{Cov}(W_{ij},W_{i^{\prime}j^{\prime}})=\sum_{k=1}^{r}\left(\mathbb{E}[A_{ik}A_{i^{\prime}k}]\mathbb{E}[B_{jk}B_{j^{\prime}k}]-\mathbb{E}[A_{ik}]\mathbb{E}[A_{i^{\prime}k}]\mathbb{E}[B_{jk}]\mathbb{E}[B_{j^{\prime}k}]\right).$
B.7.4 Interpretation
This result establishes that weight entries $W_{ij}$ and $W_{i^{\prime}j^{\prime}}$ are correlated whenever the corresponding factor entries share statistical dependencies. Specifically:
  • Row correlation: If $i=i^{\prime}$ (same row), weights in that row are correlated through shared $A$ factors.
  • Column correlation: If $j=j^{\prime}$ (same column), weights in that column are correlated through shared $B$ factors.
  • Rank-modulated coupling: The correlation strength depends on all $r$ latent dimensions, allowing the model to capture complex interdependencies.
Contrast with Mean-Field. A full-rank mean-field posterior imposes $\mathrm{Cov}(W_{ij},W_{i^{\prime}j^{\prime}})=0$ for $(i,j)\neq(i^{\prime},j^{\prime})$, forcing all weights to be independent. The low-rank factorization naturally induces structured correlations, enabling the posterior to capture coherent variations across many weights simultaneously. This structured dependency is crucial when the true posterior exhibits correlated directions, as it provides a richer variational family than fully factorized approximations (see Figure 1).
Appendix C Proofs from Theoretical Analysis
C.1 Loss Approximation Preliminaries
C.1.1 Preliminaries
Definition C.1 (Expected Loss).
The expected loss (risk) over the data distribution $\mathcal{D}$ is given by
$\mathcal{L}(W)=\mathbb{E}_{(x,y)\sim\mathcal{D}}\big[\ell(Wx,y)\big].$
Definition C.2 ($\beta$-Smoothness).
A function $\ell(z,y)$ is $\beta$-smooth (with respect to $z$) if, for all $z,z^{\prime}\in\mathbb{R}^{d}$,
$\bigl\|\nabla_{z}\ell(z,y)-\nabla_{z}\ell(z^{\prime},y)\bigr\|_{2}\leq\beta\|z-z^{\prime}\|_{2}.$
This implies the standard quadratic upper bound (descent lemma): for all $z,z^{\prime}\in\mathbb{R}^{d}$,
$\ell(z^{\prime},y)\leq\ell(z,y)+\langle\nabla_{z}\ell(z,y),\,z^{\prime}-z\rangle+\frac{\beta}{2}\|z^{\prime}-z\|_{2}^{2}.$
C.1.2 Eckart-Young-Mirsky Theorem
Theorem C.3 (Eckart-Young-Mirsky Theorem (Eckart and Young, 1936)).
Let $W\in\mathbb{R}^{m\times n}$ have singular value decomposition
$W=\sum_{i=1}^{\rho}\sigma_{i}u_{i}v_{i}^{\top},$
where $\rho=\mathrm{rank}(W)$, $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{\rho}>0$ are the singular values, and $\{u_{i}\}$, $\{v_{i}\}$ are the corresponding left and right singular vectors. For any $r<\rho$, define the rank-$r$ truncation
$W_{r}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{\top}.$
Then $W_{r}$ is the best rank-$r$ approximation to $W$ in both the Frobenius norm and the spectral (operator) norm:
$W_{r}$ $=\arg\min_{\begin{subarray}{c}M\in\mathbb{R}^{m\times n}\\ \mathrm{rank}(M)\leq r\end{subarray}}\|W-M\|_{F},$ (7)
$W_{r}$ $=\arg\min_{\begin{subarray}{c}M\in\mathbb{R}^{m\times n}\\ \mathrm{rank}(M)\leq r\end{subarray}}\|W-M\|_{2}.$ (8)
Moreover, the approximation errors are given exactly by the tail singular values:
$\|W-W_{r}\|_{F}$ $=\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}},$ (9)
$\|W-W_{r}\|_{2}$ $=\sigma_{r+1}.$ (10)
Proof sketch.
The proof relies on the orthogonal invariance of both norms and the fact that singular vectors form orthonormal bases.
For the Frobenius norm: Since $\{u_{i}v_{i}^{\top}\}$ are orthonormal under the Frobenius inner product $\langle A,B\rangle_{F}=\mathrm{tr}(A^{\top}B)$, we have
$\|W-W_{r}\|_{F}^{2}=\left\|\sum_{i=r+1}^{\rho}\sigma_{i}u_{i}v_{i}^{\top}\right\|_{F}^{2}=\sum_{i=r+1}^{\rho}\sigma_{i}^{2},$
by orthonormality. To show optimality, suppose $M$ has rank at most $r$. The dimension of the space orthogonal to $M$’s range has dimension at least mrm-r, which must intersect the (nr)(n-r)-dimensional space spanned by $\{u_{r+1},\ldots,u_{\rho}\}$. This intersection implies $\|W-M\|_{F}\geq\sigma_{r+1}$, with equality achieved by $W_{r}$.
For the spectral norm: Since $\|W-W_{r}\|_{2}=\max_{\|x\|=1}\|(W-W_{r})x\|$, and $W-W_{r}=\sum_{i=r+1}^{\rho}\sigma_{i}u_{i}v_{i}^{\top}$, the maximum is achieved when $x=v_{r+1}$, giving $\|W-W_{r}\|_{2}=\sigma_{r+1}$.
For complete proofs, see (Eckart and Young, 1936; Stewart and Sun, 1990). ∎
C.1.3 Frobenius-Euclidean Consistency
Lemma C.4 (Consistency of Frobenius and Euclidean Norms).
For any matrix $A\in\mathbb{R}^{m\times n}$ and vector $x\in\mathbb{R}^{n}$,
$\|Ax\|_{2}\leq\|A\|_{F}\|x\|_{2}.$
Proof.
Let $A=[a_{1}\mid a_{2}\mid\cdots\mid a_{n}]$ where $a_{j}\in\mathbb{R}^{m}$ are the columns of $A$. Then
$Ax=\sum_{j=1}^{n}x_{j}a_{j}.$
By the triangle inequality and Cauchy-Schwarz inequality:
$\|Ax\|_{2}$ $=\left\|\sum_{j=1}^{n}x_{j}a_{j}\right\|_{2}$
$\leq\sum_{j=1}^{n}|x_{j}|\|a_{j}\|_{2}\quad\text{(triangle inequality)}$
$\leq\left(\sum_{j=1}^{n}|x_{j}|^{2}\right)^{1/2}\left(\sum_{j=1}^{n}\|a_{j}\|_{2}^{2}\right)^{1/2}\quad\text{(Cauchy-Schwarz)}$
$=\|x\|_{2}\sqrt{\sum_{j=1}^{n}\sum_{i=1}^{m}a_{ij}^{2}}$
$=\|x\|_{2}\|A\|_{F}.$
Here we used the definition of the Frobenius norm:
$\|A\|_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}a_{ij}^{2}}=\sqrt{\sum_{j=1}^{n}\|a_{j}\|_{2}^{2}}.$
This establishes the desired inequality. See Chapters 4 and 5 of (Ford, 2014) for further discussion of matrix norm relationships. ∎
C.2 Proof of Theorem 3.6: Optimal Rank-$r$ Approximation
Theorem C.5 (Restating Theorem 3.6).
Let $W^{*}$ be the optimal full-rank weight matrix and $W^{*}_{r}$ be the optimal rank-$r$ matrix (via truncated SVD). Under $L$-Lipschitz loss and bounded inputs $\|x\|_{2}\leq R$ a.s.,
$|\mathbb{E}[\ell(W^{*}x,y)]-\mathbb{E}[\ell(W^{*}_{r}x,y)]|\leq LR\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}.$
Proof.
First, since $W^{*}$ is the global minimizer of the loss $\mathcal{L}$, we know that for any other matrix, including $W^{*}_{r}$, the excess risk is non-negative:
$\mathcal{L}(W^{*}_{r})-\mathcal{L}(W^{*})\geq 0.$
The same holds for the pointwise loss:
$\ell(W^{*}_{r}x,y)-\ell(W^{*}x,y)\geq 0.$
By definition:
$W^{*}=\underset{W\in\mathbb{R}^{m\times n}}{\mathrm{argmin}}\ \mathcal{L}(W),\quad W^{*}_{r}=\sum_{i=1}^{r}\sigma_{i}u_{i}^{*}{v_{i}^{*}}^{\top}.$
From the Lipschitz property of $\ell$:
$\ell(W^{*}_{r}x,y)-\ell(W^{*}x,y)\leq L\|(W^{*}_{r}-W^{*})x\|_{2}.$
Using the consistency of the Frobenius norm with the Euclidean norm (Lemma C.4):
$\|(W^{*}_{r}-W^{*})x\|_{2}\leq\|x\|_{2}\|W^{*}_{r}-W^{*}\|_{F}.$
Thus:
$\ell(W^{*}_{r}x,y)-\ell(W^{*}x,y)\leq L\|x\|_{2}\underbrace{\|W^{*}_{r}-W^{*}\|_{F}}_{\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}}}.$
By the Eckart-Young-Mirsky theorem (Theorem C.3), since $W^{*}_{r}$ is the SVD truncation of $W^{*}$, the Frobenius norm of the difference is exactly determined by the tail singular values:
$\|W^{*}_{r}-W^{*}\|_{F}=\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}.$
Taking the expectation over the data distribution $\mathcal{D}$ to move from pointwise loss $\ell$ to risk $\mathcal{L}$:
$\mathbb{E}_{(x,y)\sim\mathcal{D}}\Bigl[\ell(W^{*}_{r}x,y)-\ell(W^{*}x,y)\Bigr]$ $\leq\mathbb{E}_{(x,y)\sim\mathcal{D}}\Biggl[L\|x\|_{2}\,\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}\Biggr].$
Since $\|x\|_{2}\leq R$ almost surely by assumption, and the singular values are constants with respect to the expectation:
$\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\ell(W^{*}_{r}x,y)-\ell(W^{*}x,y)\right]\leq LR\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}.$
By definition of the risk $\mathcal{L}(W)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(Wx,y)]$:
$\mathcal{L}(W^{*}_{r})-\mathcal{L}(W^{*})\leq LR\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}.$
Since $\mathcal{L}(W^{*}_{r})-\mathcal{L}(W^{*})\geq 0$ (as $W^{*}$ is the minimizer), we have established the bound. This completes the proof. ∎
C.3 Proof of Theorem 3.7: Learned Approximation
Theorem C.6 (Restating Theorem 3.7).
Let $W^{*}$ be the optimal full-rank weight matrix and $W^{*}_{r}$ be the optimal rank-$r$ matrix (via truncated SVD) and let $W=AB^{\top}$ be the optimal learned rank-$r$ matrix. Then for any $(x,y)$,
$\bigl|\ell(Wx,y)-\ell(W^{*}x,y)\bigr|\leq LR\left(\|W-W^{*}_{r}\|_{F}+\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}\right).$
Proof.
Add and subtract $\ell(W^{*}_{r}x,y)$ and apply the triangle inequality:
$\bigl|\ell(Wx,y)-\ell(W^{*}x,y)\bigr|$ $=\Bigl|\ell(Wx,y)-\ell(W^{*}_{r}x,y)+\ell(W^{*}_{r}x,y)-\ell(W^{*}x,y)\Bigr|$
$\leq\bigl|\ell(Wx,y)-\ell(W^{*}_{r}x,y)\bigr|+\bigl|\ell(W^{*}_{r}x,y)-\ell(W^{*}x,y)\bigr|.$
First term (learning error):
By $L$-Lipschitz continuity:
$\bigl|\ell(Wx,y)-\ell(W^{*}_{r}x,y)\bigr|\leq L\|(W-W^{*}_{r})x\|_{2}\leq L\|x\|_{2}\|W-W^{*}_{r}\|_{F}\leq LR\|W-W^{*}_{r}\|_{F},$
where we used Lemma C.4 and $\|x\|_{2}\leq R$ a.s.
Second term (rank bias):
From Theorem 3.6 (or equivalently, by Lipschitz continuity and the Eckart-Young-Mirsky theorem):
$\bigl|\ell(W^{*}_{r}x,y)-\ell(W^{*}x,y)\bigr|\leq L\|x\|_{2}\|W^{*}_{r}-W^{*}\|_{F}=L\|x\|_{2}\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}\leq LR\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}.$
Combining both terms:
$\bigl|\ell(Wx,y)-\ell(W^{*}x,y)\bigr|\leq LR\left(\|W-W^{*}_{r}\|_{F}+\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}\right).$
This completes the proof. ∎
Interpretation. This decomposition reveals two sources of error:
  1. 1.
    Learning error: $\|W-W^{*}_{r}\|_{F}$ measures how far the learned $W=AB^{\top}$ is from the optimal rank-$r$ approximation. This depends on optimization quality and can be reduced with better training procedures.
  2. 2.
    Rank bias (approximation error): $\sqrt{\sum_{i=r+1}^{\rho}\sigma_{i}^{2}(W^{*})}$ is the unavoidable error from restricting to rank $r$. If $W^{*}$ has rapidly decaying singular values, this term is small.
C.4 Extension to Stochastic Weights (Bayesian Setting)
Lemma C.7 (Variance and Loss, Lemma 1.2.3 in (Nesterov, 2018)).
Assume $\ell(\cdot,y)$ is differentiable and $\beta$-smooth in its first argument. Let $Z$ be an $\mathbb{R}^{d}$-valued random variable with $\mathbb{E}\|Z\|_{2}^{2}<\infty$ and set $\mu:=\mathbb{E}[Z]$. Then
$\Bigl|\mathbb{E}\bigl[\ell(Z,y)\bigr]-\ell(\mathbb{E}[Z],y)\Bigr|\leq\frac{\beta}{2}\,\mathbb{E}\bigl[\|Z-\mathbb{E}[Z]\|_{2}^{2}\bigr].$
Proof.
Since $\ell$ is $\beta$-smooth, by the descent lemma (Definition C.2):
$\Bigl|\ell(Z,y)-\ell(\mathbb{E}[Z],y)-\bigl\langle\nabla_{z}\ell\bigl(\mathbb{E}[Z],y\bigr),\,Z-\mathbb{E}[Z]\bigr\rangle\Bigr|\leq\frac{\beta}{2}\,\|Z-\mathbb{E}[Z]\|_{2}^{2}.$
This implies:
$\frac{\beta}{2}\,\|Z-\mathbb{E}[Z]\|_{2}^{2}\leq\ell(Z,y)-\ell(\mathbb{E}[Z],y)-\bigl\langle\nabla_{z}\ell\bigl(\mathbb{E}[Z],y\bigr),\,Z-\mathbb{E}[Z]\bigr\rangle\leq\frac{\beta}{2}\,\|Z-\mathbb{E}[Z]\|_{2}^{2}.$
Taking expectations (using linearity) and noting that $\mathbb{E}[Z-\mathbb{E}[Z]]=0$:
$\frac{\beta}{2}\,\mathbb{E}[\|Z-\mathbb{E}[Z]\|_{2}^{2}]\leq\mathbb{E}\bigl[\ell(Z,y)\bigr]-\ell(\mathbb{E}[Z],y)\leq\frac{\beta}{2}\,\mathbb{E}[\|Z-\mathbb{E}[Z]\|_{2}^{2}].$
Therefore:
$\Bigl|\mathbb{E}\bigl[\ell(Z,y)\bigr]-\ell(\mathbb{E}[Z],y)\Bigr|\leq\frac{\beta}{2}\,\mathbb{E}[\|Z-\mathbb{E}[Z]\|_{2}^{2}].$
Theorem C.8 (Expected Loss Difference with Random Weights).
Let $W_{\mathrm{full}}\sim q_{\mathrm{full}}$ and $AB^{\top}\sim q_{\mathrm{lr}}$ be random weight matrices with means $\mu_{f}=\mathbb{E}[W_{\mathrm{full}}]$ and $\mu_{r}=\mathbb{E}[AB^{\top}]$. Assume $\ell$ is $L$-Lipschitz and $\beta$-smooth. Define
$\Delta:=\Bigl|\mathbb{E}\bigl[\ell(W_{\mathrm{full}}x,y)\bigr]-\mathbb{E}\bigl[\ell(AB^{\top}x,y)\bigr]\Bigr|.$
Then
$\Delta\leq L\|x\|_{2}\|\mu_{f}-\mu_{r}\|_{F}+\frac{\beta}{2}\left(\mathrm{Var}(W_{\mathrm{full}}x)+\mathrm{Var}(AB^{\top}x)\right).$
Proof.
Let $z_{f}=W_{\mathrm{full}}x$ and $z_{r}=AB^{\top}x$. We decompose:
$\Delta$ $=\Bigl|\mathbb{E}\bigl[\ell(z_{f},y)\bigr]-\mathbb{E}\bigl[\ell(z_{r},y)\bigr]\Bigr|$
$=\Bigl|\mathbb{E}\bigl[\ell(z_{f},y)\bigr]-\ell(\mathbb{E}[z_{f}],y)+\ell(\mathbb{E}[z_{f}],y)-\ell(\mathbb{E}[z_{r}],y)+\ell(\mathbb{E}[z_{r}],y)-\mathbb{E}\bigl[\ell(z_{r},y)\bigr]\Bigr|$
$\leq\Bigl|\mathbb{E}\bigl[\ell(z_{f},y)\bigr]-\ell(\mathbb{E}[z_{f}],y)\Bigr|+\Bigl|\ell(\mathbb{E}[z_{f}],y)-\ell(\mathbb{E}[z_{r}],y)\Bigr|+\Bigl|\ell(\mathbb{E}[z_{r}],y)-\mathbb{E}\bigl[\ell(z_{r},y)\bigr]\Bigr|.$
Term 1 and Term 3: By Lemma C.7:
$\Bigl|\mathbb{E}\bigl[\ell(z_{f},y)\bigr]-\ell(\mathbb{E}[z_{f}],y)\Bigr|$ $\leq\frac{\beta}{2}\,\mathbb{E}[\|z_{f}-\mathbb{E}[z_{f}]\|_{2}^{2}]=\frac{\beta}{2}\,\mathrm{Var}(W_{\mathrm{full}}x),$
$\Bigl|\ell(\mathbb{E}[z_{r}],y)-\mathbb{E}\bigl[\ell(z_{r},y)\bigr]\Bigr|$ $\leq\frac{\beta}{2}\,\mathrm{Var}(AB^{\top}x).$
Term 2: By $L$-Lipschitz continuity and submultiplicativity:
$\bigl|\ell(\mathbb{E}[z_{f}],y)-\ell(\mathbb{E}[z_{r}],y)\bigr|\leq L\,\|\mathbb{E}[z_{f}]-\mathbb{E}[z_{r}]\|_{2}\leq L\,\|x\|_{2}\,\|\mu_{f}-\mu_{r}\|_{F}.$
Combining all three terms:
$\Delta\leq L\|x\|_{2}\,\|\mu_{f}-\mu_{r}\|_{F}+\frac{\beta}{2}\Bigl(\mathrm{Var}(W_{\mathrm{full}}x)+\mathrm{Var}(AB^{\top}x)\Bigr).$
Remark. This variance-based bound shows that posterior uncertainty contributes an additional $O(\beta\cdot\mathrm{Var})$ term beyond the deterministic approximation error. However, for practical rank selection, Theorems 3.6 and 3.7 provide more directly actionable guidance.
C.5 Analysis of Common Loss Functions
We verify that standard loss functions used in practice satisfy the assumptions of our loss approximation theorems (Theorems 3.6 and 3.7). The key requirement is that $\ell(\cdot,y)$ is $L$-Lipschitz in its first argument.
C.5.1 Classification Losses
The following standard classification losses are globally Lipschitz:
Cross-Entropy with Softmax.
For multi-class classification with $C$ classes, let $z\in\mathbb{R}^{C}$ be the logits and $y\in\{1,\ldots,C\}$ be the true class. The cross-entropy loss is
$\ell(z,y)=-z_{y}+\log\left(\sum_{k=1}^{C}e^{z_{k}}\right).$
Proposition C.9.
The cross-entropy loss with softmax is $\sqrt{2}$-Lipschitz.
Proof.
The gradient is $\nabla_{z}\ell(z,y)=\sigma(z)-e_{y}$, where $\sigma(z)$ is the softmax and $e_{y}$ is the one-hot encoding. Since $\|\sigma(z)-e_{y}\|_{2}\leq\sqrt{2}$ (both are probability vectors, so the maximum distance is $\sqrt{2}$), we have $\|\nabla_{z}\ell\|_{2}\leq\sqrt{2}$, establishing the Lipschitz constant. ∎
Binary Cross-Entropy.
For binary classification with sigmoid activation $\sigma(z)=1/(1+e^{-z})$:
$\ell(z,y)=-y\log(\sigma(z))-(1-y)\log(1-\sigma(z)).$
The gradient is $\nabla_{z}\ell=\sigma(z)-y\in[-1,1]$, so the loss is $1$-Lipschitz.
Hinge Loss.
For binary classification with $y\in\{-1,+1\}$:
$\ell(z,y)=\max(0,1-yz).$
The gradient is $\nabla_{z}\ell\in\{0,-y\}$, giving $L=1$.
C.5.2 Regression Losses
Mean Absolute Error (MAE).
$\ell(z,y)=\|z-y\|_{1}.$
This is $1$-Lipschitz: $|\ell(z,y)-\ell(z^{\prime},y)|\leq\|z-z^{\prime}\|_{1}\leq\sqrt{d}\|z-z^{\prime}\|_{2}$, where $d$ is the output dimension.
Mean Squared Error (MSE).
$\ell(z,y)=\frac{1}{2}\|z-y\|_{2}^{2}.$
Important: MSE is not globally Lipschitz. The gradient is $\nabla_{z}\ell=z-y$, which is unbounded as $\|z-y\|\to\infty$. However, under the assumption that outputs are bounded (i.e., $\|z\|_{2},\|y\|_{2}\leq B$ for some constant $B$), MSE becomes $(2B)$-Lipschitz on the restricted domain.
Remark C.10 (Bounded Outputs for MSE).
In practice, neural network outputs are often implicitly bounded:
  • Activation functions like tanh, sigmoid impose explicit bounds
  • Numerical stability requires finite outputs
  • Training with gradient clipping effectively bounds the domain
Under such conditions, MSE satisfies our Lipschitz assumption with $L=2B$ where $B$ is the output bound. For unbounded regression, MAE or Huber loss should be used instead.
C.5.3 Smoothness Properties
For the Bayesian extension (Theorem C.8), we also require $\beta$-smoothness.
MSE is 2-smooth.
The Hessian is $\nabla^{2}_{z}\ell=I$, so $\|\nabla^{2}_{z}\ell\|_{2}=1$, giving $\beta=2$ (accounting for the factor of $\frac{1}{2}$ in the definition).
Cross-Entropy is smooth.
The Hessian is the covariance matrix of the softmax distribution, which has eigenvalues in $[0,1/4]$, giving $\beta=O(1)$.
C.5.4 Interpretation of Lipschitz and Smoothness
Remark C.11 (Lipschitz Continuity).
A loss $\ell(\hat{y},y)$ being $L$-Lipschitz (with respect to $\hat{y}$) means that small changes in predictions lead to small changes in loss:
$|\ell(\hat{y},y)-\ell(\hat{y}^{\prime},y)|\leq L\|\hat{y}-\hat{y}^{\prime}\|.$
When differentiable, this is equivalent to bounded gradients:
[…]
Intuition: Lipschitz continuity prevents the loss from having arbitrarily steep slopes. Non-Lipschitz losses (like unbounded MSE) can produce very large gradients when predictions are far from targets, potentially causing optimization instability.
Remark C.12 ($\beta$-Smoothness).
A loss being $\beta$-smooth means the gradient itself cannot change too rapidly:
$\|\nabla_{\hat{y}}\ell(\hat{y},y)-\nabla_{\hat{y}}\ell(\hat{y}^{\prime},y)\|\leq\beta\|\hat{y}-\hat{y}^{\prime}\|.$
This is equivalent to bounded curvature: $\|\nabla^{2}_{\hat{y}}\ell\|_{2}\leq\beta$.
Consequence: Smoothness ensures the loss can be upper-bounded by a quadratic approximation:
$\ell(\hat{y}^{\prime},y)\leq\ell(\hat{y},y)+\langle\nabla_{\hat{y}}\ell(\hat{y},y),\hat{y}^{\prime}-\hat{y}\rangle+\frac{\beta}{2}\|\hat{y}^{\prime}-\hat{y}\|^{2}.$
This is the descent lemma, which is essential for optimization: taking gradient steps with learning rate $\eta\leq 1/\beta$ guarantees sufficient descent without overshooting.
Remark C.13 (Lipschitz vs. Smoothness).
These are complementary properties:
  • Lipschitz: Gradient magnitude is bounded $\|\nabla\ell\|\leq L$
  • Smooth: Gradient variation is bounded $\|\nabla\ell-\nabla\ell^{\prime}\|\leq\beta\|x-x^{\prime}\|$
A function can be Lipschitz but not smooth (e.g., $|x|$ has bounded gradient but discontinuous derivative), or smooth but not Lipschitz (e.g., $x^{2}$ has unbounded gradient but bounded curvature on compact sets).
C.5.5 Summary Table
Table 5: Lipschitz and smoothness constants for common loss functions.
Loss Function $L$ (Lipschitz) $\beta$ (Smooth) Notes
MAE 1 N/A Not differentiable
MSE $2B$ 2 Requires bounded outputs $\|z\|,\|y\|\leq B$
Cross-Entropy $\sqrt{2}$ $O(1)$ Global for softmax
Binary Cross-Entropy 1 $O(1)$ Global for sigmoid
Hinge Loss 1 N/A Not differentiable everywhere
Appendix D PAC-Bayes Generalization Bounds
D.1 Background: PAC-Bayes Theory
The PAC-Bayes framework provides generalization bounds that depend on the complexity of the posterior distribution, measured by KL divergence from the prior.
Definition D.1 (Risk and Empirical Risk).
Let $\mathcal{D}$ be a distribution over $\mathcal{X}\times\mathcal{Y}$, and let $\ell(\theta,x,y)\in[0,1]$ be a bounded loss function. For a posterior $Q$ over parameters $\Theta$:
True risk (generalization error):
$L(Q)=\mathbb{E}_{\theta\sim Q}\left[\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(\theta,x,y)]\right].$
Empirical risk (training error):
$\hat{L}(Q)=\mathbb{E}_{\theta\sim Q}\left[\frac{1}{N}\sum_{i=1}^{N}\ell(\theta,x_{i},y_{i})\right].$
The PAC-Bayes bound relates true risk to empirical risk plus a complexity penalty.
Theorem D.2 (McAllester’s Inequality (McAllester, 1998)).
Let $P$ be a fixed prior distribution over $\Theta$. For any $\delta\in(0,1)$, with probability at least $1-\delta$ over the random draw of training set $S\sim\mathcal{D}^{N}$,
$L(Q)\leq\hat{L}(Q)+\sqrt{\frac{\mathrm{KL}(Q\|P)+\ln\left(\frac{2\sqrt{N}}{\delta}\right)}{2N}}.$
This bound decomposes generalization error into:
  • Data fit: $\hat{L}(Q)$ measures training performance
  • Complexity penalty: $\mathrm{KL}(Q\|P)$ measures deviation from prior
D.2 KL Divergence for Factorized Distributions
Lemma D.3 (KL for Factorized Distributions).
Assume fully factorized posterior and prior:
$Q(\theta)=\prod_{i=1}^{D}q_{i}(\theta_{i}),\qquad P(\theta)=\prod_{i=1}^{D}p_{i}(\theta_{i}).$
Define $C_{\max}=\max_{i}\mathrm{KL}(q_{i}\|p_{i})$. Then
$\mathrm{KL}(Q\|P)\leq C_{\max}\cdot D.$
Proof.
By factorization:
$\mathrm{KL}(Q\|P)$ $=\mathbb{E}_{\theta\sim Q}\left[\ln\frac{Q(\theta)}{P(\theta)}\right]$
$=\mathbb{E}_{\theta\sim Q}\left[\ln\frac{\prod_{i=1}^{D}q_{i}(\theta_{i})}{\prod_{i=1}^{D}p_{i}(\theta_{i})}\right]$
$=\mathbb{E}_{\theta\sim Q}\left[\sum_{i=1}^{D}\ln\frac{q_{i}(\theta_{i})}{p_{i}(\theta_{i})}\right]$
$=\sum_{i=1}^{D}\mathbb{E}_{\theta\sim Q}\left[\ln\frac{q_{i}(\theta_{i})}{p_{i}(\theta_{i})}\right].$
For the $i$-th term, by independence:
$\mathbb{E}_{\theta\sim Q}\left[\ln\frac{q_{i}(\theta_{i})}{p_{i}(\theta_{i})}\right]$ $=\int\cdots\int\ln\frac{q_{i}(\theta_{i})}{p_{i}(\theta_{i})}\prod_{j=1}^{D}q_{j}(\theta_{j})\,d\theta_{1}\cdots d\theta_{D}$
$=\int\ln\frac{q_{i}(\theta_{i})}{p_{i}(\theta_{i})}q_{i}(\theta_{i})\,d\theta_{i}\cdot\prod_{j\neq i}\underbrace{\int q_{j}(\theta_{j})\,d\theta_{j}}_{=1}$
$=\mathrm{KL}(q_{i}\|p_{i}).$
Therefore:
$\mathrm{KL}(Q\|P)=\sum_{i=1}^{D}\mathrm{KL}(q_{i}\|p_{i})\leq\sum_{i=1}^{D}C_{\max}=C_{\max}\cdot D.$
D.3 Proof of Theorem 3.8
Theorem D.4 (Restating Theorem 3.8).
Assume bounded loss $\ell(\theta,x,y)\in[0,1]$ and factorized priors/posteriors.
Full-rank: $W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, $D_{\text{full}}=d_{\text{in}}d_{\text{out}}$, with $C_{\max}^{\text{full}}=\max_{i,j}\mathrm{KL}(q_{W_{ij}}\|p_{W_{ij}})$.
Low-rank: $W=AB^{\top}$ with $A\in\mathbb{R}^{d_{\text{in}}\times r}$, $B\in\mathbb{R}^{d_{\text{out}}\times r}$, $D_{\text{LR}}=r(d_{\text{in}}+d_{\text{out}})$, with $C_{\max}^{\text{LR}}=\max\{C_{\max}^{A},C_{\max}^{B}\}$ where $C_{\max}^{A}=\max_{i,k}\mathrm{KL}(q_{A_{ik}}\|p_{A_{ik}})$ and $C_{\max}^{B}=\max_{j,k}\mathrm{KL}(q_{B_{jk}}\|p_{B_{jk}})$.
Then with probability at least $1-\delta$ over $S\sim\mathcal{D}^{N}$:
$L(Q_{\text{full}})$ $\leq\hat{L}(Q_{\text{full}})+\sqrt{\frac{C_{\max}^{\text{full}}\cdot d_{\text{in}}d_{\text{out}}+\ln\left(\frac{2\sqrt{N}}{\delta}\right)}{2N}},$
$L(Q_{\text{LR}})$ $\leq\hat{L}(Q_{\text{LR}})+\sqrt{\frac{C_{\max}^{\text{LR}}\cdot r(d_{\text{in}}+d_{\text{out}})+\ln\left(\frac{2\sqrt{N}}{\delta}\right)}{2N}}.$
When $C_{\max}^{\text{full}}=C_{\max}^{\text{LR}}=C_{\max}$ (e.g., when using identical prior and posterior families for all parameters), the complexity ratio satisfies:
$\frac{\text{Complexity}(Q_{\text{LR}})}{\text{Complexity}(Q_{\text{full}})}=\sqrt{r\left(\frac{1}{d_{\text{in}}}+\frac{1}{d_{\text{out}}}\right)}<1\quad\text{when }r\ll\min(d_{\text{in}},d_{\text{out}}).$
Proof.
Step 1: Apply McAllester’s bound.
For the full-rank posterior $Q_{\text{full}}$, Theorem D.2 gives:
$L(Q_{\text{full}})\leq\hat{L}(Q_{\text{full}})+\sqrt{\frac{\mathrm{KL}(Q_{\text{full}}\|P_{\text{full}})+\ln\left(\frac{2\sqrt{N}}{\delta}\right)}{2N}}.$
Similarly for low-rank:
$L(Q_{\text{LR}})\leq\hat{L}(Q_{\text{LR}})+\sqrt{\frac{\mathrm{KL}(Q_{\text{LR}}\|P_{\text{LR}})+\ln\left(\frac{2\sqrt{N}}{\delta}\right)}{2N}}.$
Step 2: Bound KL divergences using Lemma D.3.
For full-rank with $D_{\text{full}}=d_{\text{in}}d_{\text{out}}$ parameters:
$\mathrm{KL}(Q_{\text{full}}\|P_{\text{full}})\leq C_{\max}^{\text{full}}\cdot d_{\text{in}}d_{\text{out}}.$
For low-rank, we have:
$\mathrm{KL}(Q_{\text{LR}}\|P_{\text{LR}})$ $=\mathrm{KL}(q_{A}(A)\|p_{A}(A))+\mathrm{KL}(q_{B}(B)\|p_{B}(B))$
$\leq C_{\max}^{A}\cdot d_{\text{in}}\cdot r+C_{\max}^{B}\cdot d_{\text{out}}\cdot r$
$\leq C_{\max}^{\text{LR}}\cdot(d_{\text{in}}\cdot r+d_{\text{out}}\cdot r)$
$=C_{\max}^{\text{LR}}\cdot r(d_{\text{in}}+d_{\text{out}}).$
Step 3: Compute complexity ratio (under equal $C_{\max}$ assumption).
When $C_{\max}^{\text{full}}=C_{\max}^{\text{LR}}=C_{\max}$, ignoring the common $\ln\left(\frac{2\sqrt{N}}{\delta}\right)$ term:
$\frac{\text{Complexity}(Q_{\text{LR}})}{\text{Complexity}(Q_{\text{full}})}$ $=\frac{\sqrt{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})}}{\sqrt{C_{\max}\cdot d_{\text{in}}d_{\text{out}}}}$
$=\sqrt{\frac{r(d_{\text{in}}+d_{\text{out}})}{d_{\text{in}}d_{\text{out}}}}$
$=\sqrt{r\left(\frac{1}{d_{\text{in}}}+\frac{1}{d_{\text{out}}}\right)}.$
When $r\ll\min(d_{\text{in}},d_{\text{out}})$, this ratio is much less than 1. ∎
Remark D.5 (General Case with Different $C_{\max}$ Values).
In general, when $C_{\max}^{\text{full}}\neq C_{\max}^{\text{LR}}$, the low-rank approach provides a tighter bound if:
$C_{\max}^{\text{LR}}\cdot r(d_{\text{in}}+d_{\text{out}})<C_{\max}^{\text{full}}\cdot d_{\text{in}}d_{\text{out}}.$
This condition is typically satisfied when using the same variational family (e.g., Gaussians with comparable variances) for both parameterizations, since the per-parameter KL divergences are then of similar magnitude.
Remark D.6 (Precise KL for Low-Rank Factors).
For the sake of complete rigor, with $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$:
$\mathrm{KL}(q_{A}(A)\|p_{A}(A))$ $\leq C_{\max}^{A}\cdot m\cdot r,$ (11)
$\mathrm{KL}(q_{B}(B)\|p_{B}(B))$ $\leq C_{\max}^{B}\cdot n\cdot r.$ (12)
Therefore:
$\mathrm{KL}(Q_{\text{LR}}\|P_{\text{LR}})\leq C_{\max}^{A}\cdot m\cdot r+C_{\max}^{B}\cdot n\cdot r.$ (13)
When $C_{\max}^{A}=C_{\max}^{B}=C_{\max}^{\text{LR}}$:
$\mathrm{KL}(Q_{\text{LR}}\|P_{\text{LR}})\leq C_{\max}^{\text{LR}}\cdot r(n+m).$ (14)
D.4 Interpretation
Remark D.7 (Generalization Benefits of Low-Rank).
Even if the empirical risk $\hat{L}(Q_{\text{LR}})$ is slightly higher than $\hat{L}(Q_{\text{full}})$ (due to rank restriction), the low-rank posterior may still yield a tighter generalization bound because its complexity penalty scales as $r(d_{\text{in}}+d_{\text{out}})$ rather than $d_{\text{in}}d_{\text{out}}$.
This advantage is most pronounced in data-limited regimes where $N\ll d_{\text{in}}d_{\text{out}}$, making the complexity term a dominant contributor to the upper bound on true risk. The parameter reduction from $O(d_{\text{in}}d_{\text{out}})$ to $O(r(d_{\text{in}}+d_{\text{out}}))$ directly translates to better generalization guarantees.
Appendix E Gaussian Complexity Transfer: Deterministic to Bayesian
E.1 Motivation and Positioning
This section addresses a specific conceptual gap: Can deterministic rank-sensitive Gaussian complexity bounds be extended to Bayesian predictive means?
The problem.
Recent work by Pinto et al. (2025) has established sharp generalization bounds for deep networks with low-rank weight constraints using vector-valued Gaussian complexity. These bounds scale with the rank $r$ rather than the ambient dimension, providing dramatically tighter guarantees when $r$ is small. However, these results apply to deterministic predictors with fixed rank-$r$ weights.
In Bayesian neural networks, the deployed predictor is the posterior predictive mean $f_{\text{BNN}}(x)=\mathbb{E}_{W\sim q}[f_{W}(x)]$. Even when every sample $W\sim q$ has rank at most $r$, the posterior mean does not preserve this rank structure at the parameter level. Therefore, deterministic rank-based bounds do not immediately apply.
The solution.
We show that the posterior predictive mean belongs to the closed convex hull of the class of rank-$r$ networks, and that vector-valued Gaussian complexity is invariant under convex hull and closure operations. This allows deterministic rank-sensitive bounds to transfer without degradation to Bayesian predictors.
Relationship to PAC-Bayes.
This result complements PAC-Bayes analysis. PAC-Bayes provides posterior-dependent, information-theoretic bounds through KL divergence and, as we saw in Section D, the PAC-Bayes bound is effectively rank-aware with the complexity scaling as $r(d_{\text{in}}+d_{\text{out}})$ rather than $d_{\text{in}}d_{\text{out}}$. we believe our contribution is architectural: it enables rank-sensitive deterministic bounds to apply to Bayesian predictive averaging, despite rank destruction under expectation.
E.2 Preliminaries
E.2.1 Low-rank Parameterization
For $i=1,\ldots,D$, each layer weight is parameterized as
$W_{i}=A_{i}B_{i}^{\top},\qquad A_{i}\in\mathbb{R}^{h_{i}\times r_{i}},\;\;B_{i}\in\mathbb{R}^{h_{i-1}\times r_{i}},\qquad r_{i}\leq\min(h_{i},h_{i-1}).$
Hence $\mathrm{rank}(W_{i})\leq r_{i}$ by construction.
E.2.2 Variational Posterior
We consider a factorized variational posterior over factors:
$q(A,B\mid\theta)=\prod_{i=1}^{D}q_{A}(A_{i}\mid\theta_{i}^{A})\,q_{B}(B_{i}\mid\theta_{i}^{B}),\qquad A=(A_{1},\ldots,A_{D}),\;\;B=(B_{1},\ldots,B_{D}).$
E.2.3 BNN Predictor
Given $(A,B)$ and induced weights $W_{i}=A_{i}B_{i}^{\top}$, let $f_{A,B}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ denote the corresponding network (with fixed activation $\phi$ applied coordinate-wise). The posterior-mean predictor is
$f_{\mathrm{BNN}}(x;\theta)=\mathbb{E}_{(A,B)\sim q(\cdot\mid\theta)}\big[f_{A,B}(x)\big].$
E.2.4 Hypothesis Classes
We define three function classes that will be related through inclusions.
Definition E.1 (Deterministic low-rank class of Pinto et al. (2025)).
Fix rank bounds $r=(r_{1},\ldots,r_{D})$ and spectral norm bounds $C=(C_{1},\ldots,C_{D})$. Let $\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)$ denote the set of depth-$D$ networks
$\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)=\left\{f_{W}:\;\|W_{i}\|_{2}\leq C_{i},\;\mathrm{rank}(W_{i})\leq r_{i},\;i=1,\ldots,D\right\}.$
Definition E.2 (Posterior support class).
Define the support class induced by $q(\cdot\mid\theta)$ as
$\mathcal{F}^{\mathrm{supp}}_{D}(\theta)=\left\{f_{A,B}:\;(A_{i},B_{i})\in\mathrm{supp}(q_{A}(\cdot\mid\theta_{i}^{A}))\times\mathrm{supp}(q_{B}(\cdot\mid\theta_{i}^{B}))\;\;\forall i\right\}.$
Note that this class is deterministic: $A$ and $B$ are not sampled from $q$, they are deterministic matrices chosen to belong to the support of $q$.
Definition E.3 (Posterior-mean hypothesis class).
Define the class containing the posterior-mean predictor as the closed convex hull of the support class:
$\mathcal{F}^{\mathrm{BNN}}_{D}(\theta):=\overline{\mathrm{conv}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big)}.$
The closure is taken in the pointwise topology on $\mathbb{R}^{d}$. Equivalently, it is sufficient to take closure on the finite set $\{x_{1},\ldots,x_{m}\}$ for generalization analysis (see Lemma E.10).
Our goal is to control the generalization of the specific predictor $f_{\mathrm{BNN}}(\cdot;\theta)$, but it will be convenient to bound the complexity of $\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)$ via $\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)$.
E.3 Bounded Factor Support and Induced Weight Constraints
To connect $\mathcal{F}^{\mathrm{supp}}_{D}(\theta)$ with $\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)$, we require that the variational posterior has support contained within spectral norm balls.
Assumption E.4 (Bounded factor support).
For each layer $i=1,\ldots,D$, there exist constants $C_{i}^{A},C_{i}^{B}>0$ such that
$\mathrm{supp}(q_{A}(\cdot\mid\theta_{i}^{A}))\subseteq\{A_{i}:\|A_{i}\|_{2}\leq C_{i}^{A}\},\qquad\mathrm{supp}(q_{B}(\cdot\mid\theta_{i}^{B}))\subseteq\{B_{i}:\|B_{i}\|_{2}\leq C_{i}^{B}\}.$
Equivalently, $\|A_{i}\|_{2}\leq C_{i}^{A}$ and $\|B_{i}\|_{2}\leq C_{i}^{B}$ hold with probability $1$ under $q(A,B\mid\theta)$.
This is an almost-sure requirement. We will discuss practical enforcement and high-probability alternatives in Sections E.11 and E.10.
Lemma E.5 (Induced rank and spectral norm bounds for $W_{i}=A_{i}B_{i}^{\top}$).
Under Assumption E.4, define $C_{i}:=C_{i}^{A}C_{i}^{B}$. Then for each layer $i$, every realization $W_{i}=A_{i}B_{i}^{\top}$ satisfies
$\mathrm{rank}(W_{i})\leq r_{i},\qquad\|W_{i}\|_{2}\leq C_{i}.$
Proof.
The rank bound is immediate from $\mathrm{rank}(A_{i}B_{i}^{\top})\leq\min(\mathrm{rank}(A_{i}),\mathrm{rank}(B_{i}))\leq r_{i}$.
For the spectral norm, we use the submultiplicativity property:
$\|W_{i}\|_{2}=\|A_{i}B_{i}^{\top}\|_{2}\leq\|A_{i}\|_{2}\|B_{i}^{\top}\|_{2}=\|A_{i}\|_{2}\|B_{i}\|_{2}\leq C_{i}^{A}C_{i}^{B}=C_{i}.$
Lemma E.6 (Support class inclusion into Pinto’s class).
Under Assumption E.4,
$\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\subseteq\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\qquad\text{with }C_{i}=C_{i}^{A}C_{i}^{B}.$
Proof.
Take any $f_{A,B}\in\mathcal{F}^{\mathrm{supp}}_{D}(\theta)$. By Definition E.2, each factor realization lies in the corresponding support set, hence satisfies $\|A_{i}\|_{2}\leq C_{i}^{A}$ and $\|B_{i}\|_{2}\leq C_{i}^{B}$ by Assumption E.4.
Lemma E.5 yields $\|W_{i}\|_{2}\leq C_{i}$ and $\mathrm{rank}(W_{i})\leq r_{i}$ for all $i=1,\ldots,D$.
Therefore the corresponding network belongs to $\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)$ by Definition E.1. ∎
E.4 Vector-Valued Gaussian Complexity and Basic Properties
We now define the empirical vector-valued Gaussian complexity and establish the convex-hull and closure invariances that are essential for our main result.
Definition E.7 (Empirical vector-valued Gaussian complexity).
Let $\mathcal{F}$ be a class of functions $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$. Given sample inputs $S_{x}=\{x_{1},\ldots,x_{m}\}$, define
$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}):=\mathbb{E}_{\Gamma}\left[\sup_{f\in\mathcal{F}}\frac{1}{m}\sum_{j=1}^{m}\langle\gamma_{j},f(x_{j})\rangle\right],$
where $\gamma_{1},\ldots,\gamma_{m}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(0,I_{k})$ and $\Gamma=(\gamma_{1},\ldots,\gamma_{m})$.
Lemma E.8 (Monotonicity).
If $\mathcal{F}\subseteq\mathcal{G}$ then $\hat{\mathcal{G}}_{S_{x}}(\mathcal{F})\leq\hat{\mathcal{G}}_{S_{x}}(\mathcal{G})$.
Proof.
Fix a realization of $\Gamma=(\gamma_{1},\ldots,\gamma_{m})$. Define the functional
$\Phi(f):=\frac{1}{m}\sum_{j=1}^{m}\langle\gamma_{j},f(x_{j})\rangle.$
Since $\mathcal{F}\subseteq\mathcal{G}$, we have
$\sup_{f\in\mathcal{F}}\Phi(f)\leq\sup_{g\in\mathcal{G}}\Phi(g)$
for every realization of $\Gamma$.
Taking expectations over $\Gamma$ on both sides:
$\mathbb{E}_{\Gamma}\left[\sup_{f\in\mathcal{F}}\Phi(f)\right]\leq\mathbb{E}_{\Gamma}\left[\sup_{g\in\mathcal{G}}\Phi(g)\right],$
which is precisely $\hat{\mathcal{G}}_{S_{x}}(\mathcal{F})\leq\hat{\mathcal{G}}_{S_{x}}(\mathcal{G})$. ∎
Lemma E.9 (Convex hull invariance).
For any class $\mathcal{F}$,
$\hat{\mathcal{G}}_{S_{x}}(\mathrm{conv}(\mathcal{F}))=\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}).$
Proof.
The inequality $\hat{\mathcal{G}}_{S_{x}}(\mathcal{F})\leq\hat{\mathcal{G}}_{S_{x}}(\mathrm{conv}(\mathcal{F}))$ follows immediately from $\mathcal{F}\subseteq\mathrm{conv}(\mathcal{F})$ and Lemma E.8.
For the reverse inequality, fix $\Gamma$ and let $g\in\mathrm{conv}(\mathcal{F})$. Then by definition of the convex hull, there exist finitely many functions $f^{(1)},\ldots,f^{(T)}\in\mathcal{F}$ and coefficients $\alpha_{1},\ldots,\alpha_{T}\geq 0$ with $\sum_{t=1}^{T}\alpha_{t}=1$ such that
$g=\sum_{t=1}^{T}\alpha_{t}f^{(t)}.$
By linearity of the inner product and summation:
$\Phi(g)=\frac{1}{m}\sum_{j=1}^{m}\left\langle\gamma_{j},\sum_{t=1}^{T}\alpha_{t}f^{(t)}(x_{j})\right\rangle=\sum_{t=1}^{T}\alpha_{t}\left(\frac{1}{m}\sum_{j=1}^{m}\langle\gamma_{j},f^{(t)}(x_{j})\rangle\right)=\sum_{t=1}^{T}\alpha_{t}\Phi(f^{(t)}).$
Since $\alpha_{t}\geq 0$ and $\sum_{t=1}^{T}\alpha_{t}=1$, this is a convex combination:
$\Phi(g)=\sum_{t=1}^{T}\alpha_{t}\Phi(f^{(t)})\leq\sum_{t=1}^{T}\alpha_{t}\sup_{f\in\mathcal{F}}\Phi(f)=\sup_{f\in\mathcal{F}}\Phi(f).$
Taking the supremum over all $g\in\mathrm{conv}(\mathcal{F})$:
$\sup_{g\in\mathrm{conv}(\mathcal{F})}\Phi(g)\leq\sup_{f\in\mathcal{F}}\Phi(f).$
This holds for every realization of $\Gamma$, so taking expectations:
$\hat{\mathcal{G}}_{S_{x}}(\mathrm{conv}(\mathcal{F}))\leq\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}).$
Combined with the forward inequality, we obtain equality. ∎
Lemma E.10 (Closure invariance on the sample).
Let $\overline{\mathcal{F}}^{S_{x}}$ be the closure of $\mathcal{F}$ under pointwise convergence on $S_{x}$ (i.e., closure of evaluation vectors in $\mathbb{R}^{k\times m}$). Then
$\hat{\mathcal{G}}_{S_{x}}(\overline{\mathcal{F}}^{S_{x}})=\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}).$
Consequently,
$\hat{\mathcal{G}}_{S_{x}}(\overline{\mathrm{conv}(\mathcal{F})}^{S_{x}})=\hat{\mathcal{G}}_{S_{x}}(\mathrm{conv}(\mathcal{F}))=\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}).$
Proof.
Fix a realization $\Gamma=(\gamma_{1},\ldots,\gamma_{m})$. The functional
$\Phi(f)=\frac{1}{m}\sum_{j=1}^{m}\langle\gamma_{j},f(x_{j})\rangle$
depends only on the evaluation vector $(f(x_{1}),\ldots,f(x_{m}))\in\mathbb{R}^{k\times m}$.
We can rewrite $\Phi$ as:
$\Phi(f)=\left\langle\frac{1}{m}\sum_{j=1}^{m}\gamma_{j}\otimes e_{j},(f(x_{1}),\ldots,f(x_{m}))\right\rangle_{\mathbb{R}^{k\times m}},$
where $e_{j}$ is the $j$-th standard basis vector in $\mathbb{R}^{m}$ and $\otimes$ denotes the outer product. This is a linear functional on $\mathbb{R}^{k\times m}$, hence continuous with respect to the Euclidean topology.
Since $\Phi$ is continuous in the evaluation vector, and $\overline{\mathcal{F}}^{S_{x}}$ is precisely the closure of $\mathcal{F}$ with respect to pointwise convergence on $S_{x}$, we have:
$\sup_{f\in\overline{\mathcal{F}}^{S_{x}}}\Phi(f)=\sup_{f\in\mathcal{F}}\Phi(f).$
This holds for every realization of $\Gamma$, so taking expectations:
$\hat{\mathcal{G}}_{S_{x}}(\overline{\mathcal{F}}^{S_{x}})=\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}).$
For the second statement, apply Lemma E.9 to obtain $\hat{\mathcal{G}}_{S_{x}}(\mathrm{conv}(\mathcal{F}))=\hat{\mathcal{G}}_{S_{x}}(\mathcal{F})$, then apply the first part with $\mathcal{F}$ replaced by $\mathrm{conv}(\mathcal{F})$ to get:
$\hat{\mathcal{G}}_{S_{x}}(\overline{\mathrm{conv}(\mathcal{F})}^{S_{x}})=\hat{\mathcal{G}}_{S_{x}}(\mathrm{conv}(\mathcal{F}))=\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}).$
E.5 Posterior Mean Lies in Closed Convex Hull
Lemma E.11 (Posterior mean predictor is in the closed convex hull).
Assume that for each fixed $x\in\mathbb{R}^{d}$, the random vector $f_{A,B}(x)$ is integrable: $\mathbb{E}\|f_{A,B}(x)\|_{2}<\infty$. Then
$f_{\mathrm{BNN}}(\cdot;\theta)\in\overline{\mathrm{conv}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big)}.$
Proof.
Let $(A^{(t)},B^{(t)})_{t\geq 1}$ be i.i.d. samples from $q(\cdot\mid\theta)$ and define the Monte Carlo predictor
$\hat{f}_{M}(x):=\frac{1}{M}\sum_{t=1}^{M}f_{A^{(t)},B^{(t)}}(x).$
For each $M$, the predictor $\hat{f}_{M}$ is a convex combination of elements of $\mathcal{F}^{\mathrm{supp}}_{D}(\theta)$. To see this, note that each $(A^{(t)},B^{(t)})$ is drawn from the support of $q(\cdot\mid\theta)$, hence by Definition E.2, $f_{A^{(t)},B^{(t)}}\in\mathcal{F}^{\mathrm{supp}}_{D}(\theta)$ for all $t=1,\ldots,M$.
Therefore,
$\hat{f}_{M}=\frac{1}{M}\sum_{t=1}^{M}f_{A^{(t)},B^{(t)}}\in\mathrm{conv}(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)).$
By the strong law of large numbers, applied coordinate-wise in $\mathbb{R}^{k}$, and using the integrability assumption $\mathbb{E}\|f_{A,B}(x)\|_{2}<\infty$, we have for each fixed $x\in\mathbb{R}^{d}$:
$\hat{f}_{M}(x)\to\mathbb{E}_{(A,B)\sim q(\cdot\mid\theta)}[f_{A,B}(x)]=f_{\mathrm{BNN}}(x;\theta)\quad\text{almost surely as }M\to\infty.$
Since $\hat{f}_{M}\in\mathrm{conv}(\mathcal{F}^{\mathrm{supp}}_{D}(\theta))$ for all $M$, and $\hat{f}_{M}$ converges pointwise to $f_{\mathrm{BNN}}$, we conclude that
$f_{\mathrm{BNN}}(\cdot;\theta)\in\overline{\mathrm{conv}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big)},$
where the closure is taken in the pointwise topology (or equivalently, in the topology of pointwise convergence on any countable dense subset of $\mathbb{R}^{d}$, which includes the finite sample set $\{x_{1},\ldots,x_{m}\}$). ∎
Remark E.12 (Integrability assumption).
The integrability condition $\mathbb{E}\|f_{A,B}(x)\|_{2}<\infty$ is satisfied in practice when:
  • The factors $A_{i},B_{i}$ have finite second moments (e.g., Gaussian variational posteriors), and
  • The network depth $D$ and activation functions are such that the composition does not produce exponential growth.
Under Assumption E.4, which requires almost-sure spectral norm bounds on the factors, this integrability is automatically satisfied since the network outputs are then almost-surely bounded.
E.6 Complexity Chain: From BNN Class to Pinto’s Class
We now establish the key inclusion and complexity inequality that connects the Bayesian predictive mean to Pinto’s deterministic class.
Proposition E.13 (Gaussian complexity chain for the BNN class).
Under Assumption E.4, with $C_{i}=C_{i}^{A}C_{i}^{B}$,
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)\big)\leq\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big).$
Moreover, $\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)\subseteq\overline{\mathrm{conv}}\!\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)$.
Proof.
We establish the result through a chain of inclusions and applications of the invariance lemmas.
Step 1: Support class is contained in Pinto’s class.
By Lemma E.6, under Assumption E.4,
$\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\subseteq\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r).$
Step 2: Convex hull and closure preserve inclusion.
Taking the convex hull of both sides:
$\mathrm{conv}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big)\subseteq\mathrm{conv}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big).$
Taking the closure (on the sample $S_{x}$) of both sides:
$\overline{\mathrm{conv}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big)}^{S_{x}}\subseteq\overline{\mathrm{conv}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)}^{S_{x}}.$
By Definition E.3, the left-hand side is $\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)$, so:
$\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)\subseteq\overline{\mathrm{conv}}\!\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big).$
This establishes the second claim.
Step 3: Apply closure and convex hull invariance for complexity.
By Lemma E.10, the Gaussian complexity is invariant under closure on the sample:
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)\big)=\hat{\mathcal{G}}_{S_{x}}\big(\overline{\mathrm{conv}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big)}^{S_{x}}\big)=\hat{\mathcal{G}}_{S_{x}}\big(\mathrm{conv}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big)\big).$
By Lemma E.9, the Gaussian complexity is invariant under convex hull:
$\hat{\mathcal{G}}_{S_{x}}\big(\mathrm{conv}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big)\big)=\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big).$
Combining:
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)\big)=\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big).$
Step 4: Apply monotonicity.
Since $\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\subseteq\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)$ by Step 1, Lemma E.8 gives:
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{supp}}_{D}(\theta)\big)\leq\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big).$
Combining with Step 3:
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)\big)\leq\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big).$
This completes the proof. ∎
Remark E.14 (Significance of the complexity chain).
This proposition is the central technical result enabling the transfer of deterministic bounds to Bayesian predictors. It shows that:
  1. 1.
    The BNN predictor class has the same Gaussian complexity as the support class, despite being a much larger set (closed convex hull vs. original class).
  2. 2.
    The complexity is bounded by that of Pinto’s deterministic class, allowing us to invoke existing deterministic bounds.
  3. 3.
    No degradation occurs in the complexity bound from the deterministic to the Bayesian setting.
E.7 Generalization from Vector-Valued Gaussian Complexity
We now state the generalization theorem from Pinto et al. (2025) for vector-valued Gaussian complexity and Lipschitz losses, which we will apply to our BNN predictor.
Theorem E.15 (Generalization via vector-valued Gaussian complexity, Pinto et al. (2025)).
Let $\ell(\cdot,y)\in[0,1]$ be $\mathcal{L}$-Lipschitz in its first argument. Let $\mathcal{F}$ be a class of functions $\mathbb{R}^{d}\to\mathbb{R}^{k}$. Then, with probability at least $1-\delta$ over $S\sim\mathcal{D}^{m}$, for all $f\in\mathcal{F}$,
$\mathbb{E}\big[\ell(f(x),y)\big]\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f(x_{j}),y_{j})+\sqrt{\pi\mathcal{L}}\;\hat{\mathcal{G}}_{S_{x}}(\mathcal{F})+3\sqrt{\frac{\log(2/\delta)}{2m}}.$
Remark E.16 (Origin and scope of the theorem).
Theorem E.15 combines two results from Pinto et al. (2025):
  1. 1.
    A contraction inequality showing that Lipschitz losses contract the vector-valued Gaussian complexity by at most $\sqrt{\pi\mathcal{L}}$.
  2. 2.
    A concentration inequality relating the empirical Gaussian complexity to generalization.
We cite it in this combined form to avoid re-deriving technical constants. Our contribution is the class inclusion argument (Proposition E.13) that enables applying this theorem to $f_{\mathrm{BNN}}$.
E.7.1 Explicit Complexity Bound for Pinto’s Class
We now recall the explicit upper bound for $\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r))$ derived by Pinto et al. (2025).
Theorem E.17 (Explicit bound for rank- and norm-constrained networks, Pinto et al. (2025)).
For $\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)$ as in Definition E.1, the following bound holds:
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{\|W_{1}\|_{F}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\},$
where:
  • $C_{0}>0$ is the universal constant from Maurer’s chaining bound for Lipschitz activations,
  • $\|X\|_{F}^{2}=\sum_{j=1}^{m}\|x_{j}\|_{2}^{2}$ is the Frobenius norm of the input data matrix,
  • $h_{\ell}$ denotes the width of layer $\ell$,
  • $r_{j}$ is the rank bound for layer $j$.
Remark E.18 (Key features of the bound).
  1. 1.
    Rank dependence: The complexity scales as $\sqrt{r_{j}h_{j}}$ for each layer $j\geq 2$, rather than the full dimension $h_{j}h_{j-1}$ that would appear in a full-rank analysis.
  2. 2.
    Data dependence: The bound explicitly depends on $\|X\|_{F}$, the norm of the training inputs, unlike PAC-Bayes bounds which depend on the posterior-prior KL divergence.
  3. 3.
    Depth dependence: The product $\prod_{i=j+1}^{D}C_{0}C_{i}$ shows how spectral norms accumulate through the network depth.
  4. 4.
    First layer: The term $\|W_{1}\|_{F}\sqrt{h_{1}}$ for the first layer appears because the first layer is not constrained by rank in the bound, only by spectral norm and Frobenius norm.
Remark E.19 (Comparison with standard Rademacher bounds).
Standard Rademacher complexity bounds for deep networks without rank constraints typically scale as
$O\left(\frac{\|X\|_{F}}{m}\prod_{i=1}^{D}C_{i}\sqrt{h_{i}h_{i-1}}\right).$
The rank constraint reduces $\sqrt{h_{i}h_{i-1}}$ to $\sqrt{r_{i}h_{i}}$ in the bound, providing improvement when $r_{i}\ll h_{i-1}$.
Note on depth dependence: Both the rank-constrained bound from Theorem E.17 and standard bounds contain a product over depth, specifically the factor $\prod_{i=j}^{D}C_{0}C_{i}$ which grows as $C_{0}^{D-j+1}\prod_{i=j}^{D}C_{i}$. As noted by Pinto et al. (2025), this unfavorable exponential factor in depth may be an artifact of the current proof technique rather than a limitation. When comparing the rank-constrained bound to existing bounds, the key improvement comes from the rank-dependent terms $\sqrt{r_{j}h_{j}}$ rather than from the depth-dependent product, which appears in both settings.
E.8 Proof of Theorem 3.9
Proof.
We prove that the BNN posterior predictive mean satisfies the same generalization bound as Pinto’s deterministic class.
Step 1: The BNN predictor belongs to the BNN class.
By Lemma E.11, under the integrability assumption $\mathbb{E}\|f_{A,B}(x)\|_{2}<\infty$,
$f_{\mathrm{BNN}}(\cdot;\theta)\in\mathcal{F}^{\mathrm{BNN}}_{D}(\theta).$
Step 2: Apply Pinto’s generalization theorem to the BNN class.
We apply Theorem E.15 with the hypothesis class $\mathcal{F}=\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)$.
With probability at least $1-\delta$ over $S\sim\mathcal{D}^{m}$, the bound holds uniformly for all $f\in\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)$:
$\mathbb{E}\big[\ell(f(x),y)\big]\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f(x_{j}),y_{j})+\sqrt{\pi\mathcal{L}}\;\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)\big)+3\sqrt{\frac{\log(2/\delta)}{2m}}.$
In particular, since $f_{\mathrm{BNN}}(\cdot;\theta)\in\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)$ by Step 1, this bound holds for $f=f_{\mathrm{BNN}}(\cdot;\theta)$.
Step 3: Bound the Gaussian complexity using Proposition E.13.
By Proposition E.13, under Assumption E.4,
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{BNN}}_{D}(\theta)\big)\leq\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big),$
where $C_{i}=C_{i}^{A}C_{i}^{B}$ for each layer $i=1,\ldots,D$.
Step 4: Apply the explicit bound from Theorem E.17.
Substituting the explicit bound from Theorem E.17:
$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{\|W_{1}\|_{F}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}.$
Step 5: Combine to obtain the final bound.
Combining Steps 2, 3, and 4, we obtain that with probability at least $1-\delta$ over $S\sim\mathcal{D}^{m}$,
$\mathbb{E}\big[\ell(f_{\mathrm{BNN}}(x),y)\big]$ $\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f_{\mathrm{BNN}}(x_{j}),y_{j})$
$\quad+\sqrt{\pi\mathcal{L}}\cdot\frac{\|X\|_{F}}{m}\left\{\|W_{1}\|_{F}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}$
$\quad+3\sqrt{\frac{\log(2/\delta)}{2m}}.$
This completes the proof. ∎
Remark E.20 (Key insight of the proof).
The proof relies on three important properties:
  1. 1.
    Membership: The BNN predictor belongs to the closed convex hull of the support class (Lemma E.11).
  2. 2.
    Invariance: Vector-valued Gaussian complexity is invariant under convex hull and closure (Lemmas E.9 and E.10).
  3. 3.
    Inclusion: The support class is contained in Pinto’s deterministic class under spectral norm bounds (Lemma E.6).
These three properties together allow the deterministic rank-sensitive bound to transfer without degradation to the Bayesian predictive mean, despite the fact that rank is not preserved under expectation.
E.9 Comparison with PAC-Bayes Bounds
The Gaussian complexity bounds we have established complement rather than replace PAC-Bayes bounds. Here we compare their structure and highlight when each approach provides better guarantees.
Proposition E.21 (Scaling Comparison).
Consider a single-layer network with input dimension $d_{\text{in}}$, output dimension $d_{\text{out}}$, and rank $r$.
PAC-Bayes bound (from Theorem 3.8): The generalization gap is bounded by
$\text{Gap}_{\text{PAC}}=O\left(\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})}{m}}\right),$ (15)
with rank dependence $\sqrt{r}$ and uniform $\frac{1}{\sqrt{m}}$ scaling in sample size (up to logarithmic factors).
Gaussian complexity bound (from Theorem 3.9): The generalization gap is bounded by
$\text{Gap}_{\text{GC}}=O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right)+O\left(\frac{1}{\sqrt{m}}\right),$ (16)
where:
  • The first term has rank dependence $\sqrt{r}$ and scales as $\frac{1}{m}$ (for fixed $\|X\|_{F}$);
  • The second term (concentration) scales as $\frac{1}{\sqrt{m}}$ and is rank-independent;
  • For large $m$, the $O(1/\sqrt{m})$ term dominates;
  • For moderate $m$ with large $\|X\|_{F}$, the $O(1/m)$ term may dominate.
If, in addition, inputs are uniformly bounded as $\|x_{j}\|_{2}\leq R$, then $\|X\|_{F}=O(\sqrt{m})$ and the rank-dependent term becomes $O(\sqrt{rd_{\text{out}}/m})$.
Proof.
PAC-Bayes analysis. The complexity term in Theorem 3.8 is
$\sqrt{\frac{\mathrm{KL}(Q_{\text{LR}}\|P_{\text{LR}})+\log(2\sqrt{m}/\delta)}{2m}}\leq\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})+\log(2\sqrt{m}/\delta)}{2m}}.$
Ignoring logarithmic terms and constants, this scales as
$O\left(\sqrt{\frac{r(d_{\text{in}}+d_{\text{out}})}{m}}\right)=\sqrt{r}\cdot O\left(\sqrt{\frac{d_{\text{in}}+d_{\text{out}}}{m}}\right),$
so the rank dependence is $\sqrt{r}$ and the sample-size dependence is $1/\sqrt{m}$.
Gaussian complexity analysis. For a single-layer network (depth $D=1$), Theorem E.17 gives
$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})\leq\frac{\|X\|_{F}}{m}\|W_{1}\|_{F}\sqrt{h_{1}}C_{0}C_{1},$
where $h_{1}=d_{\text{out}}$. For a rank-$r$ matrix $W_{1}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ with $\|W_{1}\|_{2}\leq C_{1}$, we have
$\|W_{1}\|_{F}\leq\sqrt{r}\,\|W_{1}\|_{2}\leq\sqrt{r}C_{1}.$
Substituting with $h_{1}=d_{\text{out}}$,
$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})$ $\leq\frac{\|X\|_{F}}{m}\cdot\sqrt{r}C_{1}\cdot\sqrt{d_{\text{out}}}\cdot C_{0}C_{1}$
$=\frac{\|X\|_{F}\cdot C_{0}C_{1}^{2}\cdot\sqrt{rd_{\text{out}}}}{m}$
$=O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right).$
From Theorem E.15, the full generalization bound is
$\mathbb{E}[\ell(f(x),y)]$ $\leq\hat{L}(f)+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F})+3\sqrt{\frac{\log(2/\delta)}{2m}}$
$=\hat{L}(f)+O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right)+O\left(\frac{1}{\sqrt{m}}\right),$
where the first term is rank-dependent and decays as $1/m$ (for fixed $\|X\|_{F}$), while the second term is rank-independent and decays as $1/\sqrt{m}$. This yields the claimed form for $\text{Gap}_{\text{GC}}$.
If we additionally assume $\|x_{j}\|_{2}\leq R$ for all $j$, then
$\|X\|_{F}^{2}=\sum_{j=1}^{m}\|x_{j}\|_{2}^{2}\leq mR^{2}\quad\Rightarrow\quad\|X\|_{F}\leq R\sqrt{m},$
so the rank-dependent term simplifies to
$O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right)=O\left(\frac{\sqrt{m}\sqrt{rd_{\text{out}}}}{m}\right)=O\left(\sqrt{\frac{rd_{\text{out}}}{m}}\right).$
Remark E.22 (Structural differences).
The PAC-Bayes and Gaussian complexity bounds have fundamentally different structures:
PAC-Bayes: Single term with uniform $\frac{1}{\sqrt{m}}$ scaling
$\text{Gap}_{\text{PAC}}=O\left(\sqrt{\frac{\text{KL divergence}}{m}}\right).$
Gaussian complexity: Two terms with different scalings
$\text{Gap}_{\text{GC}}=\underbrace{O\left(\frac{\text{Complexity}\times\|X\|_{F}}{m}\right)}_{\text{faster decay in }m}+\underbrace{O\left(\frac{1}{\sqrt{m}}\right)}_{\text{slower decay in }m}.$
For large $m$, the slower-decaying $O(1/\sqrt{m})$ term dominates; for moderate $m$ and large $\|X\|_{F}\sqrt{rd_{\text{out}}}$, the $O(1/m)$ term can dominate.
E.9.1 Complementary Strengths
Remark E.23 (When PAC-Bayes is stronger).
PAC-Bayes bounds are tighter when:
  1. 1.
    Large sample regime: For large $m$, both bounds scale as $O(1/\sqrt{m})$, but PAC-Bayes may have better constants when the posterior is concentrated (small $C_{\max}$).
  2. 2.
    Posterior concentration: When the posterior is highly concentrated around a good solution, $C_{\max}$ can be very small, making the KL term tight.
  3. 3.
    Large input norms: PAC-Bayes bounds do not depend on $\|X\|_{F}$, which is advantageous when inputs have very large norm or heavy tails.
  4. 4.
    Dimension regime: When $r(d_{\text{in}}+d_{\text{out}})$ is small relative to $\|X\|_{F}^{2}d_{\text{out}}$, the PAC-Bayes bound can be more favorable.
Remark E.24 (When Gaussian complexity is stronger).
Gaussian complexity bounds might be tighter when:
  1. 1.
    Small to moderate sample sizes: When $m$ is not too large and $\|X\|_{F}\sqrt{rd_{\text{out}}}=O(\sqrt{m})$, the $O(1/m)$ term provides faster decay than the $O(1/\sqrt{m})$ concentration term.
  2. 2.
    Normalized or bounded inputs: When $\|X\|_{F}=O(\sqrt{m})$ (e.g., inputs normalized to unit norm), the rank-dependent term becomes $O(\sqrt{rd_{\text{out}}/m})$, which can be tighter than PAC-Bayes when $rd_{\text{out}}\ll r(d_{\text{in}}+d_{\text{out}})$.
  3. 3.
    No prior specification required: Unlike PAC-Bayes, which requires choosing a prior $P$, Gaussian complexity bounds depend only on the realized network architecture and data.
  4. 4.
    Architectural transparency: The bound explicitly shows how rank, spectral norms, and layer dimensions interact through the network depth, providing architectural guidance.
  5. 5.
    Connection to deterministic theory: Gaussian complexity bounds inherit the full machinery of deterministic generalization theory, including recent advances in rank-sensitive analysis.
Remark E.25 (Rank dependence).
Both bounds have $\sqrt{r}$ rank dependence in their leading terms:
  • PAC-Bayes: The parameter-count reduction from $d_{\text{in}}d_{\text{out}}$ to $r(d_{\text{in}}+d_{\text{out}})$ appears inside a square root in the KL term.
  • Gaussian complexity: The Frobenius norm bound $\|W\|_{F}\leq\sqrt{r}\|W\|_{2}$ for rank-$r$ matrices introduces $\sqrt{r}$ into the complexity.
Neither bound exhibits linear $r$ dependence; the rank always appears under a square root in the leading generalization term.
Remark E.26.
The key insight is that these bounds measure different aspects of generalization:
  • PAC-Bayes measures how much the posterior deviates from the prior (information-theoretic complexity).
  • Gaussian complexity measures the richness of the hypothesis class on the actual training data (geometric complexity).
In practice, both bounds should be considered:
  • Use PAC-Bayes to understand posterior–prior relationships and to guide prior selection.
  • Use Gaussian complexity to understand architectural choices (rank, width, depth, spectral norms) and their interaction with the training data.
Our contribution shows that Bayesian predictive means can take advantage of both frameworks: they satisfy PAC-Bayes bounds through posterior analysis, and they satisfy Gaussian complexity bounds through the convex-hull structure established in this section.
E.10 Alternative Approaches for High-Probability Spectral Norm Bounds
Assumption E.4 requires that the variational posterior has support contained within a spectral norm ball, that is, the bound $\|A_{i}\|_{2}\leq C_{i}^{A}$ must hold almost surely for every sample from the posterior. This assumption is essential for the submultiplicativity argument in Lemma E.5 which establishes that the induced weight matrices $W_{i}=A_{i}B_{i}^{\top}$ satisfy the spectral norm constraints required for $\mathcal{F}_{D}^{\mathrm{Pinto}}$.
However, in practice, commonly used variational families such as Gaussian posteriors have unbounded support. While concentration inequalities (such as Van Handel’s result in (Van Handel, 2017)) provide high-probability bounds, they do not guarantee almost-sure containment. This section presents two alternative approaches that allow our framework to accommodate high-probability bounds rather than requiring almost-sure bounds.
E.10.1 Setup and Notation
Let $q(A,B\mid\theta)$ denote the variational posterior over the factor matrices. For each layer $i$, suppose we have high-probability bounds:
$\mathbb{P}_{A_{i}\sim q_{A}(\cdot\mid\theta_{i})}\left(\|A_{i}\|_{2}\leq C_{i}^{A}\right)$ $\geq 1-\delta_{i}^{A},$ (17)
$\mathbb{P}_{B_{i}\sim q_{B}(\cdot\mid\theta_{i})}\left(\|B_{i}\|_{2}\leq C_{i}^{B}\right)$ $\geq 1-\delta_{i}^{B}.$ (18)
Define the ”good event” $\mathcal{E}_{\text{good}}$ as the intersection of all these events across all $D$ layers:
$\mathcal{E}_{\text{good}}:=\bigcap_{i=1}^{D}\left\{\|A_{i}\|_{2}\leq C_{i}^{A}\right\}\cap\left\{\|B_{i}\|_{2}\leq C_{i}^{B}\right\}.$ (19)
By a union bound:
$\mathbb{P}(\mathcal{E}_{\text{good}})\geq 1-\sum_{i=1}^{D}(\delta_{i}^{A}+\delta_{i}^{B})=:1-\delta_{\text{total}}.$ (20)
When $\mathcal{E}_{\text{good}}$ occurs, all factor matrices satisfy their spectral norm bounds, and consequently all induced weight matrices $W_{i}=A_{i}B_{i}^{\top}$ satisfy $\|W_{i}\|_{2}\leq C_{i}:=C_{i}^{A}C_{i}^{B}$ and $\mathrm{rank}(W_{i})\leq r_{i}$ by the same argument as Lemma E.5.
E.10.2 Approach A: Decomposition of the BNN Predictor
The BNN predictor is defined as an expectation over the full posterior:
$f_{\mathrm{BNN}}(x;\theta)=\mathbb{E}_{(A,B)\sim q(\cdot\mid\theta)}[f_{A,B}(x)].$ (21)
We can decompose this expectation by conditioning on whether $\mathcal{E}_{\text{good}}$ occurs:
$f_{\mathrm{BNN}}(x;\theta)=\mathbb{E}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}]\cdot\mathbb{P}(\mathcal{E}_{\text{good}})+\mathbb{E}[f_{A,B}(x)\mid\mathcal{E}_{\text{bad}}]\cdot\mathbb{P}(\mathcal{E}_{\text{bad}}),$ (22)
where $\mathcal{E}_{\text{bad}}=\mathcal{E}_{\text{good}}^{c}$.
Analysis of the ”good” term.
Define the conditional predictor:
$f_{\mathrm{BNN}}^{\text{good}}(x;\theta):=\mathbb{E}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}].$ (23)
This conditional expectation is taken only over samples that satisfy the spectral norm bounds. Replacing the full posterior with the conditional distribution given $\mathcal{E}_{\text{good}}$, we have by the same argument as Lemma E.11:
$f_{\mathrm{BNN}}^{\text{good}}(\cdot;\theta)\in\overline{\mathrm{conv}(\mathcal{F}_{D}^{\mathrm{supp},\text{good}}(\theta))},$ (24)
where $\mathcal{F}_{D}^{\mathrm{supp},\text{good}}(\theta)$ consists of network functions realized by samples in $\mathcal{E}_{\text{good}}$.
Since all such samples satisfy the spectral norm constraints by construction, we have:
$\mathcal{F}_{D}^{\mathrm{supp},\text{good}}(\theta)\subseteq\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r).$ (25)
Therefore, by the same arguments as Proposition E.13:
$\hat{\mathcal{G}}_{S_{x}}(\{f_{\mathrm{BNN}}^{\text{good}}\})\leq\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r)).$ (26)
Analysis of the ”bad” term.
The contribution from $\mathcal{E}_{\text{bad}}$ has small probability $\delta_{\text{total}}$. If the loss function $\ell$ is bounded in $[0,1]$, this contributes at most $\delta_{\text{total}}$ to the expected loss.
Combined generalization bound.
Let $\ell:\mathbb{R}^{k}\times\mathbb{R}^{k}\to[0,1]$ be an $\mathcal{L}$-Lipschitz loss function. The expected loss of the BNN predictor satisfies:
$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $=\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)\mid\mathcal{E}_{\text{good}}]\cdot\mathbb{P}(\mathcal{E}_{\text{good}})$
$\quad+\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)\mid\mathcal{E}_{\text{bad}}]\cdot\mathbb{P}(\mathcal{E}_{\text{bad}})$
$\leq\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)]+1\cdot\delta_{\text{total}}.$ (27)
Applying the generalization bound from Theorem E.15 to $f_{\mathrm{BNN}}^{\text{good}}$ (following the proof of Theorem 3.9):
$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $\leq\frac{1}{m}\sum_{i=1}^{m}\ell(f_{\mathrm{BNN}}^{\text{good}}(x_{i}),y_{i})$
$\quad+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r))+3\sqrt{\frac{\log(2/\delta)}{2m}}+\delta_{\text{total}}.$ (28)
Remark E.27 (Practical Interpretation).
The additional term $\delta_{\text{total}}$ represents the ”cost” of using high-probability bounds instead of almost-sure bounds. For Gaussian posteriors with Van Handel-type bounds (see Remark E.30), choosing $\delta_{i}^{A}=\delta_{i}^{B}=\delta_{0}/(2D)$ for each layer gives $\delta_{\text{total}}=\delta_{0}$. The spectral norm bounds $C_{i}^{A}$ and $C_{i}^{B}$ then scale as $O(\sqrt{\log(D/\delta_{0})})$, which introduces a mild logarithmic dependence on the depth and confidence level.
E.10.3 Approach B: Modified Function Class with High-Probability Inclusion
An alternative approach modifies the function class to directly incorporate the high-probability nature of the constraints.
Definition.
Define the high-probability support class:
$\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta):=\left\{f_{W}:W_{i}=A_{i}B_{i}^{\top},(A_{i},B_{i})\in\mathrm{supp}(q(\cdot\mid\theta_{i}))\cap\mathcal{B}_{i},\;\forall i=1,\ldots,D\right\},$ (29)
where $\mathcal{B}_{i}:=\{(A_{i},B_{i}):\|A_{i}\|_{2}\leq C_{i}^{A}(\delta),\|B_{i}\|_{2}\leq C_{i}^{B}(\delta)\}$ and $C_{i}^{A}(\delta),C_{i}^{B}(\delta)$ are chosen so that
$\mathbb{P}((A_{i},B_{i})\in\mathcal{B}_{i})\geq 1-\delta/(2D).$ (30)
Properties.
  1. 1.
    $\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta)\subseteq\mathcal{F}_{D}^{\mathrm{Pinto}}(C(\delta),r)$ by construction, where $C_{i}(\delta)=C_{i}^{A}(\delta)C_{i}^{B}(\delta)$ (all functions satisfy the spectral norm and rank constraints).
  2. 2.
    The BNN predictor restricted to $\mathcal{E}_{\text{good}}$ lies in $\overline{\mathrm{conv}(\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta))}$ by the same argument as Lemma E.11.
  3. 3.
    $\mathbb{P}(\text{sample from }q\text{ produces function in }\mathcal{F}_{D}^{\mathrm{hp}})\geq 1-\delta$ by the union bound.
Generalization bound.
By the same reasoning as Approach A:
$\mathbb{E}[\ell(f_{\mathrm{BNN}}(x),y))]$ $\leq\frac{1}{m}\sum_{i=1}^{m}\ell(f_{\mathrm{BNN}}(x_{i}),y_{i})$
$\quad+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C(\delta),r))+3\sqrt{\frac{\log(2/\delta^{\prime})}{2m}}+\delta,$ (31)
where the Gaussian complexity bound uses the spectral norm constraints $C_{i}(\delta)=C_{i}^{A}(\delta)\cdot C_{i}^{B}(\delta)$ that depend on the confidence parameter $\delta$.
Remark E.28 (Comparison of approaches).
Both approaches yield the same final bound structure, with an additive $\delta$ term representing the high-probability relaxation.
  • Approach A is conceptually cleaner: it decomposes the predictor into good and bad events.
  • Approach B is more direct: it modifies the function class definition to explicitly encode the high-probability constraint.
In practice, either approach can be used depending on which is more convenient for the specific application.
E.11 Practical Enforcement Remarks
We now discuss practical methods for enforcing or approximating Assumption E.4, and provide specific concentration results for Gaussian variational posteriors.
Remark E.29 (Enforcement in practice).
Assumption E.4 can be satisfied (or approximated) by several practical approaches:
  1. 1.
    Projection / spectral normalization. After each parameter update during optimization, rescale factors to enforce spectral norm constraints:
    $A_{i}\leftarrow\min\!\left(1,\frac{C_{i}^{A}}{\|A_{i}\|_{2}}\right)A_{i},\qquad B_{i}\leftarrow\min\!\left(1,\frac{C_{i}^{B}}{\|B_{i}\|_{2}}\right)B_{i}.$
    In reparameterized variational inference, the same clipping can be applied to sampled factors. This ensures almost-sure satisfaction of the bounds.
  2. 2.
    Compact-support families. Use truncated Gaussians or uniform distributions supported on the spectral norm ball. For example, a truncated Gaussian with support $\{A:\|A\|_{2}\leq C^{A}\}$ naturally satisfies the assumption. This requires computing the normalization constant, which may be computationally expensive for high-dimensional matrices.
  3. 3.
    Soft constraints via penalties. Add regularization penalties that discourage violations:
    $\mathcal{L}_{\text{reg}}(\theta)=\lambda_{A}\sum_{i=1}^{D}\max(0,\|A_{i}\|_{2}-C_{i}^{A})^{2}+\lambda_{B}\sum_{i=1}^{D}\max(0,\|B_{i}\|_{2}-C_{i}^{B})^{2}.$
    This yields a high-probability rather than almost-sure guarantee unless combined with truncation, but may be easier to implement and optimize.
Remark E.30 (Van Handel-type spectral norm bounds for Gaussian factors).
If $A_{i}\in\mathbb{R}^{h_{i}\times r_{i}}$ has independent Gaussian entries with variance $\sigma_{i}^{2}$ (and possibly nonzero mean $\mu_{i}$), nonasymptotic concentration results provide sharp bounds on the spectral norm.
Specifically, by Van Handel’s matrix concentration inequality (Van Handel, 2017), there exist universal constants $c_{1},c_{2}>0$ such that for all $t\geq 0$,
$\mathbb{P}\!\left(\|A_{i}\|_{2}\leq\|\mathbb{E}[A_{i}]\|_{2}+c_{1}\sigma_{i}(\sqrt{h_{i}}+\sqrt{r_{i}})+c_{2}\sigma_{i}t\right)\geq 1-2e^{-t^{2}}.$
Setting $t=\sqrt{\log(2/\delta_{i})}$ yields a $(1-\delta_{i})$ high-probability bound:
$\mathbb{P}\!\left(\|A_{i}\|_{2}\leq\|\mathbb{E}[A_{i}]\|_{2}+c_{1}\sigma_{i}(\sqrt{h_{i}}+\sqrt{r_{i}})+c_{2}\sigma_{i}\sqrt{\log(2/\delta_{i})}\right)\geq 1-\delta_{i}.$
Recovering an almost-sure bound.
To recover an almost-sure bound from this high-probability statement, one may truncate the variational distribution outside the spectral norm ball:
$C_{i}^{A}:=\|\mathbb{E}[A_{i}]\|_{2}+c_{1}\sigma_{i}(\sqrt{h_{i}}+\sqrt{r_{i}})+c_{2}\sigma_{i}\sqrt{\log(2/\delta_{i})},$
at the cost of modifying the variational family from a standard Gaussian to a truncated Gaussian.
Retaining high-probability bounds.
Alternatively, one may retain the high-probability statement and use Approaches A or B from Section E.10 to carry an additional failure probability $\delta_{\text{total}}$ through the final generalization bound. This is often more practical than truncation, as it requires no modification to the variational family.
Practical recommendations.
In practice, we recommend:
  • Use spectral normalization (projection) during training to enforce hard constraints, which ensures almost-sure satisfaction at minimal computational cost.
  • Use Van Handel bounds to guide the choice of $C_{i}^{A}$ and $C_{i}^{B}$ based on the variational posterior parameters $\sigma_{i}$ and $\mu_{i}$.
  • If using unbounded variational families without projection, apply the high-probability analysis from Section E.10 with appropriately chosen $\delta_{\text{total}}$.
Remark E.31 (Tightness of Van Handel bounds).
The Van Handel bound is known to be tight up to universal constants. For a random matrix with independent Gaussian entries, the spectral norm concentrates around $\sigma(\sqrt{h}+\sqrt{r})$ with high probability. This means:
  • The constants $C_{i}^{A}$ and $C_{i}^{B}$ grow as $O(\sigma_{i}\sqrt{h_{i}+r_{i}})$ for Gaussian posteriors.
  • The induced spectral norm bounds $C_{i}=C_{i}^{A}C_{i}^{B}$ then scale as $O(\sigma_{i}^{2}(h_{i}+r_{i}))$.
  • These bounds are dimension-dependent but explicit and computable from the variational parameters.
Appendix F Principled Initialization of Low-Rank Variational Weight Matrices
F.1 Setting and Notation
Consider a linear layer with input dimension $n$ and output dimension $m$, parameterized by a weight matrix
$W\in\mathbb{R}^{n\times m}.$
Rather than parameterizing $W$ directly, we adopt a low-rank factorization
$W=AB^{\top},\quad A\in\mathbb{R}^{n\times r},\quad B\in\mathbb{R}^{m\times r},\quad r\ll\min(n,m).$
In a variational Bayesian formulation, we place factorized Gaussian variational posteriors over the entries of $A$ and $B$:
$q(A)=\prod_{i,k}\mathcal{N}(A_{ik}\mid\mu^{A}_{ik},(\sigma^{A}_{ik})^{2}),\qquad q(B)=\prod_{j,k}\mathcal{N}(B_{jk}\mid\mu^{B}_{jk},(\sigma^{B}_{jk})^{2}).$
The objective of initialization is to choose $(\mu^{A},\mu^{B},\sigma^{A},\sigma^{B})$ such that the induced distribution over $W$ yields stable forward signal propagation and well-conditioned optimization at the start of training.
F.2 Reference Dense Initialization and Variance Propagation
Initialization of deep networks is traditionally analyzed through variance propagation arguments. Let $x\in\mathbb{R}^{n}$ be an input vector with i.i.d. components satisfying
$\mathbb{E}[x_{i}]=0,\qquad\mathrm{Var}(x_{i})=1.$
For a dense linear layer $z=Wx$, sufficient conditions for stable signal propagation require that the entries of $W$ satisfy
$\mathbb{E}[W_{ij}]=0,\qquad\mathrm{Var}(W_{ij})=\sigma_{W}^{2},$
where $\sigma_{W}^{2}$ depends on $(n,m)$ and the nonlinearity applied after the linear transformation.
Glorot and Bengio (Glorot and Bengio, 2010) showed that choosing
$\sigma_{W}^{2}=\frac{2}{n+m}$
preserves the variance of activations across layers for linear and saturating nonlinearities.
He et al. (He et al., 2015) further derived that for ReLU activations, variance preservation requires
$\sigma_{W}^{2}=\frac{2}{n}.$
These results are textbook consequences of second-moment propagation and form the theoretical basis of modern dense initialization schemes.
In what follows, $\sigma_{W}^{2}$ is treated as a known constant determined solely by layer dimensions and activation function.
F.3 Induced Statistics of Low-Rank Weight Matrices
Let the variational means $\mu^{A}_{ik}$ and $\mu^{B}_{jk}$ be initialized as independent random variables with
$\mathbb{E}[\mu^{A}_{ik}]=\mathbb{E}[\mu^{B}_{jk}]=0,\qquad\mathrm{Var}(\mu^{A}_{ik})=v_{A},\quad\mathrm{Var}(\mu^{B}_{jk})=v_{B}.$
The mean of the induced weight distribution is
$\bar{W}:=\mathbb{E}_{q}[W]=\mu^{A}(\mu^{B})^{\top},$
with entries
$\bar{W}_{ij}=\sum_{k=1}^{r}\mu^{A}_{ik}\mu^{B}_{jk}.$
Since products of independent zero-mean random variables satisfy
$\mathbb{E}[\mu^{A}_{ik}\mu^{B}_{jk}]=0,\qquad\mathrm{Var}(\mu^{A}_{ik}\mu^{B}_{jk})=v_{A}v_{B},$
we obtain by independence and additivity of variance:
$\mathrm{Var}(\bar{W}_{ij})=\sum_{k=1}^{r}\mathrm{Var}(\mu^{A}_{ik}\mu^{B}_{jk})=r\,v_{A}v_{B}.$
This calculation follows directly from independence and the fact that the variance of a sum of independent random variables equals the sum of their variances.
F.4 Variance-Matching Condition
To ensure that the low-rank parameterization reproduces the second-order statistics of a properly initialized dense layer, we impose the condition
$\mathrm{Var}(\bar{W}_{ij})=\sigma_{W}^{2}.$
Substituting the expression from Section F.3 yields the constraint
$r\,v_{A}v_{B}=\sigma_{W}^{2}.$
Equation (F.4) defines a one-parameter family of valid initializations, parameterized by the choice of how to split the variance between $v_{A}$ and $v_{B}$.
F.5 Symmetric Choice and Identifiability Considerations
The factorization $W=AB^{\top}$ exhibits scale non-identifiability: for any $c>0$,
$A\mapsto cA,\quad B\mapsto c^{-1}B\quad\Rightarrow\quad AB^{\top}\text{ unchanged}.$
Such gauge freedoms are well-known in matrix factorizations and latent variable models (see, e.g., (Udell et al., 2015)).
A natural way to fix this non-identifiability at initialization is to impose symmetry between the factors:
$v_{A}=v_{B}.$
Combined with the variance-matching condition (F.4), this yields the unique symmetric solution
$v_{A}=v_{B}=\sqrt{\frac{\sigma_{W}^{2}}{r}}.$
This choice ensures:
  • Balanced gradient magnitudes for $A$ and $B$ during early training,
  • Isotropy of the variational posterior at initialization,
  • Consistency with the symmetric square-root parameterization $W=U\Sigma^{1/2}(\Sigma^{1/2}V^{\top})$ arising from the singular value decomposition.
F.6 Concrete Initializer Families
F.6.1 Gaussian Initializers
If the variational means are initialized as
$\mu^{A}_{ik},\mu^{B}_{jk}\sim\mathcal{N}(0,s^{2}),$
then $v_{A}=v_{B}=s^{2}$, and the symmetric variance-matching condition yields
$s=\left(\frac{\sigma_{W}^{2}}{r}\right)^{1/4}.$
Example: For He initialization with ReLU activations, $\sigma_{W}^{2}=2/n$, so:
$s=\left(\frac{2}{nr}\right)^{1/4}.$
F.6.2 Uniform Initializers
If instead
$\mu^{A}_{ik},\mu^{B}_{jk}\sim\mathcal{U}(-a,a),$
then by standard probability theory (Casella and Berger, 2002),
$\mathrm{Var}(\mathcal{U}(-a,a))=\frac{a^{2}}{3}.$
Substituting $v_{A}=v_{B}=a^{2}/3$ into the symmetric variance-matching condition (F.4):
$r\left(\frac{a^{2}}{3}\right)^{2}=\sigma_{W}^{2},$
which gives
$a=\sqrt{3}\left(\frac{\sigma_{W}^{2}}{r}\right)^{1/4}.$
Example: For Glorot initialization, $\sigma_{W}^{2}=2/(n+m)$, so:
$a=\sqrt{3}\left(\frac{2}{r(n+m)}\right)^{1/4}.$
F.7 Initialization of Variational Scales
The variational standard deviations $\sigma^{A}_{ik}$ and $\sigma^{B}_{jk}$ control posterior uncertainty and the magnitude of the KL divergence term in variational inference (Blundell et al., 2015).
To prevent stochastic sampling noise from dominating the likelihood signal at early stages of optimization, we initialize them as a fixed fraction of the mean scale:
$\sigma^{A}_{ik}=\sigma^{B}_{jk}=\eta\left(\frac{\sigma_{W}^{2}}{r}\right)^{1/4},\quad\eta\in(0,1).$
This choice ensures that the signal-to-noise ratio of $W$ at initialization remains $O(1)$, a condition commonly required for stable variational optimization (Graves, 2011).
Practical recommendation: A typical choice is $\eta=0.1$ or $\eta=0.01$, which provides moderate initial uncertainty without overwhelming the signal from the data.
F.8 Summary
The proposed initialization strategy:
  • Reproduces the second-order statistics of dense variance-preserving initializations (Glorot and Bengio, 2010; He et al., 2015),
  • Resolves scale non-identifiability inherent to low-rank factorizations,
  • Preserves computational and parametric efficiency,
  • Requires no pretrained full-rank weights.
It therefore provides a theoretically principled foundation for initializing low-rank variational neural network layers.
F.9 SVD Initialization from Pretrained Weights
Let $W^{*}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ denote the weight matrix of a converged deterministic model that minimizes the empirical risk. The variational objective for a low-rank Bayesian layer is
$\mathcal{L}_{\text{VI}}=\mathbb{E}_{q(W)}[\ell(W)]+\beta\,\text{KL}(q(W)\,\|\,p(W)),$ (32)
where $\ell(W)$ is the data loss and $\beta$ is the KL weight.
Since $W^{*}$ minimizes $\ell(W)$, initializing the posterior mean near $W^{*}$ places the optimization in a region where:
  • The expected data loss $\mathbb{E}_{q(W)}[\ell(W)]$ is already small;
  • The optimizer can focus on balancing the KL penalty and fine-tuning under the rank constraint.
This warm-start might, theoretically, reduces the risk of converging to poor local minima far from $W^{*}$.
To initialize a low-rank Bayesian layer from a pretrained deterministic weight matrix $W_{\text{full}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, we use the truncated singular value decomposition (SVD).
Let $W_{\text{full}}=U\Sigma V^{\top}$ be the SVD. The best rank-$r$ approximation is
$W_{r}=U_{r}\Sigma_{r}V_{r}^{\top},$ (33)
where $U_{r}\in\mathbb{R}^{d_{\text{out}}\times r}$, $V_{r}\in\mathbb{R}^{d_{\text{in}}\times r}$, and $\Sigma_{r}=\text{diag}(\sigma_{1},\ldots,\sigma_{r})$ contains the top $r$ singular values.
We initialize the mean parameters of the low-rank factorization $W=AB^{\top}$ by splitting the singular values symmetrically:
$\mu^{A}=U_{r}\Sigma_{r}^{1/2},\qquad\mu^{B}=V_{r}\Sigma_{r}^{1/2}.$ (34)
This ensures $\mu^{A}(\mu^{B})^{\top}=W_{r}$, warm-starting the Bayesian model from the optimal rank-$r$ approximation to the pretrained weights. Variance parameters are initialized using principled variance-preserving schemes.
Appendix G Experimental Details
This section provides complete implementation details to ensure full reproducibility of all experiments.
G.1 Hardware and Software
Compute Infrastructure:
  • GPU: NVIDIA H100 NVL with 95GB memory
  • CUDA Driver Version: 580.95.05
  • Operating System: Linux
  • Python: 3.10.12
Software Dependencies:
  • TensorFlow: 2.15.0
  • TensorFlow Probability: (bundled with TF 2.15.0)
  • NumPy: standard version compatible with TF 2.15.0
  • Scikit-learn: for preprocessing and metrics
  • Pandas: for data handling
  • DuckDB: for efficient preprocessing of large MIMIC-III tables
TensorFlow Configuration:
  • Precision: float32 (mixed precision explicitly disabled)
  • GPU memory: Growth mode enabled (prevents allocating all memory at startup)
  • Memory allocator: CUDA async allocator (TF_GPU_ALLOCATOR=cuda_malloc_async)
  • Deterministic operations: Enabled via TF_DETERMINISTIC_OPS=1 and TF_CUDNN_DETERMINISTIC=1
Random Seeds: All experiments run with 4 or 5 different random seeds. For each seed, we set: Python’s built-in random, NumPy’s random generator, and TensorFlow’s random operations. All reported metrics show mean $\pm$ standard deviation across these runs.
G.2 Evaluation Metrics
We report the following metrics depending on models and tasks:
Discrimination Performance:
  • AUROC (Area Under ROC Curve): Measures ability to rank positive samples higher than negative samples across all decision thresholds. Computed via scikit-learn’s roc_auc_score. We use AUROC instead of accuracy because accuracy is misleading for imbalanced datasets (e.g., predicting all negatives on MIMIC-III yields 91.6% accuracy but misses all deaths).
  • AUPR-Error: Area under precision-recall curve for detecting prediction errors. Higher values indicate better error detection via uncertainty.
  • AUPR-Success: Area under precision-recall curve for detecting correct predictions via low uncertainty.
Calibration Quality:
  • NLL (Negative Log-Likelihood): Measures quality of predicted probabilities. Lower is better. Computed as binary cross-entropy between true labels and predicted probabilities.
  • ECE (Expected Calibration Error): Measures alignment between predicted confidences and actual accuracy. We bin predictions by confidence and measure the gap between average confidence and accuracy in each bin. We test 6 binning configurations (equal-width vs. equal-mass binning with 10, 15, or 20 bins) and report the best (lowest) ECE for each model.
Out-of-Distribution Detection:
  • AUROC-OOD: AUROC for separating in-distribution from OOD samples using mutual information (MI) as the uncertainty score. MI measures epistemic (model) uncertainty and is computed from 512 forward passes with different weight samples.
  • AUPR-OOD: AUPR for OOD detection using MI.
  • MI Ratio: Mean MI on OOD data divided by mean MI on in-distribution data. Higher values indicate better OOD uncertainty.
All metrics computed using scikit-learn where applicable. For Bayesian models, predictions use 512 Monte Carlo weight samples.
G.3 MIMIC-III Mortality Prediction
G.3.1 Dataset and Preprocessing
Data source. MIMIC-III v1.4 (Johnson et al., 2016), binary ICU mortality prediction. Dataset sizes: 40,406 train (8.4% mortality), 4,490 test (8.6% mortality), 5,357 OOD newborn admissions (1.1% mortality).
Features. 44 continuous features: demographics (age capped at 90, gender), ICU characteristics (length of stay, time to admission), vital signs (heart rate, blood pressure, respiratory rate, temperature, SpO2), and laboratory results (CBC, metabolic panel, liver function, coagulation). Mean and standard deviation computed per ICU stay where applicable.
Preprocessing. Following Ruhe et al. (2019): (i) merge vital signs and labs from CHARTEVENTS/LABEVENTS, (ii) outlier removal via 8×IQR rule, (iii) median imputation, (iv) MinMax scaling to [0,1] fit on training set. Class weights: 1.0 (survived), 11.88 (deceased).
G.3.2 Models and Training
Architecture. All models use 2-layer MLP: 44 → 128 (ReLU) → 128 (ReLU) → 1 (sigmoid).
Model Parameters Initialization
Deterministic 20.5K Glorot uniform
Deep Ensemble (5×) 102.7K Glorot uniform, different seeds
Full-Rank BBB 41.1K $\mu\sim\mathcal{U}(-0.2,0.2)$, $\rho\sim\mathcal{U}(-5,-4)$
Low-Rank ($r=15$) 13.6K He/Glorot with damping
LR-SVD ($r=15$) 13.6K SVD of pretrained deterministic
Rank-1 Mult. 20.5K Glorot + small perturbations
Table 6: Model specifications. Low-rank damping: 0.32 for $r\leq 5$, 0.55 for $r>5$.
Rank choice We first do an ablation study to find the optimal combination of ranks across layers. If a deterministic network is available we can look at the singular values decay to gain intuition for the appropriate range or to validate the chosen rank obtained through ablation (See Figures 9 and 10).
Figure 9: Singular value decay and low-rank approximation error for deterministic neural network weight matrices. Layer 0 ($44\times 128$, left): Singular values decay rapidly (top-left, log scale). Red dashed line at $r=15$ indicates selected rank. Bottom-left shows energy retention curve: rank $r=15$ achieves $73.9\%$ of total Frobenius norm energy. Layer 1 ($128\times 128$, right): Similar rapid spectral decay (top-right). Red dashed line marks $r=15$. Bottom-right: cumulative energy curve reaches $74.6\%$ at $r=15$. Dotted horizontal line at $75\%$ energy (reference threshold).
Training. Adam optimizer ($\alpha=10^{-3}$, default $\beta$), batch size 64. Bayesian models: 256 epochs with ELBO objective using KL scaling $\lambda=0.5/N_{\text{batches}}\approx 7.9\times 10^{-4}$, scale-mixture prior $p(w)=0.5\mathcal{N}(0,1)+0.5\mathcal{N}(0,e^{-6})$ (Blundell et al., 2015), single weight sample per forward pass during training. Deterministic baseline: 32 epochs to prevent overfitting. Deep ensemble: 32 epochs per member with different seeds.
Inference. Bayesian models: 512 Monte Carlo samples for predictive mean and uncertainty. Deep ensemble: average over 5 members. Epistemic uncertainty via mutual information: $\mathbb{I}(y;\theta|x,\mathcal{D})=H[y|x,\mathcal{D}]-\mathbb{E}_{\theta}[H[y|x,\theta]]$.
G.3.3 Evaluation Metrics
Standard binary classification metrics computed on predictive mean: AUROC, AUPR. Uncertainty metrics: AUPR-Err (precision-recall for misclassification detection using predictive entropy), AUC/AUPR-OOD (detection of OOD samples using mutual information), AUPR-In (correctly identifying in-distribution samples). Calibration: negative log-likelihood (NLL), expected calibration error (ECE). All results averaged over 5 independent runs with different random seeds.
G.4 Beijing PM2.5 Time-Series Forecasting
G.4.1 Dataset and Preprocessing
Data source. Beijing PM2.5 air quality dataset, next-hour concentration forecasting. Dataset sizes: 29,213 train, 7,553 validation, 7,553 test. OOD set: Guangzhou PM2.5 data (distributional shift in pollution and meteorology).
Features. 15 continuous features per timestep over a T=24T{=}24 hour sliding window: PM2.5 concentration, wind direction/speed, temperature, pressure, cumulative snow/rain hours, and derived time encodings. Input shape: $(N,24,15)$.
Preprocessing. (i) Sliding-window extraction with horizon-1 targets; (ii) StandardScaler (zero mean, unit variance) fit on flattened training data only; (iii) targets independently standardized. All predictions inverse-transformed before evaluation.
G.4.2 Models and Training
Architecture. All models use a custom 2-layer LSTM (H=64H{=}64) built from scratch following Fortunato et al. (2017), with explicit time-step unrolling rather than tf.keras.layers.LSTM. Each LSTM layer contains two packed-gate projections (x_to_gates: $d_{\mathrm{in}}\to 4H$ with bias; h_to_gates: $H\to 4H$ without bias) computing all four gates (input, forget, cell candidate, output) in a single matrix multiply. Forget gate bias initialized to $1.0$ (Jozefowicz et al., 2015). Output: $\mathbf{h}_{T}$ from the top layer fed to a linear head producing a scalar forecast.
Model Layer Type Rank Initialization
Deterministic Dense Glorot (input), Orthogonal (recurrent)
Deep Ensemble ($5\times$) Dense Glorot/Orthogonal, different seeds
Full-Rank BBB DenseVariational $\mu\sim\mathcal{U}(-0.2,0.2)$, $\rho\sim\mathcal{U}(-3,-2)$
Low-Rank (r=[14,20]r{=}[14,20]) LowRankDenseVar. [14,20] Adaptive with damping (Sec. F)
LR-SVD r=[14,20]r{=}[14,20]) LowRankDenseVar. [14,20] SVD of pretrained deterministic
Rank-1 Mult. Rank1DenseVar. 1 Glorot base + small perturbations
Table 7: LSTM model specifications. Architecture: 2-layer LSTM (H=64H{=}64), input $(24,15)$, linear output head. Bayesian layers use $W=AB^{\top}$ (low-rank) or $(1{+}s)\odot xW_{0}\odot(1{+}r)$ (rank-1) (Dusenberry et al., 2020a). Low-rank output head uses rank=1{}=1.
Bayesian weight caching. Following Algorithm 2 of Fortunato et al. (2017), variational layers sample weights once at t=0t{=}0 (use_cached=False) and reuse them for $t{=}1,\ldots,T{-}1$ (use_cached=True). A callback clears caches at each batch boundary. KL loss is added only at t=0t{=}0.
Training. Adam optimizer ($\alpha=10^{-3}$, default $\beta$), batch size 64. Bayesian models: up to 150 epochs with ELBO objective using KL scaling $\lambda=0.09/N_{\mathrm{train}}\approx 3.08\times 10^{-6}$, scale-mixture prior $p(w)=0.5\mathcal{N}(0,1)+0.5\mathcal{N}(0,e^{-6})$ (Blundell et al., 2015), single weight sample per forward pass. Early stopping: patience 30 on validation MAE. Learning rate reduced by 0.5 on plateau (patience 5, min $10^{-6}$). Deep ensemble: 5 members, up to 150 epochs each (patience 10), seeds $42,\ldots,46$.
Inference. Bayesian models: 150 Monte Carlo samples for predictive mean and uncertainty. Deep ensemble: average over 5 members; noise variance pooled from validation residuals. Prediction intervals: $\hat{y}\pm z_{0.975}\hat{\sigma}$.
G.4.3 Evaluation Metrics
Point prediction: MAE, RMSE, $R^{2}$ in original scale. Uncertainty: NLL, CRPS, ECE, PICP (target 0.95), MPIW, interval score. OOD detection (Guangzhou): AUROC, AUPR, FPR@95%TPR, TPR@5%FPR using predictive std. Selective prediction at retention rates $\{100\%,95\%,\ldots,70\%\}$. All results averaged over runs.
G.5 SST-2 Sentiment Classification
G.5.1 Dataset and Preprocessing
Data source. SST‑2 for in‑distribution (ID) sentiment classification and AGNews for out‑of‑distribution (OOD) detection. SST‑2 is loaded from HuggingFace glue/sst2; AGNews is loaded from ag_news (test split).
Splits and sizes. The current preprocessed artifacts used in this experiment contain:
  • SST‑2 train: 67,349 examples
  • SST‑2 dev: 872 examples
  • AGNews test (OOD): 7,600 examples (4 classes, 1,900 each)
OOD is capped by OOD_SLICE = 7600 during evaluation.
Tokenization. The pipeline uses bert-base-uncased via AutoTokenizer. Each text is converted to a string (empty string for missing values), then tokenized with: add_special_tokens=True, max_length=64, padding=’max_length’, truncation=True, return_attention_mask=True, return_tensors=’np’.
Saved arrays. For each split, the following NumPy arrays are written: _ids.npy*\_ids.npy, _mask.npy*\_mask.npy, _labels.npy*\_labels.npy. The pipeline writes to SAVE_DIR_processed_data_agnews/. All arrays are loaded by the notebook through the config modules.
Labels. SST‑2 labels are binary. AGNews labels are 4‑class (0–3) and are stored but not used for classification; OOD detection treats AGNews as a single out‑of‑distribution set.
G.5.2 Models and Training
Architecture. All models share the same tiny pre‑norm Transformer encoder: $d_{\text{model}}=256$, $n_{\text{layers}}=4$, $n_{\text{heads}}=4$, $d_{\text{ff}}=512$, max length 64, CLS pooling, and 2‑class softmax. Dropout is disabled (rate 0.0).
Model Params Init (abbr.)
Det. 9.92M Glorot (Dense), Unif (Emb)
DE (5×) 49.61M Det. init, diff seeds
FR-BBB 19.84M $\mu\sim\mathcal{U}(-0.2,0.2)$, $\rho\sim\mathcal{U}(-5,-4)$
LR-BBB ($r{=}16,r_{e}{=}10$) 1.47M G/H adapt., damp=0.85; $\rho$ rand; $b:\mu{=}0,\rho{=}{-}5$
LR-SVD ($r{=}16,r_{e}{=}10$) 1.47M SVD init from Det.
R1-BBB 10.01M $W_{0}$ Glorot; $b,r,s:\mu\sim\mathcal{U}(-0.2,0.2),\rho\sim\mathcal{U}(-5,-4)$
Table 8: SST‑2 model specs. Abbr.: Det.=deterministic, DE=deep ensemble, FR=full‑rank, LR=low‑rank, R1=rank‑1, G/H=Glorot/He, Unif=uniform, rand=random.
Training. Batch size 64. Learning rate $2\times 10^{-4}$ with cosine decay (alpha =0.01=0.01) and weight decay $0.01$. Baseline uses Adam; Bayesian models use Adam to avoid decay on variational parameters. Epochs: baseline 20, Bayesian 20, deep ensemble 20 per member. Early stopping on validation loss (patience 5) and ReduceLROnPlateau (patience 3, factor 0.5, min LR $10^{-6}$).
KL scaling and annealing. KL weight is computed as $\lambda=0.001/N_{\text{batches}}$. Bayesian models use KL annealing with 2 zero‑KL epochs and linear warm‑up to full KL by epoch 4.
Seeds. Main experiment uses seed 42. Multi‑seed robustness uses {42, 123, 456, 2026}; all RNGs (Python, NumPy, TensorFlow) are set for each run.
G.5.3 Evaluation Metrics
Classification performance. Accuracy and AUROC on SST‑2 dev.
Calibration. Negative log‑likelihood (NLL) and Brier score. ECE is selected via a best‑configuration search over equal‑width vs. equal‑mass binning with 10/15/20 bins; the best configuration is then applied to all models.
Uncertainty and OOD detection. Bayesian models use Monte Carlo sampling with n_samples=200 (main experiment). Deep ensemble uncertainty is computed from member disagreement (variance/STD). Metrics include AUPR‑Error, AUPR‑Success, AUROC‑OOD, AUPR‑In/OOD, and MI ratio (mean MI on OOD divided by mean MI on ID). The evaluation reports both MI‑based and STD‑based OOD metrics.
Appendix H Extended Results and analysis from Main Experiments
H.1 Experimental Insights and Practical Considerations
Calibration–OOD Detection Tradeoff.
Our results reveal a consistent pattern: low-rank models achieve superior OOD detection while exhibiting moderately higher NLL compared to Deep Ensembles (e.g., MIMIC-III: 0.433 vs. 0.300; SST-2: 0.527 vs. 0.434). We hypothesize this tradeoff stems from the singular posterior geometry established in Section 3.1. Deep Ensembles maintain $M$ independent point estimates, each highly optimized for likelihood on training data, yielding sharp in-distribution predictions. In contrast, our low-rank posterior concentrates on the rank-$r$ manifold $\mathcal{R}_{r}$, enforcing structured weight correlations (Lemma 3.5) that propagate uncertainty more coherently across the input space. This enables better distinction between in-distribution and OOD inputs, precisely because the model maintains broader epistemic uncertainty rather than collapsing to overconfident predictions (see Figure 18).
This represents a tradeoff between predictive sharpness (rewarded by NLL) and epistemic honesty (rewarded by OOD detection). Notably, Low-Rank Gaussian achieves best-in-class AUPR-In (0.824 on MIMIC-III), suggesting its uncertainty estimates remain reliable for identifying trustworthy predictions. For safety-critical applications where knowing what you don’t know matters more than marginal log-likelihood improvements, this tradeoff favors low-rank approaches.
Ensembling as a potential mitigation. The calibration gap may be addressed by ensembling multiple low-rank models. Given the parameter efficiency of our approach (13.6K vs. 112K for Deep Ensemble on MIMIC-III), one could ensemble 5–8 low-rank models while remaining below the parameter budget of a single Deep Ensemble, potentially recovering calibration benefits while preserving superior OOD detection. We leave systematic exploration of low-rank ensembles to future work.
SVD Initialization.
Theoretically, warm-starting from the truncated SVD of a learned deterministic weight matrix should improve optimization by initializing near a good solution (Appendix F.9). Empirically, LR-SVD does achieve gains on specific metrics, higher AUROC on MIMIC-III (0.898 vs 0.895) and best AUPR-Succ on SST-2 (0.923 vs 0.917), but these improvements are modest, and performance on other metrics shows no benefit or slight degradation. Thus, SVD initialization is justified only when a pretrained deterministic weight matrix already exists or when training one is inexpensive. Otherwise, the improvements do not warrant the additional computational burden of first training a deterministic model.
Computational Efficiency and Scaling.
Parameter reduction does not uniformly translate to wall-clock speedups. On our smaller-scale experiments (MIMIC-III, Beijing Air Quality), where architectures contain tens of thousands of parameters, we observe minimal training time differences, particularly on GPU. This reflects a fundamental characteristic of GPU computation: modern accelerators are optimized for large matrix multiplications that saturate parallel compute units. Two smaller matrix multiplications ($A\in\mathbb{R}^{m\times r}$ times $B^{\top}\in\mathbb{R}^{r\times n}$) incur overhead from multiple kernel launches and underutilize GPU parallelism compared to a single large multiplication, even when the total FLOP count is lower. Additionally, the variational formulation requires sampling two factor matrices per layer, adding memory transactions that dominate compute time at small scales.
The efficiency benefits manifest at larger scales. On SST-2, our low-rank Transformer trains in 8.2 minutes compared to 23.1 minutes for Full-Rank BBB, approaching the 7.7-minute baseline of the deterministic model. This is achieved with 1.5M parameters versus 19.8M for Full-Rank BBB ($13\times$ reduction). We view this as evidence that low-rank variational inference scales favorably: as architectures grow, the $O(r(m+n))$ versus $O(mn)$ parameter complexity translates increasingly into practical speedups, suggesting even greater benefits for large-scale models where full-rank BNNs become prohibitive.
Role of KL Scaling and Annealing.
The KL weight $\beta$ in the ELBO objective $\mathcal{L}=\mathbb{E}_{q}[\log p(\mathcal{D}|W)]-\beta\cdot\mathrm{KL}(q\|p)$ controls the strength of posterior regularization toward the prior, and its setting significantly impacts all evaluation metrics.
Effect on predictive performance and calibration. Higher $\beta$ values (stronger regularization) constrain the posterior to remain closer to the prior, which may limit in-distribution accuracy but can yield better-calibrated uncertainty estimates. Lower $\beta$ values allow the posterior to be dominated by the likelihood, typically concentrating around likelihood-maximizing weights, improving accuracy but risking overconfidence.
Effect on OOD detection. Stronger KL regularization (higher $\beta$) generally improves OOD detection by preventing posterior collapse to near-deterministic weights. When the posterior retains meaningful variance, epistemic uncertainty can distinguish familiar from unfamiliar inputs. Conversely, very low $\beta$ can yield posteriors so concentrated that mutual information approaches zero, eliminating the model’s ability to express epistemic uncertainty.
We employ KL annealing, gradually increasing $\beta$ from zero over early epochs, to balance these considerations. This allows the model to first find a good likelihood basin before the regularization penalty takes effect, avoiding the underfitting that can occur when strong KL penalties are applied from initialization. Across experiments, we found this schedule crucial for stable training, particularly in deeper architectures (Transformers, LSTMs) where early KL dominance can prevent learning altogether. Our $\beta$ values (ranging from $0.001/N_{\text{batches}}$ to $0.5/N_{\text{batches}}$ depending on architecture) were selected via validation performance, balancing the accuracy–uncertainty tradeoff for each task. As with any regularization hyperparameter, $\beta$ requires tuning for each architecture and dataset; we do not claim our chosen values are optimal.
Practical Rank Selection.
The rank $r$ can be treated as a standard hyperparameter tuned via ablation. We recommend running ablation studies with reduced computational budget, fewer training epochs and fewer MC samples during validation, to efficiently identify a suitable rank range. In our experiments, this direct approach was our primary method. The singular value decay analysis (Figures 7915) serves as an optional validation step that can narrow the search range when a pretrained deterministic model happens to be available, but it is not required. Practitioners without the time or compute to first train a deterministic network can proceed directly to rank ablation. As Figure 10 illustrates, rank selection meaningfully impacts the calibration–OOD detection tradeoff, making this ablation step important for achieving desired performance characteristics.
Statistical Considerations.
While mean improvements for low-rank models are consistent in direction across random seeds, we note that some confidence intervals overlap between methods (e.g., Table 1 AUPR-OOD). The observed patterns, particularly the calibration–OOD tradeoff, replicate across all four experimental settings, lending confidence to the qualitative conclusions. However, larger-scale studies with more seeds would be needed to establish statistical significance for smaller effect sizes.
H.2 MIMIC-III
Rank Selection via Ablation Study
Figure 10: Trade-off between OOD detection (AUROC OOD MI) and calibration (NLL, ECE) for low-rank Gaussian models with varying rank pairs $(r_{1},r_{2})$.
Rank selection is performed through ablation studies with reduced computational budget (fewer epochs, fewer MC samples). Figure 10 shows the trade-off between OOD detection (AUROC, y-axis) and calibration (NLL/ECE, x-axis) for different rank configurations on MIMIC-III. Higher ranks generally improve OOD detection but degrade calibration, while lower ranks maintain better calibration at reduced OOD sensitivity. The configuration $(r_{1}=15,r_{2}=15)$ occupies the Pareto frontier, achieving AUROC $\approx 0.72$ while maintaining NLL $\approx 0.33$ and ECE $\approx 0.12$ (vertical dashed lines mark baseline thresholds). When pretrained weights are available, singular value decay analysis (Figure 9) can optionally guide rank selection by revealing the effective dimensionality of weight matrices. The characteristic exponential decay pattern shows that weight matrices naturally reside in lower-dimensional subspaces, providing post-hoc justification for the empirically selected ranks.
Figure 11: OOD detection performance on MIMIC-III. Low-Rank Gaussian ($r=15$) achieves AUPR OOD of 0.81, outperforming Full-Rank BNN (0.75) and Deep Ensemble (0.77) while using $\sim 30\%$ of full rank parameters and 12% of deep ensemble parameters
Table 9: Training time comparison on MIMIC-III (2-layer MLP). Despite 70% parameter reduction, low-rank methods incur overhead from dual sampling and matrix multiplications.
Method Training Time (s)
Deterministic Dense 52.5
Deep Ensemble (5 models) 221.1
Full-Rank BBB 502.4
Rank-1 Multiplicative 503.2
Low-Rank Gaussian SVD init 524.0
Low-Rank Gaussian ($r=15$) 532.2
Computational Efficiency at Small Scale
Table 9 shows training times for MIMIC-III experiments. Despite achieving 70% parameter reduction (from 44.8K to 13.6K parameters), low-rank methods exhibit training times comparable to Full-Rank BBB (532s vs. 502s). This stems from the computational overhead of performing two sampling operations and two matrix multiplications ($\mathbf{A}\in\mathbb{R}^{m\times r}$ times $\mathbf{B}^{\top}\in\mathbb{R}^{r\times n}$) rather than a single operation. Modern GPUs achieve higher throughput with one large matrix multiplication dispatched to a single kernel, whereas the factored form requires two kernel launches with intermediate memory transfers. Consequently, at small to medium scales, parameter reduction translates primarily to memory savings rather than wall-clock speedups. However, the reduced memory footprint enables deployment on resource-constrained devices and facilitates cheaper ensembling strategies. As we demonstrate in Section G.5, efficiency benefits emerge at larger scales where the parameter reduction outweighs the factorization overhead.
Parameter-Efficiency Adjusted Performance
Figure 12: Parameter-efficiency adjusted performance across models. Metrics normalized to $[0,1]$ and scaled by $\sqrt{\text{min\_params}/\text{params}}$ to reward computational efficiency. Low-Rank Gaussian $(r=15)$ achieves the best efficiency-performance trade-off, particularly excelling in AUPR metrics while maintaining compact parameterization.
Raw performance comparisons can mislead when models have different parameter scales. Figure 12 applies efficiency adjustment using $\sqrt{\text{min\_params}/\text{params}}$, penalizing larger models while preventing complete suppression via square-root dampening . Low-Rank Gaussian $(r=15)$ emerges as the winner, achieving maximal coverage with particular strength in AUPR In Domain MI and AUPR OOD MI . This suggests low-rank parameterization captures essential uncertainty structure without requiring full parameter space .
Full-Rank BBB suffers notable reduction despite good raw performance due to $\sim$3× parameter count . Deep Ensemble show the biggest drop(efficiency factors $\approx 0.14$). The Rank-1 Multiplicative model demonstrates moderate efficiency-adjusted performance, balancing extreme compactness with reasonable metrics. This analysis supports low-rank architectures for resource-constrained deployment requiring both performance and efficiency .
H.3 Beijing PM2.5 time-series forecasting.
Rank Selection via Ablation Study
Figure 13: Trade-off between OOD detection (AUROC, y-axis) and calibration metrics (NLL and ECE, x-axis) for low-rank LSTM models with varying rank pairs $(r_{\text{ih}},r_{\text{hh}})$ on Beijing air quality forecasting. Points labeled with rank configurations; color indicates KL weight. Dashed lines mark baseline thresholds.
Figure 14: Trade-off between OOD detection (AUROC, y-axis) and predictive performance (MAE and RMSE, x-axis) for low-rank LSTM models with varying rank configurations on Beijing air quality forecasting. Lower MAE/RMSE indicates better forecasting accuracy.
Rank selection is performed through ablation studies with reduced computational budget (fewer epochs, fewer MC samples). Figures 13 and 14 show the multi-dimensional trade-offs between OOD detection (AUROC, y-axis) and both calibration (NLL/ECE) and predictive performance (MAE/RMSE) for different rank configurations $(r_{\text{ih}},r_{\text{hh}})$ on the Beijing dataset. Higher ranks generally improve OOD detection but can degrade both calibration and forecasting accuracy. The configuration $(r_{\text{ih}}=14,r_{\text{hh}}=20)$ achieves strong OOD detection (AUROC $\approx 0.66$) while maintaining competitive calibration and predictive performance (vertical dashed lines mark baseline thresholds). The KL weight $\beta=0.25$ provides better separation of the Pareto frontier compared to lower values. When pretrained weights are available, singular value decay analysis (Figure 15) can optionally guide rank selection by revealing the effective dimensionality of weight matrices, providing post-hoc justification for the empirically selected ranks.
Figure 15: Singular value decay of pretrained LSTM weight matrices. we can observe fast decay for both x-to-gate and h-to-gate layers
Calibration Analysis
Figure 16 presents the calibration curves for all uncertainty-aware models, with corresponding AUC values in Table 10. The Low-Rank Bayesian model achieves the best calibration (AUC = 0.104), closely followed by Full-Rank BBB (AUC = 0.110), indicating that both models produce well-calibrated predictive intervals that closely match their nominal coverage levels. The SVD-initialized variant shows slightly degraded calibration (AUC = 0.126), suggesting a trade-off between OOD detection capability and calibration quality. Both Rank-1 and Deep Ensemble exhibit substantially worse calibration (AUC $\approx$ 0.30), with curves deviating significantly from the diagonal, indicating overconfident predictions where observed coverage falls below expected coverage across most confidence levels.
Figure 16: Calibration curves for Bayesian LSTM variants. The dashed diagonal represents perfect calibration. Models closer to the diagonal exhibit better-calibrated uncertainty estimates.
Table 10: Calibration Error (AUC) comparison. Best bold, second underlined.
[…]
Full-Rank BBB 0.110
Low-Rank (r=r{=}14,20) 0.104
Low-Rank (SVD) 0.126
Rank-1 Mult. 0.302
Deep Ensemble (5) 0.307
Selective Prediction Analysis
Selective prediction, where models abstain from predictions on uncertain inputs, is critical for safety-critical applications. We evaluate how well each model’s uncertainty estimates enable effective selective prediction by progressively filtering samples with highest predictive variance.
Table 11 presents performance at three retention levels. At 100% retention, Deep Ensemble achieves best MAE (10.40), reflecting its superior calibration. However, the key insight emerges when filtering uncertain predictions: Bayesian methods demonstrate substantially larger improvements, indicating their uncertainty estimates better identify difficult samples.
At 80% retention, Low-Rank achieves 8.71 MAE (17.4% improvement from baseline), outperforming Deep Ensemble’s 9.21 MAE (11.4% improvement). Notably, Low-Rank’s absolute performance at 80% retention (8.71) surpasses Deep Ensemble’s baseline performance (10.40), demonstrating that Bayesian uncertainty can compensate for modest calibration gaps through effective abstention.
The pattern is consistent across retention levels: Bayesian methods show 10-11% improvement at 90% retention compared to Deep Ensemble’s 9%, and maintain this advantage at 80%. This suggests the rank constraint induces uncertainty estimates that correlate more strongly with prediction difficulty, even when those uncertainties are less well-calibrated in absolute terms.
Table 11: Selective prediction performance on Beijing Air Quality test set. Values show MAE (lower better) and improvement percentage from 100% retention baseline. Best performance at each retention level in bold.
Model 100% Retention 90% Retention 80% Retention
Full-Rank BBB 10.66 (0.0%) 9.46 (11.2%) 8.82 (17.2%)
Low-Rank BBB 10.55 (0.0%) 9.43 (10.6%) 8.71 (17.4%)
Low-Rank SVD 10.77 (0.0%) 9.50 (11.8%) 9.06 (15.8%)
Rank-1 BBB 10.80 (0.0%) 9.45 (12.5%) 9.42 (12.8%)
Deep Ensemble 10.40 (0.0%) 9.45 (9.1%) 9.21 (11.4%)
Implications for Practice. These results have important implications for deployment scenarios:
  • In likelihood-focused settings requiring predictions on all inputs, Deep Ensemble remains preferable due to superior calibration.
  • In safety-critical settings where abstention is acceptable, Low-Rank methods provide better risk-return tradeoffs, achieving lower error rates on retained predictions.
  • The crossover point occurs around 85-90% retention, suggesting practitioners can use retention rate requirements to guide method selection.
This selective prediction advantage helps explain why Low-Rank methods achieve competitive OOD detection despite calibration gaps: the uncertainty signal effectively identifies distributional shift, even if absolute confidence levels are miscalibrated.
H.4 SST-2 sentiment classification
Computation efficiency
Table 12 reports training times for each model on SST-2. Low-Rank BBB variants train in approximately 8 minutes, approaching the 7.7-minute baseline of the deterministic Transformer despite performing variational inference with Monte Carlo sampling. This represents a $2.9\times$ speedup over Full-Rank BBB (23.08 min) and $8\times$ over Deep Ensemble (64.72 min).
The efficiency gain stems directly from parameter reduction: Low-Rank BBB uses 1.5M parameters versus 19.8M for Full-Rank BBB and 49.6M for Deep Ensemble. At this scale, the $O(r(m+n))$ versus $O(mn)$ complexity difference translates into meaningful wall-clock improvements, validating our theoretical analysis. Rank-1 BBB occupies a middle ground (12.52 min), reflecting its intermediate parameterization.
These results demonstrate that low-rank factorization enables practical Bayesian uncertainty quantification at training costs comparable to deterministic baselines, a key requirement for adoption in resource-constrained settings.
Table 12: Average Training Time per Model (in minutes)
Model Avg. Time (min)
Baseline Transformer 7.70
Low-Rank SVD Init 7.95
Low-Rank Random Init 8.22
Rank-1 BBB 12.52
Full-Rank BBB 23.08
Deep Ensemble 64.72
Parameter-Efficiency Adjusted Performance
Figure 17 presents metrics adjusted by parameter efficiency.
Low-Rank BBB (red) achieves the largest coverage area, dominating across AUROC-OOD, AUPR-Success, AUPR-In, and MI Ratio while maintaining competitive accuracy and NLL. Low-Rank SVD (purple) shows a similar profile with slightly reduced OOD metrics. Both low-rank variants substantially outperform all other approaches on efficiency-adjusted scores.
Deep Ensemble (green) and Full-Rank BBB (orange) collapse toward the center despite strong raw performance, reflecting their $33\times$ and $13\times$ parameter overhead respectively. Rank-1 BBB (brown) shows moderate efficiency-adjusted performance but remains dominated by the low-rank variants across most metrics. The deterministic baseline (blue) scores well only on parameter efficiency, with no uncertainty quantification capability.
[…]
Figure 17: Parameter-efficiency adjusted performance across models. Metrics normalized to $[0,1]$ and scaled by $\sqrt{\text{param\_efficiency}}$. Low-Rank BBB (red) achieves the best overall tradeoff, dominating OOD detection and uncertainty metrics while maintaining competitive accuracy.
Uncertainty Distribution Analysis
Figures 18 and 19 visualize the distribution of predictive uncertainty for in-distribution (SST-2) versus out-of-distribution (AGNews) samples across models. Several patterns emerge.
Low-Rank BBB exhibits clear separation between in-domain and OOD distributions for both predictive standard deviation and mutual information. The OOD samples (orange) concentrate at higher uncertainty values, enabling effective detection.
Deep Ensemble shows modest separation, with a substantial distribution overlap between OOD samples and ID samples.
Full-Rank BBB displays considerable overlap between distributions, explaining its weaker OOD detection performance. The model struggles to assign higher uncertainty to unfamiliar inputs.
Rank-1 BBB exhibits near-degenerate uncertainty distributions with values concentrated near zero, consistent with its low MI ratio (Table 3). This confirms that rank-1 perturbations provide insufficient epistemic expressiveness.
Low-Rank SVD BBB shows similar separation patterns to Low-Rank BBB with random initialization, though with slightly narrower distributions overall.
Notably, predictive standard deviation provides marginally better OOD separation than mutual information for most Bayesian models. Table 13 confirms this empirically: STD-based metrics slightly outperform MI-based metrics across all models, though differences are small. This suggests that for these architectures, total predictive variance captures OOD signal as effectively as the epistemic component isolated by mutual information.
Table 13: OOD detection metrics using Mutual Information (MI) vs Predictive Standard Deviation (STD) as uncertainty measures. Results averaged over 4 seeds. Higher is better for all metrics.
AUROC-OOD AUPR-OOD AUPR-In
Model MI STD MI STD MI STD
Deterministic .500 .500 .897 .897 .103 .103
Deep Ensemble .658 .663 .930 .932 .267 .282
Full-Rank BBB .622 .623 .921 .921 .222 .223
Low-Rank BBB .641 .642 .919 .920 .302 .303
LR-SVD BBB .616 .619 .903 .904 .273 .274
Rank-1 BBB .613 .615 .904 .905 .273 .273
Figure 18: Predictive standard deviation distributions for in-distribution (blue) vs OOD (orange) samples. Low-Rank BBB and Deep Ensemble show clear separation; Rank-1 BBB shows near-degenerate distributions.
Figure 19: Mutual information distributions for in-distribution (blue) vs OOD (orange) samples. Similar patterns to STD, with Deep Ensemble showing strongest separation.
Appendix I Additional Experimental Results
This section presents additional experiments and extended results that complement the main paper findings.
I.1 MNIST Experiment
We conduct a controlled empirical study comparing several Bayesian neural network (BNN) parameterizations on MNIST classification task. This experiment serves as a reproduction of the foundational Bayes by Backprop (BBB) framework from  (Blundell et al., 2015), with the addition of modern low-rank factorization approaches and a rank-1 multiplicative variant. (Dusenberry et al., 2020a).
I.1.1 Experimental Setup
Dataset and Preprocessing.
We use the full MNIST dataset (LeCun et al., 1998) with 60,000 training samples and 10,000 test samples. Following  (Blundell et al., 2015), pixel values are normalized to the range $[0,1]$ by dividing by 126, and each $28\times 28$ image is flattened to a 784-dimensional feature vector. The training set is further split into 50,000 for training and 10,000 for validation; the original test set remains held out. No data augmentation is applied.
Neural Network Architecture.
All models employ an identical MLP architecture:
$784\to 1200\to 1200\to 10,$
where the first two hidden layers use ReLU activations and the output layer is linear (logits). This setup matches the standard Blundell et al. baseline to ensure fair comparison.
Models Under Comparison.
We evaluate four distinct Bayesian parameterizations, all using the same prior and KL scaling:
  1. 1.
    Full-Rank Gaussian (BBB). Following (Blundell et al., 2015), we place an independent diagonal Gaussian variational posterior on each weight and bias:
    $q(W_{ij}\mid\mu_{ij},\sigma_{ij}^{2})=\mathcal{N}(\mu_{ij},\sigma_{ij}^{2}),$
    reparameterized as $W=\mu+\sigma\odot\epsilon$ with $\epsilon\sim\mathcal{N}(0,I)$. Biases are either stochastic Gaussians or deterministic (both tested); we report the deterministic-bias variant for parameter efficiency.
  2. 2.
    Low-Rank Gaussian Factorization. We impose a rank-$r$ factorization $W\approx AB^{\top}$ where $A\in\mathbb{R}^{n\times r}$ and $B\in\mathbb{R}^{m\times r}$ with $r\ll\min(n,m)$. Each entry of $A$ and $B$ has a diagonal Gaussian posterior: $q(A_{ik})=\mathcal{N}(\mu^{A}_{ik},(\sigma^{A}_{ik})^{2})$ and similarly for $B$. The bias remains deterministic. We sweep ranks $r\in\{10,25,50\}$ for hidden layers and report the rank that best balances accuracy and calibration.
  3. 3.
    Low-Rank Laplace Factorization. Identical to the Gaussian factorization but with Laplace posteriors on $A$ and $B$:
    $q(A_{ik}\mid\mu^{A}_{ik},b^{A}_{ik})=\text{Laplace}(\mu^{A}_{ik},b^{A}_{ik}),$
    where the scale parameter is $b=\text{softplus}(\rho)$ to ensure positivity. Laplace posteriors have heavier tails than Gaussians, potentially improving uncertainty calibration in low-rank settings.
  4. 4.
    Rank-1 Multiplicative (Dusenberry et al., 2020). Following Dusenberry et al. (Dusenberry et al., 2020a), we fix a deterministic base weight matrix $\overline{W}$ and apply a rank-1 stochastic multiplicative perturbation:
    $W^{\prime}=\overline{W}\odot(s\otimes r),$
    where $s\in\mathbb{R}^{n}$ and $r\in\mathbb{R}^{m}$ are stochastic vectors with diagonal Gaussian posteriors. This parameterization dramatically reduces the stochastic parameter count compared to full-rank while maintaining expressiveness through the deterministic base.
Shared Prior.
All stochastic parameters (weights, factors, or vectors) use the same scale-mixture Gaussian prior:
$p(w)=\pi\mathcal{N}(0,\sigma_{1}^{2})+(1-\pi)\mathcal{N}(0,\sigma_{2}^{2}),$
with fixed hyperparameters $\pi=0.5$, $\sigma_{1}=1.0$, and $\sigma_{2}=\exp(-6)\approx 0.00248$. This prior encourages sparsity by placing probability mass on both a standard Gaussian and a narrow Gaussian; it is identical across all four models to ensure fair comparison.
Training Procedure.
The models are trained by minimizing the ELBO:
$\mathcal{L}=\mathbb{E}_{q(\theta)}[-\log p(y\mid x,\theta)]+\text{kl\_scale}\cdot\text{KL}(q(\theta)\|p(\theta)),$
where $\text{kl\_scale}=1/N_{\text{train}}$ (i.e., the reciprocal of the training set size). To avoid posterior collapse due to a large KL term early in training, we employ a linear warm-up schedule: kl_scale is ramped from 0 to $1/N_{\text{train}}$ over the first 20 epochs, then held constant.
Training uses Adam optimizer with learning rate $10^{-3}$, batch size 128, and 50 epochs. All random seeds are fixed at 42 for reproducibility. Gradient computation is deterministic (cuDNN disabled) to ensure consistent behavior across runs.
Evaluation Protocol.
At test time, we estimate predictive uncertainty and calibration by drawing $T=50$ stochastic forward passes from the learned variational posterior. From these samples, we compute:
  • Accuracy: Classification accuracy using the mean predicted class.
  • Negative Log-Likelihood (NLL): $\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\log\hat{p}(y_{i}\mid x_{i})$, where $\hat{p}$ is the mean predictive distribution.
  • Brier Score: $\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\left\|\hat{p}_{i}-y_{i}^{\text{one-hot}}\right\|_{2}^{2}$, measuring squared prediction error on probabilities.
  • Expected Calibration Error (ECE): Binning predictions into equal-mass or equal-width bins and computing the average gap between confidence and empirical accuracy.
  • Mutual Information (MI): $H(\bar{p})-\mathbb{E}_{t\in[T]}[H(p^{(t)})]$, where $\bar{p}$ is the mean predictive distribution and $H$ is Shannon entropy. MI quantifies the epistemic uncertainty contributed by weight sample variation.
I.1.2 Parameter Counts and Computational Complexity
The four models exhibit different parameter efficiency:
Model Hidden Rank(s) Output Rank Total Params Compression
[…]
Table 14: Parameter counts for the four BNN variants on MNIST. Full-Rank and Rank-1 parameters count both means and rho (via softplus) scales; low-rank also includes deterministic biases. Compression factors are relative to Full-Rank BBB.
The dramatic reduction in low-rank variants arises because low-rank replaces $O(nm)$ parameters (for a full weight matrix of size $n\times m$) with $O(r(n+m))$ parameters per layer. For the 1200-unit hidden layers and rank $r=25$, this yields approximately $25(1200+1200)=60,000$ parameters per layer versus $2\times 1,200^{2}\approx 2.88$ million for the full-rank Gaussian.
Forward-pass complexity is similarly reduced: full-rank requires $O(bnm)$ multiplications for batch size $b$, while low-rank requires $O(b(nr+rm))=O(br(n+m))$. For $r=25$, this is a speedup of roughly $2r\approx 50\times$ for the matrix multiplication, though the full pipeline (sampling factors, bias, activation) has a smaller overall speedup.
I.1.3 Results
Test Set Performance.
Model Accuracy NLL Brier ECE MI
Full-Rank BBB $98.15\%$ $0.0795$ $0.0048$ $0.0082$ $0.1243$
Low-Rank Gaussian $97.32\%$ $0.0924$ $0.0059$ $0.0104$ $0.0956$
Low-Rank Laplace $97.21\%$ $0.0607$ $0.0043$ $0.0117$ $0.0823$
Rank-1 Multiplicative $98.21\%$ $0.1156$ $0.0071$ $0.0187$ $0.2104$
[…]
  • Calibration: Full-Rank BBB achieves the lowest ECE ($0.0082$), indicating superior alignment between predicted confidence and empirical accuracy. Low-Rank Gaussian and Laplace variants show slightly degraded calibration (ECE $\approx 0.01$$0.012$), while Rank-1 exhibits notably poorer calibration (ECE $0.0187$).
  • NLL: Low-Rank Laplace achieves the best NLL ($0.0607$), suggesting that heavier tails improve likelihood estimation. However, this does not translate to better ECE, highlighting a subtle trade-off between likelihood and calibration.
  • Accuracy: All models achieve high accuracy (97–98%). Rank-1 achieves the highest ($98.21\%$), but this comes at the cost of calibration.
  • Epistemic Uncertainty: Full-Rank BBB exhibits the highest MI ($0.1243$), indicating stronger epistemic uncertainty from posterior variation. Low-rank variants show more conservative epistemic estimates (MI $\approx 0.08$$0.10$), consistent with reduced posterior expressiveness.
Calibration stability.
To verify that calibration differences are genuine and not artifacts of the ECE binning scheme, we compute ECE under four configurations: equal-mass and equal-width binning with $B\in\{10,15,30\}$ bins.
Model EM, B=15 EW, B=15 EM, B=10 EW, B=30
Full-Rank BBB $0.0082$ $0.0088$ $0.0075$ $0.0095$
Low-Rank Gaussian $0.0104$ $0.0118$ $0.0096$ $0.0126$
Low-Rank Laplace $0.0117$ $0.0131$ $0.0108$ $0.0144$
Rank-1 Multiplicative $0.0187$ $0.0201$ $0.0172$ $0.0219$
Table 16: ECE under different binning schemes. EM = equal-mass, EW = equal-width. Across all binning choices, Full-Rank BBB maintains the lowest ECE, and the ranking of models is stable, confirming that calibration differences are not artifacts of a single binning scheme.
I.1.4 Interpretation and Trade-offs
[…]
Low-rank variants exhibit lower mutual information (MI $\approx 0.08$$0.10$ vs. $0.1243$), reflecting a more focused posterior that still produces well-calibrated uncertainty estimates. This suggests the constrained parameterization enforces a useful inductive bias. In contrast, Rank-1 multiplicative, despite achieving the highest accuracy, exhibits significantly degraded calibration (ECE $0.0187$), indicating that low-rank factorization offers superior expressiveness-to-parameter ratio.
Low-Rank Laplace achieves the best NLL ($0.0607$) but slightly higher ECE than Gaussian ($0.0117$ vs. $0.0104$), reflecting a minor trade-off between likelihood and calibration.
[…]
I.2 Fashion-MNIST Experiment
To assess generalization beyond MNIST, we conduct an identical experiment on Fashion-MNIST (Xiao et al., 2017), which presents a more challenging classification task with the same architecture and training protocol as the MNIST study.
I.2.1 Experimental Configuration
Fashion-MNIST is a 10-class image classification dataset with 60,000 training and 10,000 test samples of $28\times 28$ grayscale fashion product images. All preprocessing, architecture, prior, and training configuration are identical to the MNIST experiment: MLP $784\to 1200\to 1200\to 10$, pixel normalization by 126, same scale-mixture Gaussian prior, and identical optimizer and KL warm-up schedule. Only the dataset differs, providing a natural test of models on harder classification problems.
I.2.2 Results
Model Accuracy NLL Brier ECE MI
Full-Rank BBB $89.21\%$ $0.3583$ $0.1591$ $0.0262$ $0.0317$
Low-Rank Gaussian $88.69\%$ $0.3186$ $0.1623$ $\mathbf{0.0114}$ $0.0354$
Low-Rank Laplace $87.70\%$ $0.3421$ $0.1755$ $0.0135$ $0.0404$
Rank-1 Multiplicative $89.07\%$ $0.6928$ $0.1846$ $0.0794$ $0.0038$
Table 17: Fashion-MNIST test-set metrics (MC-50). Low-Rank Gaussian achieves the best calibration (ECE, NLL) with only 0.5% accuracy loss versus Full-Rank. Rank-1 remains poorly calibrated despite high accuracy.
Fashion-MNIST results closely mirror MNIST findings, confirming the generalizability of our conclusions:
  • Low-Rank Gaussian shows good calibration: Achieves the lowest NLL ($0.3186$) and ECE ($0.0114$), validating the MNIST observation. The accuracy penalty is minimal ($88.69\%$ vs $89.21\%$ full-rank), 0.52 percentage points difference.
  • Full-Rank BBB trades calibration for accuracy: While achieving top accuracy ($89.21\%$), its NLL ($0.3583$) and ECE ($0.0262$) are slightly under the low-rank models
  • Rank-1 miscalibration persists: Despite near-optimal accuracy ($89.07\%$), Rank-1 exhibits signs of miscalibration on this dataset(NLL $0.6928$, ECE $0.0794$) and very small epistemic uncertainty (MI $0.0038$).
  • Laplace slightly worse than Gaussian: While Laplace achieved competitive NLL on MNIST, it shows worse calibration here (ECE $0.0135$ vs $0.0114$), though it maintains higher epistemic uncertainty (MI $0.0404$).
I.2.3 Calibration stability Across Binning Schemes
To confirm that calibration differences are not artifacts of binning methodology, we compute ECE under equal-mass and equal-width schemes with $B\in\{10,15,30\}$ bins.
Model EM, B=15 EW, B=15 EM, B=10 EW, B=30
Low-Rank Gaussian $\mathbf{0.0114}$ $0.0120$ $0.0113$ $0.0139$
Low-Rank Laplace $0.0135$ $0.0135$ $0.0130$ $0.0143$
Full-Rank BBB $0.0262$ $0.0252$ $0.0243$ $0.0291$
Rank-1 Multiplicative $0.0794$ $0.0795$ $0.0794$ $0.0796$
Table 18: ECE is stable across binning schemes on Fashion-MNIST, confirming that calibration rankings are robust. Low-Rank Gaussian consistently outperforms other approaches.
I.2.4 Cross-Dataset Consistency
Comparing MNIST and Fashion-MNIST results:
  • Model rankings are consistent across datasets: The relative calibration ordering observed on MNIST is reflected on Fashion-MNIST, suggesting the patterns are not dataset-specific. However, absolute performance gaps vary, reflecting differences in task difficulty and dataset characteristics.
  • Parameter efficiency gap remains: Low-rank models maintain their 19.5× parameter advantage while staying competitive on accuracy and superior on calibration metrics.
  • Epistemic uncertainty patterns persist: Low-rank variants show conservative but well-calibrated MI; Rank-1 remains near-deterministic.
These results strengthen our main paper’s claims: low-rank Bayesian neural networks offer a practical efficiency-quality trade-off that seems to generalize across datasets. The Fashion-MNIST experiment demonstrates that the benefits of low-rank factorization extend to more challenging classification tasks with the same architectural family.
I.3 Toy Regression: Bayesian Uncertainty Quantification
We reproduce the toy regression task experiment from (Blundell et al., 2015). We compare full-rank and low-rank Bayesian Neural Networks (BNNs) on a non-linear regression task following, evaluating how low-rank factorization preserves epistemic uncertainty while reducing parameters.
I.3.1 Experimental Setup
Data Generation.
We generate $N_{\text{train}}=1024$ training samples from a non-linear function with observation noise:
$y=x+0.3\sin(2\pi(x+\varepsilon))+0.3\sin(4\pi(x+\varepsilon))+\varepsilon,$ (35)
where $x_{\text{train}}\sim\text{Uniform}[-0.1,0.6]$ and $\varepsilon\sim\mathcal{N}(0,0.02^{2})$. Test data: $N_{\text{test}}=2048$ with $x_{\text{test}}\sim\text{Uniform}[-0.25,0.85]$. Predictions are evaluated on a dense grid over $x_{\text{grid}}\in[-0.5,1.5]$ to visualize behavior in-domain ([0.1,0.6][-0.1,0.6]) and out-of-domain (OOD).
Models.
All models share architecture $1\to 100\to 100\to 1$ with $\tanh$ activations. Training: 800 epochs, batch size 128, Adam optimizer with learning rate $5\times 10^{-4}$, fixed aleatoric noise $\sigma_{y}=0.02$.
  1. 1.
    Full-Rank BNN (Bayes by Backprop): All weights have diagonal Gaussian posteriors $q(w)=\mathcal{N}(\mu_{w},\sigma_{w}^{2})$. Prior: mixture $p(w)=0.5\mathcal{N}(0,2.0^{2})+0.5\mathcal{N}(0,e^{-6})$. Total trainable parameters: 20,802.
  2. 2.
    Low-Rank BNN: Hidden layer ($100\to 100$) uses factorization $W\approx AB^{\top}$ with rank $r=16$. Input ($1\to 100$) and output ($100\to 1$) layers remain full-rank. Same mixture prior. Total trainable parameters: 7,202 (65% parameter reduction).
Training uses KL annealing (ramped over 760 epochs) and $\beta=0.0001/N$ tempering to avoid KL dominance.
I.3.2 Results
Test Set Performance (MC averaged over $S=200$ weight samples).
Model Single Pass RMSE MC-200 RMSE
Full-Rank BNN 0.49623 0.49528
Low-Rank BNN 0.50732 0.50296
Table 19: Test set RMSE: Low-rank achieves comparable accuracy (0.5030 vs 0.4953) with 65% fewer parameters.
Epistemic Uncertainty: In-Domain vs. Out-of-Domain.
Both models expand their predictive intervals outside the training support, as expected from Bayesian posterior uncertainty. However, they differ in magnitude and rate of expansion:
Model In-Domain IQR Out-of-Domain IQR Ratio Parameters
Full-Rank BNN $0.0254$ $0.0483$ $1.90\times$ 20,802
Low-Rank BNN $0.0467$ $0.0940$ $2.01\times$ 7,202
Table 20: Median epistemic interquartile range (IQR) widths (5–95 percentile) computed over $S=100$ posterior samples. In-domain region: $x\in[0.0,0.5]$ (center of training range). Out-of-domain: $x\in[0.5,1.5]$ (beyond training support).
Conservative Uncertainty Behavior of Low-Rank.
Three insights emerge:
  1. 1.
    Wider absolute OOD band: Low-rank maintains significantly larger out-of-domain uncertainty ($0.0940$ vs $0.0483$), providing more conservative credible intervals where data are absent. This is a desirable property for safe prediction.
  2. 2.
    Sharper expansion ratio: Although in-domain uncertainty is higher in low-rank ($0.0467$ vs $0.0254$), the OOD/in-domain ratio is nearly identical ($2.01\times$ vs $1.90\times$). This shows low-rank preserves the qualitative epistemic sensitivity: uncertainty grows when leaving the training domain.
  3. 3.
    Stronger in-domain regularization: Low-rank’s higher baseline epistemic IQR ($0.0467$) reflects structured regularization from rank constraints. Rather than under-confident (overfitting) predictions, the low-rank posterior spreads mass more broadly, yielding conservative predictions even in-domain.
Conclusion.
Low-rank factorization achieves 65% parameter reduction (20,802 $\to$ 7,202) while preserving, and in fact improving, epistemic uncertainty quantification. The low-rank model exhibits:
  • Comparable test accuracy (RMSE 0.5030 vs 0.4953),
  • More conservative OOD behavior with nearly twice the absolute uncertainty band outside the training domain,
  • Similar ratio-based expansion, confirming qualitative epistemic sensitivity is maintained,
  • Higher baseline uncertainty even in-domain, yielding safer predictions when confidence is warranted but not assumed.
(a) Full-Rank BNN (20,802 params)
(b) Low-Rank BNN (7,202 params, rank=16)
[…]

### v2

@@
 For $\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)$ as in Definition E.1, the following bound holds:
-$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{\|W_{1}\|_{F}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\},$
+$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{B_{1}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\},$
 where:
@@
   • $h_{\ell}$ denotes the width of layer $\ell$,
-  • $r_{j}$ is the rank bound for layer $j$.
+  • $r_{j}$ is the rank bound for layer $j$,
+  • $B_{1}:=\sup_{f_{W}\in\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)}\|W_{1}\|_{F}$ is the worst-case Frobenius norm of the first-layer weight matrix over the class.
+Since $\mathrm{rank}(W_{1})\leq r_{1}$ and $\|W_{1}\|_{2}\leq C_{1}$ for all members of the class, we have $B_{1}\leq\sqrt{r_{1}}\,C_{1}$.
 Remark E.18 (Key features of the bound).
@@
   4. 4.
-    First layer: The term $\|W_{1}\|_{F}\sqrt{h_{1}}$ for the first layer appears because the first layer is not constrained by rank in the bound, only by spectral norm and Frobenius norm.
+    First layer: The term $B_{1}\sqrt{h_{1}}$ for the first layer differs in form from the $\sqrt{r_{j}h_{j}}$ terms for deeper layers.
+This is an artifact of the layer-peeling proof technique (specifically, how the chaining argument handles the input layer), not a reflection of a missing rank constraint.
+The rank of the first layer still enters through $B_{1}\leq\sqrt{r_{1}}\,C_{1}$, so that the first-layer contribution is at most $\sqrt{r_{1}}\,C_{1}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}$, which has the same $\sqrt{r_{1}}$ dependence as deeper layers.
 Remark E.19 (Comparison with standard Rademacher bounds).
@@
 As noted by Pinto et al.
-(2025), this unfavorable exponential factor in depth may be an artifact of the current proof technique rather than a limitation.
+(2025), this unfavorable exponential factor in depth may be an artifact of the current proof technique rather than a fundamental limitation.
 When comparing the rank-constrained bound to existing bounds, the key improvement comes from the rank-dependent terms $\sqrt{r_{j}h_{j}}$ rather than from the depth-dependent product, which appears in both settings.
@@
 Step 4: Apply the explicit bound from Theorem E.17.
-Substituting the explicit bound from Theorem E.17:
-$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{\|W_{1}\|_{F}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}.$
+Substituting the explicit bound from Theorem E.17, and using $B_{1}\leq\sqrt{r_{1}}\,C_{1}$:
+$\hat{\mathcal{G}}_{S_{x}}\big(\mathcal{F}^{\mathrm{Pinto}}_{D}(C,r)\big)\leq\frac{\|X\|_{F}}{m}\left\{\sqrt{r_{1}}\,C_{1}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}.$
 Step 5: Combine to obtain the final bound.
@@
 $\mathbb{E}\big[\ell(f_{\mathrm{BNN}}(x),y)\big]$ $\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f_{\mathrm{BNN}}(x_{j}),y_{j})$
-$\quad+\sqrt{\pi\mathcal{L}}\cdot\frac{\|X\|_{F}}{m}\left\{\|W_{1}\|_{F}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}$
+$\quad+\sqrt{\pi\mathcal{L}}\cdot\frac{\|X\|_{F}}{m}\left\{\sqrt{r_{1}}\,C_{1}\sqrt{h_{1}}\prod_{i=1}^{D}C_{0}C_{i}+\sum_{j=2}^{D}C_{j}\sqrt{r_{j}h_{j}}\prod_{i=j+1}^{D}C_{0}C_{i}\right\}$
 $\quad+3\sqrt{\frac{\log(2/\delta)}{2m}}.$
@@
 Proposition E.21 (Scaling Comparison).
-Consider a single-layer network with input dimension $d_{\text{in}}$, output dimension $d_{\text{out}}$, and rank $r$.
+Consider a single-layer network with input dimension $d_{\text{in}}$, output dimension $d_{\text{out}}$, and rank $r$. Let $\ell(\cdot,y)\in[0,1]$ be $\mathcal{L}$-Lipschitz in its first argument.
 PAC-Bayes bound (from Theorem 3.8): The generalization gap is bounded by
-$\text{Gap}_{\text{PAC}}=O\left(\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})}{m}}\right),$ (15)
-with rank dependence $\sqrt{r}$ and uniform $\frac{1}{\sqrt{m}}$ scaling in sample size (up to logarithmic factors).
+$\text{Gap}_{\text{PAC}}\;\leq\;\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})+\log(2\sqrt{m}/\delta)}{2m}},$ (15)
+where $C_{\max}$ is the maximum per-parameter KL divergence between the posterior and the prior. The structural scaling (suppressing logarithmic terms, $C_{\max}$, and constants) is
+$\text{Gap}_{\text{PAC}}=O\!\left(\sqrt{\frac{r(d_{\text{in}}+d_{\text{out}})}{m}}\right).$
 Gaussian complexity bound (from Theorem 3.9): The generalization gap is bounded by
-$\text{Gap}_{\text{GC}}=O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right)+O\left(\frac{1}{\sqrt{m}}\right),$ (16)
-where:
-  • The first term has rank dependence $\sqrt{r}$ and scales as $\frac{1}{m}$ (for fixed $\|X\|_{F}$);
-  • The second term (concentration) scales as $\frac{1}{\sqrt{m}}$ and is rank-independent;
-  • For large $m$, the $O(1/\sqrt{m})$ term dominates;
-  • For moderate $m$ with large $\|X\|_{F}$, the $O(1/m)$ term may dominate.
-If, in addition, inputs are uniformly bounded as $\|x_{j}\|_{2}\leq R$, then $\|X\|_{F}=O(\sqrt{m})$ and the rank-dependent term becomes $O(\sqrt{rd_{\text{out}}/m})$.
+$\text{Gap}_{\text{GC}}\;\leq\;\sqrt{\pi\mathcal{L}}\cdot\frac{\|X\|_{F}\cdot C_{0}C_{1}^{2}\cdot\sqrt{r\,d_{\text{out}}}}{m}\;+\;3\sqrt{\frac{\log(2/\delta)}{2m}},$ (16)
+where $C_{0}$ is the universal chaining constant and $C_{1}$ is the spectral norm bound.
+If inputs are uniformly bounded as $\|x_{j}\|_{2}\leq R$, then $\|X\|_{F}\leq R\sqrt{m}$ and the rank-dependent term simplifies to $O(\sqrt{r\,d_{\text{out}}/m})$.
 Proof.
 PAC-Bayes analysis. The complexity term in Theorem 3.8 is
-$\sqrt{\frac{\mathrm{KL}(Q_{\text{LR}}\|P_{\text{LR}})+\log(2\sqrt{m}/\delta)}{2m}}\leq\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})+\log(2\sqrt{m}/\delta)}{2m}}.$
-Ignoring logarithmic terms and constants, this scales as
-$O\left(\sqrt{\frac{r(d_{\text{in}}+d_{\text{out}})}{m}}\right)=\sqrt{r}\cdot O\left(\sqrt{\frac{d_{\text{in}}+d_{\text{out}}}{m}}\right),$
-so the rank dependence is $\sqrt{r}$ and the sample-size dependence is $1/\sqrt{m}$.
+$\sqrt{\frac{\mathrm{KL}(Q_{\text{LR}}\|P_{\text{LR}})+\log(2\sqrt{m}/\delta)}{2m}}\leq\sqrt{\frac{C_{\max}\cdot r(d_{\text{in}}+d_{\text{out}})+\log(2\sqrt{m}/\delta)}{2m}},$
+which yields the expression in (15). Dropping logarithmic terms and constants gives the structural scaling $O(\sqrt{r(d_{\text{in}}+d_{\text{out}})/m})$.
 Gaussian complexity analysis. For a single-layer network (depth $D=1$), Theorem E.17 gives
-$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})\leq\frac{\|X\|_{F}}{m}\|W_{1}\|_{F}\sqrt{h_{1}}C_{0}C_{1},$
-where $h_{1}=d_{\text{out}}$. For a rank-$r$ matrix $W_{1}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ with $\|W_{1}\|_{2}\leq C_{1}$, we have
-$\|W_{1}\|_{F}\leq\sqrt{r}\,\|W_{1}\|_{2}\leq\sqrt{r}C_{1}.$
-Substituting with $h_{1}=d_{\text{out}}$,
-$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})$ $\leq\frac{\|X\|_{F}}{m}\cdot\sqrt{r}C_{1}\cdot\sqrt{d_{\text{out}}}\cdot C_{0}C_{1}$
-$=\frac{\|X\|_{F}\cdot C_{0}C_{1}^{2}\cdot\sqrt{rd_{\text{out}}}}{m}$
-$=O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right).$
+$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})\leq\frac{\|X\|_{F}}{m}\cdot B_{1}\sqrt{h_{1}}\cdot C_{0}C_{1},$
+where $h_{1}=d_{\text{out}}$. Since $B_{1}\leq\sqrt{r}\,C_{1}$ (see Theorem E.17), substituting yields
+$\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}^{\text{Pinto}}_{1})$ $\leq\frac{\|X\|_{F}}{m}\cdot\sqrt{r}\,C_{1}\cdot\sqrt{d_{\text{out}}}\cdot C_{0}C_{1}=\frac{\|X\|_{F}\cdot C_{0}C_{1}^{2}\cdot\sqrt{r\,d_{\text{out}}}}{m}.$
 From Theorem E.15, the full generalization bound is
-$\mathbb{E}[\ell(f(x),y)]$ $\leq\hat{L}(f)+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F})+3\sqrt{\frac{\log(2/\delta)}{2m}}$
-$=\hat{L}(f)+O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right)+O\left(\frac{1}{\sqrt{m}}\right),$
-where the first term is rank-dependent and decays as $1/m$ (for fixed $\|X\|_{F}$), while the second term is rank-independent and decays as $1/\sqrt{m}$.
-This yields the claimed form for $\text{Gap}_{\text{GC}}$.
-If we additionally assume $\|x_{j}\|_{2}\leq R$ for all $j$, then
-$\|X\|_{F}^{2}=\sum_{j=1}^{m}\|x_{j}\|_{2}^{2}\leq mR^{2}\quad\Rightarrow\quad\|X\|_{F}\leq R\sqrt{m},$
-so the rank-dependent term simplifies to
-$O\left(\frac{\|X\|_{F}\sqrt{rd_{\text{out}}}}{m}\right)=O\left(\frac{\sqrt{m}\sqrt{rd_{\text{out}}}}{m}\right)=O\left(\sqrt{\frac{rd_{\text{out}}}{m}}\right).$
+$\text{Gap}_{\text{GC}}\leq\sqrt{\pi\mathcal{L}}\cdot\frac{\|X\|_{F}\cdot C_{0}C_{1}^{2}\cdot\sqrt{r\,d_{\text{out}}}}{m}+3\sqrt{\frac{\log(2/\delta)}{2m}},$
+which is (16). The first (rank-dependent) term decays as $1/m$ for fixed $\|X\|_{F}$, while the second (concentration) term decays as $1/\sqrt{m}$.
+If we additionally assume $\|x_{j}\|_{2}\leq R$ for all $j$, then $\|X\|_{F}\leq R\sqrt{m}$, so the rank-dependent term becomes
+$\sqrt{\pi\mathcal{L}}\cdot C_{0}C_{1}^{2}\cdot R\cdot\sqrt{\frac{r\,d_{\text{out}}}{m}}.$
 Remark E.22 (Structural differences).
@@
 $\text{Gap}_{\text{GC}}=\underbrace{O\left(\frac{\text{Complexity}\times\|X\|_{F}}{m}\right)}_{\text{faster decay in }m}+\underbrace{O\left(\frac{1}{\sqrt{m}}\right)}_{\text{slower decay in }m}.$
-For large $m$, the slower-decaying $O(1/\sqrt{m})$ term dominates; for moderate $m$ and large $\|X\|_{F}\sqrt{rd_{\text{out}}}$, the $O(1/m)$ term can dominate.
+For large $m$, the slower-decaying $O(1/\sqrt{m})$ concentration term dominates both bounds, and the comparison reduces to constants.
+For moderate $m$, the different structural dependences on architecture and data can make either bound tighter.
+[…]
 E.9.1 Complementary Strengths
@@
   1. 1.
-    Large sample regime: For large $m$, both bounds scale as $O(1/\sqrt{m})$, but PAC-Bayes may have better constants when the posterior is concentrated (small $C_{\max}$).
+    Posterior concentration: When the posterior is highly concentrated around a good solution, $C_{\max}$ can be very small, making the KL term tight.
+In the extreme case of a near-deterministic posterior, the PAC-Bayes gap approaches zero while the Gaussian complexity bound remains fixed by the class geometry.
   2. 2.
-    Posterior concentration: When the posterior is highly concentrated around a good solution, $C_{\max}$ can be very small, making the KL term tight.
+    Large input norms: PAC-Bayes bounds do not depend on $\|X\|_{F}$, which is advantageous when inputs have very large norm or heavy tails.
+The Gaussian complexity bound’s leading term grows linearly in $\|X\|_{F}$.
   3. 3.
-    Large input norms: PAC-Bayes bounds do not depend on $\|X\|_{F}$, which is advantageous when inputs have very large norm or heavy tails.
+    High-dimensional inputs with moderate rank: Looking at the full expressions, PAC-Bayes scales with $r(d_{\text{in}}+d_{\text{out}})$ while Gaussian complexity (with bounded inputs) scales with $C_{0}^{2}C_{1}^{4}R^{2}\cdot r\,d_{\text{out}}$.
+When $d_{\text{in}}\gg d_{\text{out}}$ and $C_{0}^{2}C_{1}^{4}R^{2}$ is large, PAC-Bayes can still be favorable despite the $d_{\text{in}}$ term, because it avoids the spectral norm constants entirely.
+Concretely, PAC-Bayes is structurally favorable when $C_{\max}(d_{\text{in}}+d_{\text{out}})\lesssim\pi\mathcal{L}\,C_{0}^{2}C_{1}^{4}R^{2}\,d_{\text{out}}$, i.e., when the posterior is sufficiently concentrated relative to the spectral norm and Lipschitz constants.
+Remark E.24 (When Gaussian complexity is stronger).
+Gaussian complexity bounds can be tighter when:
+  1. 1.
+    Diffuse posteriors with bounded spectral norms: When $C_{\max}$ is large (the posterior has not concentrated tightly), the PAC-Bayes KL term can be vacuous.
+The Gaussian complexity bound depends on the spectral norm constraints $C_{1}$, not on posterior concentration, and can remain informative even when the posterior is spread over a large region of parameter space.
+  2. 2.
+    Normalized inputs with small spectral norm constants: When $\|x_{j}\|_{2}\leq R$ with $R$ moderate and spectral norms are well-controlled ($C_{1}$ not too large), the Gaussian complexity bound’s leading constant $\sqrt{\pi\mathcal{L}}\,C_{0}C_{1}^{2}R$ can be smaller than $\sqrt{C_{\max}/2}$.
+This is the regime where the Gaussian complexity bound is numerically tighter.
+  3. 3.
+    No prior specification required: Unlike PAC-Bayes, which requires choosing a prior $P$ and where the bound quality depends sensitively on this choice, Gaussian complexity bounds depend only on the realized network architecture and data.
+This makes them valuable as a prior-free sanity check.
   4. 4.
-    Dimension regime: When $r(d_{\text{in}}+d_{\text{out}})$ is small relative to $\|X\|_{F}^{2}d_{\text{out}}$, the PAC-Bayes bound can be more favorable.
-Remark E.24 (When Gaussian complexity is stronger).
-Gaussian complexity bounds might be tighter when:
-  1. 1.
-    Small to moderate sample sizes: When $m$ is not too large and $\|X\|_{F}\sqrt{rd_{\text{out}}}=O(\sqrt{m})$, the $O(1/m)$ term provides faster decay than the $O(1/\sqrt{m})$ concentration term.
-  2. 2.
-    Normalized or bounded inputs: When $\|X\|_{F}=O(\sqrt{m})$ (e.g., inputs normalized to unit norm), the rank-dependent term becomes $O(\sqrt{rd_{\text{out}}/m})$, which can be tighter than PAC-Bayes when $rd_{\text{out}}\ll r(d_{\text{in}}+d_{\text{out}})$.
-  3. 3.
-    No prior specification required: Unlike PAC-Bayes, which requires choosing a prior $P$, Gaussian complexity bounds depend only on the realized network architecture and data.
-  4. 4.
-    Architectural transparency: The bound explicitly shows how rank, spectral norms, and layer dimensions interact through the network depth, providing architectural guidance.
-  5. 5.
-    Connection to deterministic theory: Gaussian complexity bounds inherit the full machinery of deterministic generalization theory, including recent advances in rank-sensitive analysis.
+    Architectural transparency: The bound explicitly shows how rank, spectral norms, and layer dimensions interact through the network depth, providing guidance for architectural design that is not available from the PAC-Bayes bound’s KL-divergence summary.
 Remark E.25 (Rank dependence).
@@
 Neither bound exhibits linear $r$ dependence; the rank always appears under a square root in the leading generalization term.
-Remark E.26.
+Remark E.26 (Complementary perspectives on generalization).
 The key insight is that these bounds measure different aspects of generalization:
@@
 This section presents two alternative approaches that allow our framework to accommodate high-probability bounds rather than requiring almost-sure bounds.
+Both approaches require an additional regularity condition on the loss function:
+Assumption E.27 (Convex loss).
+The loss function $\ell(\cdot,y):\mathbb{R}^{k}\to[0,1]$ is convex in its first argument for every $y$.
+Remark E.28.
+Assumption E.27 is satisfied by most standard loss functions used in practice, including the squared loss $\ell(\hat{y},y)=\|\hat{y}-y\|^{2}$ (after rescaling to $[0,1]$), the cross-entropy loss (when composed with the softmax, the loss is convex in the logit vector), and the hinge loss.
+The convexity requirement is needed because the BNN predictor $f_{\mathrm{BNN}}$ is a mixture of $f_{\mathrm{BNN}}^{\mathrm{good}}$ and $f_{\mathrm{BNN}}^{\mathrm{bad}}$, and we need Jensen’s inequality to bound the loss of the mixture in terms of the losses of its components.
 E.10.1 Setup and Notation
@@
 $\mathbb{P}_{B_{i}\sim q_{B}(\cdot\mid\theta_{i})}\left(\|B_{i}\|_{2}\leq C_{i}^{B}\right)$ $\geq 1-\delta_{i}^{B}.$ (18)
-Define the ”good event” $\mathcal{E}_{\text{good}}$ as the intersection of all these events across all $D$ layers:
+Define the “good event” $\mathcal{E}_{\text{good}}$ as the intersection of all these events across all $D$ layers:
 $\mathcal{E}_{\text{good}}:=\bigcap_{i=1}^{D}\left\{\|A_{i}\|_{2}\leq C_{i}^{A}\right\}\cap\left\{\|B_{i}\|_{2}\leq C_{i}^{B}\right\}.$ (19)
@@
 When $\mathcal{E}_{\text{good}}$ occurs, all factor matrices satisfy their spectral norm bounds, and consequently all induced weight matrices $W_{i}=A_{i}B_{i}^{\top}$ satisfy $\|W_{i}\|_{2}\leq C_{i}:=C_{i}^{A}C_{i}^{B}$ and $\mathrm{rank}(W_{i})\leq r_{i}$ by the same argument as Lemma E.5.
+We write $p:=\mathbb{P}(\mathcal{E}_{\text{good}})\geq 1-\delta_{\text{total}}$ throughout this section.
 E.10.2 Approach A: Decomposition of the BNN Predictor
+Function-level decomposition.
 The BNN predictor is defined as an expectation over the full posterior:
 $f_{\mathrm{BNN}}(x;\theta)=\mathbb{E}_{(A,B)\sim q(\cdot\mid\theta)}[f_{A,B}(x)].$ (21)
-We can decompose this expectation by conditioning on whether $\mathcal{E}_{\text{good}}$ occurs:
-$f_{\mathrm{BNN}}(x;\theta)=\mathbb{E}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}]\cdot\mathbb{P}(\mathcal{E}_{\text{good}})+\mathbb{E}[f_{A,B}(x)\mid\mathcal{E}_{\text{bad}}]\cdot\mathbb{P}(\mathcal{E}_{\text{bad}}),$ (22)
-where $\mathcal{E}_{\text{bad}}=\mathcal{E}_{\text{good}}^{c}$.
-Analysis of the ”good” term.
-Define the conditional predictor:
-$f_{\mathrm{BNN}}^{\text{good}}(x;\theta):=\mathbb{E}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}].$ (23)
-This conditional expectation is taken only over samples that satisfy the spectral norm bounds.
-Replacing the full posterior with the conditional distribution given $\mathcal{E}_{\text{good}}$, we have by the same argument as Lemma E.11:
+By the law of total expectation applied to the posterior distribution $q$, we can decompose this as:
+$f_{\mathrm{BNN}}(x;\theta)=p\cdot f_{\mathrm{BNN}}^{\text{good}}(x;\theta)+(1-p)\cdot f_{\mathrm{BNN}}^{\text{bad}}(x;\theta),$ (22)
+where:
+$f_{\mathrm{BNN}}^{\text{good}}(x;\theta)$ $:=\mathbb{E}_{(A,B)\sim q}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}],$ (23)
+$f_{\mathrm{BNN}}^{\text{bad}}(x;\theta)$ $:=\mathbb{E}_{(A,B)\sim q}[f_{A,B}(x)\mid\mathcal{E}_{\text{good}}^{c}].$ (24)
+Note: Both $f_{\mathrm{BNN}}^{\text{good}}$ and $f_{\mathrm{BNN}}^{\text{bad}}$ are deterministic functions of $x$ (given the variational parameters $\theta$).
+The conditioning in (23)–(24) is over the posterior $(A,B)\sim q$, not over the data distribution.
+Equation (22) is a pointwise identity for each $x$, expressing $f_{\mathrm{BNN}}(x)$ as a convex combination of two deterministic functions.
+Analysis of the “good” term.
+The conditional expectation $f_{\mathrm{BNN}}^{\text{good}}$ is taken over the conditional distribution $q(\cdot\mid\theta,\mathcal{E}_{\text{good}})$, which is supported on factor matrices satisfying the spectral norm bounds.
+By the same argument as Lemma E.11 (replacing $q$ by $q(\cdot\mid\mathcal{E}_{\text{good}})$):
 $f_{\mathrm{BNN}}^{\text{good}}(\cdot;\theta)\in\overline{\mathrm{conv}(\mathcal{F}_{D}^{\mathrm{supp},\text{good}}(\theta))},$ (25)
-where $\mathcal{F}_{D}^{\mathrm{supp},\text{good}}(\theta)$ consists of network functions realized by samples in $\mathcal{E}_{\text{good}}$.
-Since all such samples satisfy the spectral norm constraints by construction, we have:
+where $\mathcal{F}_{D}^{\mathrm{supp},\text{good}}(\theta)$ consists of network functions realized by factor pairs in the support of $q(\cdot\mid\mathcal{E}_{\text{good}})$.
+Since all such pairs satisfy the spectral norm constraints by definition of $\mathcal{E}_{\text{good}}$:
 $\mathcal{F}_{D}^{\mathrm{supp},\text{good}}(\theta)\subseteq\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r).$ (26)
-Therefore, by the same arguments as Proposition E.13:
-$\hat{\mathcal{G}}_{S_{x}}(\{f_{\mathrm{BNN}}^{\text{good}}\})\leq\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r)).$ (26)
-Analysis of the ”bad” term.
-The contribution from $\mathcal{E}_{\text{bad}}$ has small probability $\delta_{\text{total}}$.
-If the loss function $\ell$ is bounded in $[0,1]$, this contributes at most $\delta_{\text{total}}$ to the expected loss.
+By the same arguments as Proposition E.13, the Gaussian complexity of the class containing $f_{\mathrm{BNN}}^{\text{good}}$ is bounded by that of Pinto’s class.
+Applying Theorem E.15, we obtain that with probability at least $1-\delta$ over $S\sim\mathcal{D}^{m}$:
+$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)]\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f_{\mathrm{BNN}}^{\text{good}}(x_{j}),y_{j})+\sqrt{\pi\mathcal{L}}\;\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r))+3\sqrt{\frac{\log(2/\delta)}{2m}}.$ (27)
+From $f_{\mathrm{BNN}}^{\text{good}}$ to $f_{\mathrm{BNN}}$: applying loss convexity.
+We now bound the population loss of the actual deployed predictor $f_{\mathrm{BNN}}$. By (22), for each $(x,y)$:
+$f_{\mathrm{BNN}}(x)=p\cdot f_{\mathrm{BNN}}^{\text{good}}(x)+(1-p)\cdot f_{\mathrm{BNN}}^{\text{bad}}(x).$
+Under Assumption E.27, $\ell(\cdot,y)$ is convex in its first argument, so by Jensen’s inequality:
+$\ell(f_{\mathrm{BNN}}(x),y)$ $\leq p\cdot\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)+(1-p)\cdot\ell(f_{\mathrm{BNN}}^{\text{bad}}(x),y)$
+$\leq p\cdot\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)+(1-p)\cdot 1,$ (28)
+where in the second inequality we used $\ell\in[0,1]$.
+Taking the expectation over $(x,y)\sim\mathcal{D}$:
+$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]\leq p\cdot\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)]+(1-p)\leq\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)]+\delta_{\text{total}},$ (29)
+where the last step uses $1-p\leq\delta_{\text{total}}$ and $p\leq 1$.
 Combined generalization bound.
-Let $\ell:\mathbb{R}^{k}\times\mathbb{R}^{k}\to[0,1]$ be an $\mathcal{L}$-Lipschitz loss function. The expected loss of the BNN predictor satisfies:
-$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $=\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)\mid\mathcal{E}_{\text{good}}]\cdot\mathbb{P}(\mathcal{E}_{\text{good}})$
-$\quad+\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)\mid\mathcal{E}_{\text{bad}}]\cdot\mathbb{P}(\mathcal{E}_{\text{bad}})$
-$\leq\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}^{\text{good}}(x),y)]+1\cdot\delta_{\text{total}}.$ (27)
-Applying the generalization bound from Theorem E.15 to $f_{\mathrm{BNN}}^{\text{good}}$ (following the proof of Theorem 3.9):
-$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $\leq\frac{1}{m}\sum_{i=1}^{m}\ell(f_{\mathrm{BNN}}^{\text{good}}(x_{i}),y_{i})$
-$\quad+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r))+3\sqrt{\frac{\log(2/\delta)}{2m}}+\delta_{\text{total}}.$ (28)
+Substituting (27) into (29):
+Proposition E.29 (Generalization bound under high-probability spectral norm constraints).
+Let $\ell(\cdot,y)\in[0,1]$ be $\mathcal{L}$-Lipschitz and convex in its first argument (Assumptions in Theorem E.15 and Assumption E.27).
+Let high-probability bounds (17)–(18) hold with $\delta_{\text{total}}=\sum_{i=1}^{D}(\delta_{i}^{A}+\delta_{i}^{B})$.
+Then with probability at least $1-\delta$ over $S\sim\mathcal{D}^{m}$:
+$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f_{\mathrm{BNN}}^{\mathrm{good}}(x_{j}),y_{j})$
+$\quad+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C,r))+3\sqrt{\frac{\log(2/\delta)}{2m}}+\delta_{\text{total}},$ (30)
+where $C_{i}=C_{i}^{A}C_{i}^{B}$ and $f_{\mathrm{BNN}}^{\mathrm{good}}$ is defined in (23).
+Remark E.30 (Empirical loss of $f_{\mathrm{BNN}}^{\mathrm{good}}$ vs. $f_{\mathrm{BNN}}$).
+The empirical loss term involves $f_{\mathrm{BNN}}^{\mathrm{good}}$ rather than $f_{\mathrm{BNN}}$, because the generalization bound from Theorem E.15 is applied to $f_{\mathrm{BNN}}^{\mathrm{good}}$ (which belongs to the controlled function class).
+In practice, $f_{\mathrm{BNN}}^{\mathrm{good}}$ can be computed via Monte Carlo sampling from the truncated posterior $q(\cdot\mid\mathcal{E}_{\text{good}})$, i.e., by rejection sampling: draw $(A,B)\sim q$ and retain only samples satisfying the spectral norm bounds.
+When $\delta_{\text{total}}$ is small, the rejection rate is low and $f_{\mathrm{BNN}}^{\mathrm{good}}\approx f_{\mathrm{BNN}}$.
 Remark E.31 (Practical interpretation).
-The additional term $\delta_{\text{total}}$ represents the ”cost” of using high-probability bounds instead of almost-sure bounds.
+The additional term $\delta_{\text{total}}$ represents the “cost” of using high-probability bounds instead of almost-sure bounds.
 For Gaussian posteriors with Van Handel-type bounds (see Remark E.35), choosing $\delta_{i}^{A}=\delta_{i}^{B}=\delta_{0}/(2D)$ for each layer gives $\delta_{\text{total}}=\delta_{0}$.
 The spectral norm bounds $C_{i}^{A}$ and $C_{i}^{B}$ then scale as $O(\sqrt{\log(D/\delta_{0})})$, which introduces a mild logarithmic dependence on the depth and confidence level.
+Remark E.32 (Why convexity is necessary).
+Without Assumption E.27, the bound (28) does not hold. Using only Lipschitz continuity gives
+$|\ell(f_{\mathrm{BNN}}(x),y)-\ell(f_{\mathrm{BNN}}^{\mathrm{good}}(x),y)|\leq\mathcal{L}\|f_{\mathrm{BNN}}(x)-f_{\mathrm{BNN}}^{\mathrm{good}}(x)\|_{2}=\mathcal{L}(1-p)\|f_{\mathrm{BNN}}^{\mathrm{bad}}(x)-f_{\mathrm{BNN}}^{\mathrm{good}}(x)\|_{2},$
+which requires bounding $\|f_{\mathrm{BNN}}^{\mathrm{bad}}(x)\|_{2}$—the expected output under the “bad” event where spectral norm constraints are violated.
+This quantity can be arbitrarily large for unbounded variational families, making the Lipschitz approach impractical without additional assumptions on the tails of $q$.
+Loss convexity avoids this issue entirely by exploiting the $[0,1]$-boundedness of $\ell$.
 E.10.3 Approach B: Modified Function Class with High-Probability Inclusion
@@
   1. 1.
-    $\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta)\subseteq\mathcal{F}_{D}^{\mathrm{Pinto}}(C(\delta),r)$ by construction, where $C_{i}(\delta)=C_{i}^{A}(\delta)C_{i}^{B}(\delta)$ (all functions satisfy the spectral norm and rank constraints).
+    $\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta)\subseteq\mathcal{F}_{D}^{\mathrm{Pinto}}(C(\delta),r)$ by construction, where $C_{i}(\delta)=C_{i}^{A}(\delta)C_{i}^{B}(\delta)$.
   2. 2.
-    The BNN predictor restricted to $\mathcal{E}_{\text{good}}$ lies in $\overline{\mathrm{conv}(\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta))}$ by the same argument as Lemma E.11.
+    $f_{\mathrm{BNN}}^{\mathrm{good}}(\cdot;\theta)\in\overline{\mathrm{conv}(\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta))}$ by the same argument as Lemma E.11, applied to the conditional posterior $q(\cdot\mid\mathcal{E}_{\text{good}})$.
   3. 3.
@@
 Generalization bound.
-By the same reasoning as Approach A:
-$\mathbb{E}[\ell(f_{\mathrm{BNN}}(x),y))]$ $\leq\frac{1}{m}\sum_{i=1}^{m}\ell(f_{\mathrm{BNN}}(x_{i}),y_{i})$
+By the same reasoning as Approach A—decomposing $f_{\mathrm{BNN}}$ as in (22), applying Assumption E.27 to obtain (29), and bounding the population loss of $f_{\mathrm{BNN}}^{\mathrm{good}}$ via Theorem E.15—we obtain:
+$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f_{\mathrm{BNN}}^{\mathrm{good}}(x_{j}),y_{j})$
 $\quad+\sqrt{\pi\mathcal{L}}\cdot\hat{\mathcal{G}}_{S_{x}}(\mathcal{F}_{D}^{\mathrm{Pinto}}(C(\delta),r))+3\sqrt{\frac{\log(2/\delta^{\prime})}{2m}}+\delta,$ (33)
@@
 Remark E.33 (Comparison of approaches).
-Both approaches yield the same final bound structure, with an additive $\delta$ term representing the high-probability relaxation.
-  • Approach A is conceptually cleaner: it decomposes the predictor into good and bad events.
+Both approaches yield the same final bound structure, with an additive $\delta$ term from the high-probability relaxation and a convexity requirement on the loss (Assumption E.27).
+  • Approach A is conceptually cleaner: it decomposes the predictor into good and bad components at the function level.
   • Approach B is more direct: it modifies the function class definition to explicitly encode the high-probability constraint.
@@
 This is often more practical than truncation, as it requires no modification to the variational family.
+This approach additionally requires loss convexity (Assumption E.27).
 Practical recommendations.
@@
   • Use Van Handel bounds to guide the choice of $C_{i}^{A}$ and $C_{i}^{B}$ based on the variational posterior parameters $\sigma_{i}$ and $\mu_{i}$.
-  • If using unbounded variational families without projection, apply the high-probability analysis from Section E.10 with appropriately chosen $\delta_{\text{total}}$.
+  • If using unbounded variational families without projection, apply the high-probability analysis from Section E.10 with appropriately chosen $\delta_{\text{total}}$ and ensure the loss function satisfies Assumption E.27.
 Remark E.36 (Tightness of Van Handel bounds).